### v1

# 2cm2cm1.5cm1.5cm

Abstract.
Finding a largest Euclidean ball in a given convex body $K$ in $\mathbb{R}^{d}$, and finding the largest volume ellipsoid in $K$ are two problems of very different nature. The first is a Euclidean problem, where similar copies of the origin centered closed unit ball are considered, whereas in the second, we search among affine copies of $\mathbf{B}^{d}$. In this paper, we interpolate between the two problems by considering ellipsoids of revolution. We generalize the problem by looking at pairs of convex bodies $K$ and $L$ and the largest volume affine image of $K$ contained in $L$, where we make certain restrictions on the affine transformation.
Key words and phrases:
John ellipsoid, body of revolution, convexity, affine image
2020 Mathematics Subject Classification:
Primary 52A23; Secondary 52A40, 46T99
1. Introduction
Finding the largest Euclidean ball contained in a convex body $K\subset\mathbb{R}^{d}$ and finding the largest volume ellipsoid contained in $K$ are problems of fundamentally different nature. The former is a Euclidean problem involving scaled copies of the origin-centered unit ball $\mathbf{B}^{d}$, while the latter involves all affine images of $\mathbf{B}^{d}$. In this paper, we study a natural interpolation between these two settings by considering ellipsoids of revolution. More generally, we investigate pairs of convex bodies $K$ and $L$ and aim to determine the largest volume affine image of $K$ contained in $L$, subject to specific restrictions on the class of admissible affine transformations. Our starting point is John’s classical theorem [Joh48, Joh14] characterizing the maximal volume ellipsoid in a convex body and its generalizations to pairs of convex bodies [Lew79, TJ89, BR02, GLMP04]. Our goal is to initiate the development of a similar theory for ellipsoids of revolution and for constrained affine images in this more general setting.
1.1. Ellipsoids of revolution
In order to introduce ellipsoids of revolution, we start with defining a special class of linear operators.
Definition 1.1.
For a linear subspace $F$ in $\mathbb{R}^{d}$, we call an invertible linear operator $A$ on $\mathbb{R}^{d}$ an $F$-operator, if both $F$ and ${F}^{\bot}$ are invariant subspaces of $A$, and the restriction $A|_{{F}^{\bot}}$ of $A$ on the orthogonal complement ${F}^{\bot}$ of $F$ is a (nonzero) multiple of the identity operator $\mathrm{Id}_{{F}^{\bot}}$ on ${F}^{\bot}$.
Note that if we take an orthonormal basis of $F$, and extend it to an orthonormal basis of $\mathbb{R}^{d}$, then the matrix of an $F$ operator takes the form
$M=\begin{pmatrix}{M_{1}}&\mathbf{0}_{s\times(d-s)}\\ \mathbf{0}_{(d-s)\times s}&\mu\mathrm{Id}_{d-s}\end{pmatrix},$
where $M_{1}\in\mathbb{R}^{s\times s},$ $\mu\in\mathbb{R}\setminus\{0\},$ $\mathbf{0}_{s\times(d-s)}$ and $\mathbf{0}_{(d-s)\times s}$ are zero matrices of the appropriate sizes, and $\mathrm{Id}_{d-s}$ denotes the identity matrix of size $d-s$.
Definition 1.2.
Let $F$ be a linear subspace in $\mathbb{R}^{d}$ and $S$ be any subset of $\mathbb{R}^{d}$. We will refer to the family of affine transformations
$\mathcal{A}_{F}=\left\{x\mapsto Ax+z:\;A\text{ is an }F\text{-operator, }z\in\mathbb{R}^{d}\right\}$
as the affine transformations with axis $F$, and the family of sets
$\mathcal{A}_{F}\!\left(S\right)=\left\{AS+z:\;A\text{ is an }F\text{-operator,}z\in\mathbb{R}^{d}\right\}$
as the positions of $S$ with axis $F$.
In particular, we call members of $\mathcal{A}_{F}\!\left(\mathbf{B}^{d}\right)$ ellipsoids of revolution with axis $F$.
The naming is justified, since if $E\in\mathcal{A}_{F}\!\left(\mathbf{B}^{d}\right)$, then $E$ is an ellipsoid all of whose non-degenerate sections by translates of ${F}^{\bot}$ are $(d-\dim{F})$-dimensional balls. Our first problem is the following.
Problem 1 (John’s problem for ellipsoids with a fixed axis of revolution).
For a given convex body (that is, a closed, convex set with non-empty interior) $K$ in $\mathbb{R}^{d}$ and a linear subspace $F$ in $\mathbb{R}^{d}$, find a largest volume ellipsoid of revolution with axis $F$ contained in $K$.
Clearly, when $F=\{\mathbf{0}\}$ is the 0-dimensional subspace, then this problem is equivalent to finding a largest volume Euclidean ball in $K$. On the other hand, when $F=\mathbb{R}^{d}$, then it is exactly John’s problem of finding the largest volume ellipsoid contained in $K$.
In John’s problem for ellipsoids with a fixed axis of revolution, the subspace $F$ is fixed. It is natural to allow for the variation of $F$, which leads us to our second problem.
Problem 2 (John’s problem for ellipsoids with any axis of revolution).
Let $K$ be a convex body in $\mathbb{R}^{d}$ and $s\in\{0,1,\dots,d\}$. Find a largest volume ellipsoid of revolution in $K$ with any $s$-dimensional axis.
1.2. A more general problem
The problems above are motivated by John’s fundamental theorem [Joh48, Joh14], where a necessary condition is given for the ball $\mathbf{B}^{d}$ to be the maximal volume ellipsoid contained in a convex body $K$. The sufficiency of the same condition was noted by K. Ball [Bal97].
A natural generalization of this question has been studied by Lewis [Lew79], V. Milman [TJ89, Theorem 14.5], Giannopoulos, Perissinaki, and Tsolomitis [GPT01], Bastero and Romance [BR02], and Gordon, Litvak, Meyer, and Pajor [GLMP04]: given convex bodies $K$ and $L$ in $\mathbb{R}^{d}$, find a maximal volume affine image of $K$ that is contained in $L$. In this work, we study the same problem under additional restrictions on the affine transformation.
Problem 3 (John’s general problem with fixed axis).
For given convex bodies $K$ and $L$, and a linear subspace $F$ in $\mathbb{R}^{d}$, find a largest volume member of $\mathcal{A}_{F}\!\left(K\right)$ contained in $L$.
Clearly, when $F=\{\mathbf{0}\}$ is the 0-dimensional subspace, then this problem is equivalent to finding a largest homothet of $K$ in $L$. On the other hand, when $F=\mathbb{R}^{d}$, then it is exactly the problem described in the paragraph preceding Problem 3.
Definition 1.3.
Let $F$ be a linear subspace in $\mathbb{R}^{d}$ and $S$ be any subset of $\mathbb{R}^{d}$. We will refer to the family of affine transformations
$\mathcal{R}_{F}=\left\{x\mapsto RAx+z:\;R\in O(d),A\text{ is an }F\text{-operator, }z\in\mathbb{R}^{d}\right\},$
where $O(d)$ denotes the orthogonal group on $\mathbb{R}^{d}$, as the rotated affine transformations with axis $F$, and the family of sets
$\mathcal{R}_{F}\!\left(S\right)=\left\{RAS+z:\;R\in O(d),A\text{ is an }F\text{-operator, }z\in\mathbb{R}^{d}\right\},$
as the rotated positions of $S$ with axis $F$.
Problem 4 (John’s general problem with rotated axis).
Let $K$ and $L$ be convex bodies in $\mathbb{R}^{d},$ and fix a linear subspace $F$ in $\mathbb{R}^{d}$. Find a largest volume rotated position of $K$ with axis $F$ that is contained in $L$.
Note a subtle but very important difference between this general case and ellipsoids. To define ellipsoids of revolution with an $s$-dimensional axis, we only need to specify the dimension $s$ of the axis, but not the axis $F$. In the general case introduced here, we fix $F$, apply an $F$-operator $A$ and then allow a rotation $R$. The advantage of this is that in this way, we can specify what ‘part’ of $K$ we allow to be changed by an affine transformation (the $F$-component), and what part (the ${F}^{\bot}$ component) is rigid, that is, it only allows similarities. A simple illustration is the task of finding the largest area isosceles triangle contained in a planar set $L$. In our framework, we simply take $K$ to be a regular triangle, and $F$ to be one of its latitude lines. Then $\mathcal{R}_{F}\!\left(K\right)$ is exactly the set of isosceles triangles.
When $F=\{\mathbf{0}\}$, then Problem 4 is the problem discussed at the beginning of this subsection considered by Lewis and others. On the other hand, when $F=\mathbb{R}^{d}$, then we are looking for the largest similar copy of $K$ in $L$.
1.3. Results
Our first goal is to find conditions of optimality in Problems 1, 2, 3 and 4 similar to John’s condition in the case of ellipsoids and the condition given by Gordon, Litvak, Meyer and Pajor [GLMP04] in the general case. We start with Problems 3 and 4, as the optimality conditions in the case of ellipsoids of revolution will be obtained as corollaries to these general results.
Let $K$ and $L$ be convex bodies in $\mathbb{R}^{d}$ with $K\subseteq L$. Following [GLMP04], we call $(v,p)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$ a contact pair of $K$ and $L$ if $v\in\operatorname{bd}K\cap\operatorname{bd}L,$ and $p\in\operatorname{bd}{K}^{\circ}\cap\operatorname{bd}{L}^{\circ}$ normalized by the condition $\left\langle p,v\right\rangle=1$. In other words, $p$ is an outer normal vector of a common support hyperplane of $K$ and $L$ at $v$, with the appropriate normalization. We use $P_{F}$ to denote the orthogonal projection onto $F$.
Our necessary condition of optimality for Problem 3 is the following.
Theorem 1.4 (Condition for John’s general problem with fixed axis).
Let $K,L\subset\mathbb{R}^{d}$ be convex bodies such that $K\subseteq L$, and let $F$ be a linear subspace in $\mathbb{R}^{d}$. Assume that $K$ is a solution to John’s general problem with fixed axis $F$ for $K$ and $L$, that is, Problem 3. Then there is
  1. (1)
    a translation vector $z\in\mathrm{int}K$,
  2. (2)
    contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $L-z$ and $K-z$,
  3. (3)
    positive weights $\upalpha_{1},\dots,\upalpha_{m}$,
such that
(1.1) $P_{F}\left(\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}\right)P_{F}=P_{F};$
(1.2) $\sum\limits_{i\in[m]}\upalpha_{i}p_{i}=\mathbf{0};$
(1.3) $\sum\limits_{i\in[m]}\upalpha_{i}P_{F}v_{i}=\mathbf{0};$
(1.4) $\operatorname{tr}\left(\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}\right)=\sum\limits_{i\in[m]}\upalpha_{i}=d.$
As a corollary, we obtain the following.
Corollary 1.5 (Containment for John’s general problem with fixed axis).
With the assumptions of Theorem 1.4 and the translation vector $z\in\mathrm{int}K$ found therein, we have
$(K-z)\cap F\subset\left(L-z\right)\cap F\subset-d\cdot P_{F}(K-z).$
In the case $F=\mathbb{R}^{d}$, Theorem 1.4 and Corollary 1.5 were obtained in [GLMP04].
Our necessary condition of optimality for Problem 4 is the following.
Theorem 1.6 (Condition for John’s general problem with rotated axis).
With the notation of Theorem 1.4, assume that $K$ is a solution to John’s general problem with rotated axis for $K$, $L$ and the subspace $F$. Then there is a translation vector $z\in\mathrm{int}K$, and there are contact pairs of $L-z$ and $K-z$, and weights as in Theorem 1.4 such that in addition to identities (1.1) – (1.4), we also have
(1.5) $\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}=\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\otimes v_{i}.$
We will state our analogous results for ellipsoids of revolution, that is for Problems 1 and 2 in Section 5. Here we state other geometric properties of maximal volume ellipsoids of revolution.
We define the inradius $\operatorname{inrad}A$ of a compact, convex set $A$ as the radius of a largest radius ball inside the affine hull of $A$ contained in $A$.
Theorem 1.7 (Properties of solutions of John’s problem for ellipsoids with a fixed axis of revolution).
Let an origin-centered ellipsoid $E$ be a solution to Problem 1 for a convex body $K$ and a linear subspace $F$ in $\mathbb{R}^{d}$ of dimension $s\in\{0,1,\dots,d\}$. Then
  1. (1)
    the inradius of $K\cap{F}^{\bot}$ satisfies
    $\operatorname{inrad}\left(K\cap{F}^{\bot}\right)\leq\frac{d}{d-s}\cdot\mathrm{radius}(E\cap{F}^{\bot}),$
  2. (2)
    the following volume bound holds:
    $\left(\frac{\operatorname{vol}\nolimits_{s}\left(K\cap F\right)}{\operatorname{vol}\nolimits_{s}\left(E\cap F\right)}\right)^{\frac{1}{s}}\leq\frac{1}{(s+1)^{\frac{d-s}{2s(d+1)}}}\sqrt{\frac{d(d+1)}{s(s+1)}}\left(\frac{\operatorname{vol}\nolimits_{s}\Delta^{s}_{J}}{\operatorname{vol}\nolimits_{s}\mathbf{B}^{s}}\right)^{\frac{1}{s}},$
    where $\Delta_{J}^{s}$ stands for the $s$-dimensional regular simplex circumscribed around $\mathbf{B}^{s}.$
  3. (3)
    the following inclusion holds:
    $E\cap F\subset K\cap F\subset d\left(E\cap F\right).$
If additionally, $K$ is centrally symmetric, then
  1. (S1)
    the inradius of $K\cap{F}^{\bot}$ satisfies
    $\operatorname{inrad}\left(K\cap{F}^{\bot}\right)\leq\sqrt{\frac{d}{d-s}}\cdot\mathrm{radius}(E\cap{F}^{\bot}),$
  2. (S2)
    the following volume bound holds:
    $\left(\frac{\operatorname{vol}\nolimits_{s}\left(K\cap F\right)}{\operatorname{vol}\nolimits_{s}\left(E\cap F\right)}\right)^{\frac{1}{s}}\leq\sqrt{\frac{d}{s}}\left(\frac{\operatorname{vol}\nolimits_{s}\Box^{s}}{\operatorname{vol}\nolimits_{s}\mathbf{B}^{s}}\right)^{\frac{1}{s}},$
    where $\Box^{s}$ stands for the standard cube $[-1,1]^{s}.$
  3. (S3)
    the following inclusion holds:
    $E\cap F\subset K\cap F\subset\sqrt{d}\left(E\cap F\right).$
This result shows that the section $K\cap F$ exhibits the same properties as a section through the center of the John ellipsoid. There is hardly any difference between the proofs of (2), (3) (and (S2), (S3)) and the corresponding statements for John ellipsoids. Everything reduces to the algebraic conditions on the projections of the contact vectors onto $F$. The proof of (1) (and (S1)) however, is more intricate.
We prove Theorems 1.4 and 1.6 in Sections 2 and 3. Then, in Section 4, we consider the problem of finding a homothety ratio such that enlarging the inner body in some sort of John position by this ratio, it will contain the outer body. In particular, we prove (3) in Theorem 1.7, and also show that no similar containment holds in the whole space $\mathbb{R}^{d}$. In Section 3.4, we state a certain converse of Theorems 1.4 and 1.6, that is, that the equations in this theorem ensure that $K$ is of locally maximum volume among certain affine images of $K$ under some additional conditions.
In Section 5, we discuss the problem of finding largest volume ellipsoids of revolution in a convex body, and in particular, we prove Theorem 1.7. We discuss the uniqueness of the solution to these John type problems in Section 6, and finally, in Section 7 outline how some of the results change for the dual question (often referred to as Löwner’s ellipsoid), where the smallest volume ellipsoid containing a convex body is considered. In Appendix A, we show that (1) of Theorem 1.7 is tight. Appendix B contains the proof of a technical fact that we use in Section 5.
1.4. Notation
We denote by $\mathbf{0}$ the origin of a vector space, and sometimes in a subscript, we specify the space, eg. $\mathbf{0}_{\mathcal{L}}$ in a space $\mathcal{L}$ of matrices. We use $\left\langle u,v\right\rangle$ to denote the standard inner product of the vectors $u,v\in\mathbb{R}^{d}$, and $\mathbf{B}^{d}$ for the closed Euclidean unit ball in $\mathbb{R}^{d}$ centered at the origin. We denote the positive hull of a set $S$ in $\mathbb{R}^{n}$ by $\operatorname{pos}\left(S\right)=\left\{\lambda x:\;\lambda\geq 0,x\in S\right\}.$ The polar of $S$ is defined as $S^{\circ}=\left\{p\in\mathbb{R}^{n}:\;\left\langle x,p\right\rangle\leq 1\,\text{ for all }x\in S\right\}$. Note that if $C$ is a positive cone (that is, the positive hull of some set), then $C^{\circ}=\left\{p\in\mathbb{R}^{n}:\;\left\langle x,p\right\rangle\leq 0\,\text{ for all }x\in C\right\}$. The group of orthogonal transformations on $\mathbb{R}^{d}$ is $O(d)$.
Recall that the polar decomposition of a square matrix $A\in\mathbb{R}^{d\times d}$ is a factorization of $A$ in the form $A=RM,$ where $R$ is an orthogonal matrix and $M$ is positive semi-definite. Moreover, if $A$ is invertible, then this representation is unique ($M=\sqrt{{A}^{\top}A}$ and $U=A{M}^{-1}$), where ${A}^{\top}$ denotes the transpose of $A$.
2. First-order necessary conditions
In this section, we reformulate Theorems 1.4 and 1.6 in terms of the (non-)existence of a certain curve in a vector space. The difference between the proofs of the two theorems lies in the choice of the vector space in which the curve (non-)exists. This condition is then translated into a statement about the separation of a point and a convex cone in the corresponding space. This is a standard technique in proving John-type theorems; cf. [Bal97].
First, we recall basic facts describing convex sets locally, and next, we introduce the vector space which parametrizes local affine transformations. In the final part of this section, we rephrase the existence of a curve in this vector space along which an ‘improved’ position of $K$ in $L$ can be found in terms of separation of cones.
2.1. The normal and the tangent cone
Definition 2.1.
Let $C$ be a convex body in $\mathbb{R}^{n}$ and $a\in\operatorname{bd}C$ be a boundary point. The normal cone $N_{K}\!\left(a\right)$ is the set consisting of the origin and all the outward normal vectors to $C$ at $a$. That is,
$N_{C}\!\left(a\right)=\left\{p:\;\left\langle p,z-a\right\rangle\leq 0\quad\forall z\in C\right\}.$
The tangent cone $\mathrm{T}_{C}\!\left(a\right)$ at $a$ is the closure of the cone $\mathrm{Pos}\ \!{\left(C-a\right)}$.
The following facts are well-known, and follow from standard separation properties of convex sets:
Lemma 2.2.
Consider a convex set $C$ in $\mathbb{R}^{d}$ and a point $a\in\operatorname{cl}\left(C\right)$. Then
${\mathrm{T}_{C}\!\left(a\right)}={\left(N_{C}\!\left(a\right)\right)}^{\circ}\quad\text{and}\quad{\left(\mathrm{T}_{C}\!\left(a\right)\right)}^{\circ}={N_{C}\!\left(a\right)}.$
In particular, $h\in\mathrm{int}\mathrm{T}_{C}\!\left(a\right)$ if and only if
$\left\langle h,p\right\rangle<0$
for all non-zero $p\in N_{C}\!\left(a\right)$.
Lemma 2.3.
Let $C$ be a convex body in $\mathbb{R}^{n}$, and $a\in C.$ Let $\gamma_{t}$ where $t\in[0,1]$ be a curve in $\mathbb{R}^{n}$ with $\gamma_{0}=a$ and
$\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}\gamma_{t}=h.$
Consider the following assertions:
  1. (i)
    $h\in\mathrm{int}\mathrm{T}_{C}\!\left(a\right);$
  2. (ii)
    $\gamma_{t}\in C$ for all $t\in[0,\tau]$ for some $\tau>0;$
  3. (iii)
    $h\in\mathrm{T}_{C}\!\left(a\right)$.
Then (i) implies (ii), and (ii) implies (iii).
2.2. Operator spaces
We use $\mathcal{M}^{d}$ and $\mathcal{M}_{\mathrm{Sym}}^{d}$ to denote the space of operators and the space of self-adjoint operators from $\mathbb{R}^{d}$ to $\mathbb{R}^{d}$, respectively. Define
=$\mathcal{R}=\left\{(M,S,g):\;M,S\in\mathcal{M}^{d},\ g\in\mathbb{R}^{d}\right\}.$
We endow $\mathcal{R}$ with an inner product as follows:
$\left\langle(M_{1},S_{1},g_{1}),(M_{2},S_{2},g_{2})\right\rangle_{\mathcal{R}}=\operatorname{tr}\left(M_{1}{M_{2}}^{\top}\right)+\operatorname{tr}\left(S_{1}{S_{2}}^{\top}\right)+\left\langle g_{1},g_{2}\right\rangle.$
2.3. A curve in $\mathcal{R}$
First, we describe certain one parameter families of affine transformations of $\mathbb{R}^{d}$. More specifically, we consider a differentiable path $\gamma:[0,t_{0})\longrightarrow\mathcal{R}$ for some $t_{0}>0$ defined by
(2.1) $\gamma_{t}=(\mathrm{Id}_{\mathbb{R}^{d}}+tH,R_{t},h_{t}),$
where $h_{0}=\mathbf{0}$, $R_{t}\in O(d)$ and $R_{0}=\mathrm{Id}_{\mathbb{R}^{d}}$.
It is well-known [Hal13, Proposition 3.24.] that by setting
(2.2) $W=\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}R_{t},$
we have that $W$ is a skew-symmetric matrix, that is, $W=-{W}^{\top}$, and conversely, for every skew symmetric matrix $W$, there is a path $R_{t}$ in the orthogonal group satisfying (2.2).
Denoting $\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}h_{t}$ by $z$, we have
(2.3) $\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}\gamma_{t}=(H,W,z).$
For any $t$, we will regard $\gamma_{t}$ as a representation of the affine transformation
(2.4) $A_{t}=R_{t}(\mathrm{Id}_{\mathbb{R}^{d}}+tH)+h_{t}$
on $\mathbb{R}^{d}$.
Throughout the section, $R_{t},A_{t},h_{t},H,W,z$ denote the quantities defined so far, all derived from a differentiable path $\gamma_{t}$ described in (2.1).
Let $K$ and $L$ be convex bodies in $\mathbb{R}^{d}$ with $K\subset L$. We are interested in two questions: first, how does the volume of $K$ change when we replace $K$ by $A_{t}(K)$? Second, under what conditions is $A_{t}(K)$ contained in $L$?
The first question is easy. We differentiate the determinant of the linear operator part of $A_{t}$.
(2.5) $\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}\det(R_{t}(\mathrm{Id}_{\mathbb{R}^{d}}+tH))=\operatorname{tr}H=$
$\left\langle\left(\mathrm{Id}_{\mathbb{R}^{d}},\mathbf{0}_{\mathcal{M}^{d}},\mathbf{0}\right),\left(H,W,z\right)\right\rangle_{\mathcal{R}}=\left\langle\left(\mathrm{Id}_{\mathbb{R}^{d}},\mathbf{0}_{\mathcal{M}^{d}},\mathbf{0}\right),\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}\gamma_{t}\right\rangle_{\mathcal{R}}.$
Setting
$\mathcal{V}_{\mathcal{R}}=\mathrm{Pos}\ \!{\left(\mathrm{Id}_{\mathbb{R}^{d}},\mathbf{0}_{\mathcal{M}^{d}},\mathbf{0}\right)},$
which is a ray in $\mathcal{R}$, we conclude the following.
Lemma 2.4.
If
$\left(H,W,z\right)\in-\mathrm{int}\left({\left(\mathcal{V}_{\mathcal{R}}\right)}^{\circ}\right),$
then the volume of $A_{t}(K)$ is strictly increasing for sufficiently small $t>0$.
Second, Lemmas 2.2 and 2.3 provide a nice tool for characterization of local inclusions along curves.
Lemma 2.5.
Let $L$ be a convex body in $\mathbb{R}^{d}$ containing the origin in its interior, and let $v$ be a boundary point of $L$. Define the perturbation $v_{t}$ of $v$ by
$v_{t}=A_{t}(v).$
Consider the following statements:
  1. (1)
    $\left\langle(H,W,z),(v\otimes p,v\otimes p,p)\right\rangle_{\mathcal{R}}<0;$
    for all $p\in{L}^{\circ}$ with $\left\langle p,v\right\rangle=1$.
  2. (2)
    $v_{t}\in L$ for all $t\in(0,\tau]$ and some $\tau>0$.
  3. (3)
    $\left\langle(H,W,z),(v\otimes p,v\otimes p,p)\right\rangle_{\mathcal{R}}\leq 0;$
    for all $p\in{L}^{\circ}$ with $\left\langle p,v\right\rangle=1$.
Then (1) implies (2), and (2) implies (3).
Proof.
By direct computation,
$\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}v_{t}=\left(H+W\right)v+z.$
Thus, for any $p\in\mathbb{R}^{d}$, we have
$\left\langle\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}v,p\right\rangle=\left\langle\left(H+W\right)v+z,p\right\rangle=\left\langle Hv,p\right\rangle+\left\langle Wv,p\right\rangle+\left\langle z,p\right\rangle=$
$\left\langle(H,W,z),(v\otimes p,v\otimes p,p)\right\rangle_{\mathcal{R}}.$
The result follows from the descriptions of the tangent cone in Lemma 2.2 and Lemma 2.3. ∎
We denote the set of contact points of $K$ and $L$ by
$\mathrm{Cont}_{\mathbb{R}^{d}}=\operatorname{bd}K\cap\operatorname{bd}L,$
and, from this point onward, assume that $K$ contains the origin in its interior.
Corollary 2.6.
If
(2.6) $\left\langle(H,W,z),(v\otimes p,v\otimes p,p)\right\rangle_{\mathcal{R}}<0$
for all $v\in\mathrm{Cont}_{\mathbb{R}^{d}}$ and for all $p\in{L}^{\circ}$ with $\left\langle p,v\right\rangle=1$, then for a sufficiently small $\varepsilon>0$, the set $A_{t}K+h_{t}$ is contained in $\mathrm{int}L$ for all $t\in(0,\varepsilon)$.
Proof.
By Lemma 2.5, for any $v\in\mathrm{Cont}_{\mathbb{R}^{d}}$, $A_{t}(v)$ is in $\mathrm{int}K$ for all $t\in(0,\varepsilon_{v})$ with some $\varepsilon_{v}>0$ depending on $v$. A standard compactness argument completes the proof. ∎
Define the set of contact operators by
$\mathrm{Cont}_{\mathcal{R}}=\left\{\left(v\otimes p,v\otimes p,p\right):\;v\in\mathrm{Cont}_{\mathbb{R}^{d}},p\in{L}^{\circ},\left\langle v,p\right\rangle=1\right\},$
and the contact operator cone by
$\mathcal{C}_{\mathcal{R}}=\mathrm{Pos}\ \!{\mathrm{Cont}_{\mathcal{R}}}.$
With this notation, Corollary 2.6 can be restated as follows.
Corollary 2.7.
If
(2.7) $(H,W,z)\in\mathrm{int}\left({\mathcal{C}_{\mathcal{R}}}^{\circ}\right)$
then for a sufficiently small $\varepsilon>0$, the set $A_{t}K+h_{t}$ is contained in $\mathrm{int}L$ for all $t\in(0,\varepsilon)$.
Definition 2.8.
Let $K\subseteq L$ be convex bodies in $\mathbb{R}^{d}$. We say that $K$ is of locally maximum volume among elements of $\mathcal{R}_{F}\!\left(K\right)$ (resp., $\mathcal{A}_{F}\!\left(K\right)$) contained in $L$, if there is no $t_{0}>0$ and differentiable curve $A_{t}:[0,t_{0})\longrightarrow\mathcal{R}_{F}$ (resp., $A_{t}:[0,t_{0})\longrightarrow\mathcal{A}_{F}$) such that $A_{0}=\mathrm{Id}_{\mathbb{R}^{d}}$, $\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}\operatorname{vol}\nolimits_{d}A_{t}(K)>0$ and $A_{t}(K)\subset\mathrm{int}L$ for all $t\in(0,t_{0})$.
For some fixed subspace $F$ of $\mathbb{R}^{d}$, we set
$\mathcal{D}=\bigg{\{}(H,W,z)\in\mathcal{R}:\;H\ \text{ is an }F\text{-operator, }W=-{W}^{\top},\ z\in\mathbb{R}^{d}\bigg{\}}.$
Lemma 2.9 (Local optimality in $\mathcal{R}_{F}\!\left(K\right)$).
If $\mathbf{0}\in\mathrm{int}K$ and $K$ is of locally maximum volume among elements of $\mathcal{R}_{F}\!\left(K\right)$ contained in $L$, then
$\left(\mathcal{V}_{\mathcal{R}}+{\mathcal{D}}^{\bot}\right)\cap\mathcal{C}_{\mathcal{R}}\neq\{\mathbf{0}_{\mathcal{R}}\}.$
Proof.
By (2.2) and (2.3), we have that if $(H,W,z)$ is in $\mathcal{D}$, then there is a curve $\gamma_{t}$ in $\mathcal{R}$ with $\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}\gamma_{t}=(H,W,z)$, and $\gamma_{0}=(\mathrm{Id}_{\mathbb{R}^{d}},\mathbf{0}_{\mathcal{M}^{d}},\mathbf{0})$ such that $A_{t}$ defined by (2.4) is a perturbation of $K$ with $A_{t}(K)\in\mathcal{R}_{F}\!\left(K\right).$
Combining this observation with Lemma 2.4 and Corollary 2.7, we obtain the following.
If $K\subseteq L$ are convex bodies in $\mathbb{R}^{d}$ with $\mathbf{0}\in\mathrm{int}L$, and
$\mathcal{D}\cap\mathrm{int}\left({\mathcal{C}_{\mathcal{R}}}^{\circ}\right)\cap\mathrm{int}\left(-{\left(\mathcal{V}_{\mathcal{R}}\right)}^{\circ}\right)\neq\{\mathbf{0}_{\mathcal{R}}\},$
then there is a curve $\gamma$ showing that $K$ is not of locally maximum volume among elements of $\mathcal{R}_{F}\!\left(K\right)$ contained in $L$.
Equivalently, if $K$ is of locally maximum volume among elements of $\mathcal{R}_{F}\!\left(K\right)$ contained in $L$, then
(2.8) $\mathcal{D}\cap\mathrm{int}\left(-{\left(\mathcal{V}_{\mathcal{R}}\right)}^{\circ}\right)\cap\mathrm{int}\left({\mathcal{C}_{\mathcal{R}}}^{\circ}\right)=\{\mathbf{0}_{\mathcal{R}}\}.$
Notice that $\mathcal{D}$ is a linear subspace in $\mathcal{R}$, and the other two sets on the left-hand side are open cones in $\mathcal{R}$. By basic separation properties of cones in Euclidean space, the proof of Lemma 2.9 is complete. ∎
What changes if we are not allowed to use a rotation in the setting of Theorem 1.4? In this case, the curve $A_{t}$ is in the subspace $\mathcal{A}_{F}$, or, equivalently, $\gamma_{t}$ is such that $R_{t}=\mathrm{Id}_{\mathbb{R}^{d}}$ for all $t$, which is equivalent to the condition that $\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}\gamma_{t}$ takes all values in
$\mathcal{D}_{0}=\bigg{\{}(H,\mathbf{0}_{\mathcal{M}^{d}},z)\in\mathcal{R}:\;H\ \text{ is an }F\text{-operator, }\ z\in\mathbb{R}^{d}\bigg{\}}.$
Thus, (2.8) changes to the following.
If $K$ is of locally maximum volume among elements of $\mathcal{A}_{F}\!\left(K\right)$ contained in $L$, then
$\mathcal{D}_{0}\cap\mathrm{int}\left(-{\left(\mathcal{V}_{\mathcal{R}}\right)}^{\circ}\right)\cap\mathrm{int}\left({\mathcal{C}_{\mathcal{R}}}^{\circ}\right)=\{\mathbf{0}_{\mathcal{R}}\}.$
This yields the following.
Lemma 2.10 (Local optimality in $\mathcal{A}_{F}\!\left(K\right)$).
If $\mathbf{0}\in\mathrm{int}K$ and $K$ is of locally maximum volume among elements of $\mathcal{A}_{F}\!\left(K\right)$ contained in $L$, then
$\left(\mathcal{V}_{\mathcal{R}}+{\mathcal{D}_{0}}^{\bot}\right)\cap\mathcal{C}_{\mathcal{R}}\neq\{\mathbf{0}_{\mathcal{R}}\}.$
In the next two sections, we prove Theorems 1.6 and 1.4, using Lemmas 2.9 and 2.10, respectively.
3. Proof of Theorems 1.6 and 1.4
We fix a non-trivial subspace $F$ of $\mathbb{R}^{d}.$ We use $P_{F}$ to denote the orthogonal projection onto $F$. Set
$\mathcal{Q}_{F}=\left\{Q\in\operatorname{Lin}\left(\mathbb{R}^{d}\right):\;P_{F}QP_{F}=\mathbf{0}_{F},\operatorname{rank}\left(P_{{F}^{\bot}}QP_{{F}^{\bot}}\right)=d-\dim{F},\operatorname{tr}\left(P_{{F}^{\bot}}QP_{{F}^{\bot}}\right)=0\right\},$
where $\mathbf{0}_{F}$ is the zero operator on $F$. In coordinates, if we consider an orthonormal basis of $F$ and extend it to an orthonormal basis of $\mathbb{R}^{d}$, then the matrix of $Q\in\mathcal{Q}_{F}$ in this basis is of the form
$Q=\begin{pmatrix}\mathbf{0}_{s\times s}&Q_{1}\\ Q_{2}&Q_{3}\end{pmatrix},$
where $\mathbf{0}_{s\times s}$ is the $s\times s$ zero matrix, $Q_{1}$ and ${Q_{2}}^{\top}$ are $s\times(d-s)$ matrices, and $Q_{3}$ is a square matrix of rank $d-s$ of zero trace.
Lemma 3.1.
The following identities hold:
${\mathcal{D}}^{\bot}=\mathcal{Q}_{F}\times\mathcal{M}_{\mathrm{Sym}}^{d}\times\{\mathbf{0}\},\quad\text{and}\quad{\mathcal{D}_{0}}^{\bot}=\mathcal{Q}_{F}\times\mathcal{M}^{d}\times\{\mathbf{0}\}.$
Proof.
Let $\left(M,S,g\right)\in\mathcal{R}$. If $\left(M,S,g\right)\in{\mathcal{D}}^{\bot}$, then
$\left\langle\left(M,S,g\right),\left(\mathbf{0}_{\mathcal{M}^{d}},\mathbf{0}_{\mathcal{M}^{d}},z\right)\right\rangle_{\mathcal{R}}=\left\langle g,z\right\rangle=0$
for all $z\in\mathbb{R}^{d}$. Hence, $g=\mathbf{0}$.
The orthogonal complement of the space of skew-symmetric matrices in $\mathcal{M}^{d}$ is $\mathcal{M}_{\mathrm{Sym}}^{d}$. Thus, $S\in\mathcal{M}_{\mathrm{Sym}}^{d}$.
Finally, let $M$ and an $F$-matrix $H$ be of the form:
$H=\begin{pmatrix}{H_{1}}&\mathbf{0}_{s\times(d-s)}\\ \mathbf{0}_{(d-s)\times s}&\mu\mathrm{Id}_{d-s}\end{pmatrix}\quad\text{and}\quad M=\begin{pmatrix}{M_{1}}&{M_{2}}\\ {M_{3}}&{M_{4}}\end{pmatrix}$
for some non-zero constant $\mu$, where $H_{1}$ and $M_{1}$ are $s\times s$ matrices, $M_{4}$ is a $(d-s)\times(d-s)$ matrix, and $M_{2}$ and ${M_{3}}^{\top}$ are $s\times(d-s)$ matrices.
Then
$\left\langle\left(M,S,\mathbf{0}\right),\left(H,\mathbf{0}_{\mathcal{M}^{d}},z\right)\right\rangle_{\mathcal{R}}=\operatorname{tr}\left(M{H}^{\top}\right)=\operatorname{tr}\left(M_{1}{H_{1}}^{\top}\right)+\mu\operatorname{tr}M_{4}.$
As $H_{1}$ and $\mu$ can be chosen arbitrarily,
$\left\langle\left(M,S,\mathbf{0}\right),\left(H,\mathbf{0}_{\mathcal{M}^{d}},z\right)\right\rangle_{\mathcal{R}}=0$
if and only if $M_{1}=\mathbf{0}_{s\times s}$ and $\operatorname{tr}M_{4}=0$, while there are no constraints on $M_{2}$ and $M_{3}$. This is equivalent to $M\in\mathcal{Q}_{F}$.
Thus,
${\mathcal{D}}^{\bot}=\mathcal{Q}_{F}\times\mathcal{M}_{\mathrm{Sym}}^{d}\times\{\mathbf{0}\}.$
The second statement of the lemma follows analogously. ∎
3.1. Slight variants of Theorem 1.6 and 1.4
In this subsection, we prove variants of these two theorems without a translation $z$, and as a price, we do not obtain (1.3).
The following result differs slightly from Theorem 1.6.
Theorem 3.2.
Let $K$, $L$ and $F$ satisfy the conditions of Theorem 1.6, and assume that the origin is in the interior of $K$, and that $K$ is of locally maximum volume among elements of $\mathcal{R}_{F}\!\left(K\right)$ contained in $L$.
Then we have contact points of $K$ and $L$ and weights satisfying (1.1), (1.2), (1.4), and (1.5).
Proof of Theorem 3.2.
Lemmas 2.9 and 3.1 yield that
$\left(\mathrm{Pos}\ \!{\{}(\mathrm{Id}_{\mathbb{R}^{d}},\mathbf{0}_{\mathcal{M}^{d}},\mathbf{0})\}+\mathcal{Q}_{F}\times\mathcal{M}_{\mathrm{Sym}}^{d}\times\{\mathbf{0}\}\right)\cap\mathcal{C}_{\mathcal{R}}\neq\{\mathbf{0}_{\mathcal{R}}\}.$
Thus, there are matrices $Q\in\mathcal{Q}_{F}$ and $S\in\mathcal{M}_{\mathrm{Sym}}^{d}$ such that $\left(\mathrm{Id}_{\mathbb{R}^{d}}+Q,S,\mathbf{0}\right)\in\mathcal{C}_{\mathcal{R}}$. That is, there are contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $K$ and $L,$ and weights $\upalpha_{1},\dots,\upalpha_{m}>0$ such that $\sum\limits_{i\in[m]}\upalpha_{i}p_{i}=\mathbf{0}$ and the matrix $M=\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}$ satisfies the conditions
  • $P_{F}MP_{F}=P_{F}(\mathrm{Id}_{\mathbb{R}^{d}}+Q)P_{F}=P_{F}\mathrm{Id}_{\mathbb{R}^{d}}P_{F}=P_{F}$,
  • $M={M}^{\top}$,
  • $\operatorname{tr}M=\operatorname{tr}\left(\mathrm{Id}_{\mathbb{R}^{d}}+Q\right)=\operatorname{tr}\mathrm{Id}_{\mathbb{R}^{d}}=d.$
The proof of Theorem 3.2 is complete. ∎
The following result differs slightly from Theorem 1.4.
Theorem 3.3.
Let $K$, $L$ and $F$ satisfy the conditions of Theorem 1.4, and assume that the origin is in the interior of $L$, and that $K$ is of locally maximum volume among elements of $\mathcal{A}_{F}\!\left(K\right)$ contained in $L$.
Then we have contact points of $K$ and $L$ and weights satisfying (1.1), (1.2), and (1.4).
The proof is essentially identical to the previous one.
Proof of Theorem 3.3.
Lemmas 2.10 and 3.1 yield that
$\left(\mathrm{Pos}\ \!{\{}(\mathrm{Id}_{\mathbb{R}^{d}},\mathbf{0}_{\mathcal{M}^{d}},\mathbf{0})\}+\mathcal{Q}_{F}\times\mathcal{M}^{d}\times\{\mathbf{0}\}\right)\cap\mathcal{C}_{\mathcal{R}}\neq\{\mathbf{0}_{\mathcal{R}}\}.$
Thus, there are matrices $Q\in\mathcal{Q}_{F}$ and $S\in\mathcal{M}^{d}$ such that $\left(\mathrm{Id}_{\mathbb{R}^{d}}+Q,S,\mathbf{0}\right)\in\mathcal{C}_{\mathcal{R}}$. That is, there are contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $K$ and $L,$ and weights $\upalpha_{1},\dots,\upalpha_{m}>0$ such that $\sum\limits_{i\in[m]}\upalpha_{i}p_{i}=\mathbf{0}$ and the matrix $M=\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}$ satisfies the conditions
  • $P_{F}MP_{F}=P_{F}(\mathrm{Id}_{\mathbb{R}^{d}}+Q)P_{F}=P_{F}\mathrm{Id}_{\mathbb{R}^{d}}P_{F}=P_{F}$, and
  • $\operatorname{tr}M=\operatorname{tr}\left(\mathrm{Id}_{\mathbb{R}^{d}}+Q\right)=\operatorname{tr}\mathrm{Id}_{\mathbb{R}^{d}}=d.$
The proof of Theorem 3.3 is complete. ∎
3.2. Rotations
This subsection is a small detour in which we clarify the geometry hidden behind the symmetry condition (1.5). The proofs of our other results do not rely on it.
Theorem 3.4.
Let $K$ and $L$ be convex bodies in $\mathbb{R}^{d}$ such that the origin is in the interior of $L\cap K,$ and there are no
  1. (1)
    contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $L$ and $K;$
  2. (2)
    positive weights $\upalpha_{1},\dots,\upalpha_{m}$
satisfying identity (1.5). Then for some $\varepsilon>0$, there is a differentiable path $R_{t},$ $t\in[0,\varepsilon]$ of orthogonal transofrmations such that $R_{t}K\subset\mathrm{int}L$ for all $t\in(0,\varepsilon].$
Proof.
There is nothing to prove if the set of contact pairs is empty. We assume that the cone $\operatorname{Cone}_{\mathrm{op}}=\mathrm{Pos}\ \!{\left\{v\otimes p:\;(v,p)\text{ is a contact pair of }K\text{ and }L\right\}}$ is non-empty.
Then the negation of identity (1.5) is equivalent to
$\mathcal{M}_{\mathrm{Sym}}^{d}\cap\operatorname{Cone}_{\mathrm{op}}=0.$
Thus, there is non-zero $W$ separating the two convex cones. However, $\mathcal{M}_{\mathrm{Sym}}^{d}$ is a subspace, hence $W\in{\left(\mathcal{M}_{\mathrm{Sym}}^{d}\right)}^{\bot},$ that is, $W$ is skew symmetric.
Set $R_{t}=e^{tW},$ $H=0,$ and ${U}^{\top}=\mathrm{Id}_{d}$ in (2.1). Clearly, $R_{t},t\in[0,1]$ is a differentiable path of orthogonal matrices in $\mathcal{M}^{d}$ satisfying (2.2). The result follows from Corollary 2.7. ∎
3.3. Search for a good center
We show how $z$ can be found to obtain (1.3) in Theorems 1.6 and 1.4.
Lemma 3.5.
Let $K\subseteq L$ be convex bodies in $\mathbb{R}^{d}$ such that the origin is in the interior of $K$, and there are
  1. (1)
    an $s$-dimensional subspace $F;$
  2. (2)
    contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $L$ and $K;$
  3. (3)
    positive weights $\upalpha_{1},\dots,\upalpha_{m}$
satisfying identities (1.1) – (1.4). Then there exists a point $z$ in the interior of $L\cap K$, contact pairs $\{(v_{1}^{\prime},p_{1}^{\prime}),\dots,(v_{m}^{\prime},p_{m}^{\prime})\}$ of $K-z$ and $L-z$, and weights $c_{1},\dots,c_{m}>0$ such that
${\sum_{i\in[m]}c_{i}v_{i}^{\prime}\otimes p_{i}^{\prime}}=\sum_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i},\quad\text{and}\quad\sum_{i\in[m]}c_{i}P_{F}v_{i}^{\prime}=\sum_{i\in[m]}c_{i}p_{i}^{\prime}=\mathbf{0}.$
Proof.
This is a purely linear algebraic observation.
We aim at finding a point $z$ in the interior of $K$ such that when we shift the origin to $z$, the shifted contact pairs $(v_{i}-z,p_{i}^{\prime})$ satisfy the required identities for $K-z$ and $L-z$.
First, observe how a contact pair transforms under translation. A pair $(v,p)$ is a contact pair of $K$ and $L$ if and only if $v$ belongs to the intersection of the boundaries of $K$ and $L$, and the half-space
$H_{p}^{-}=\left\{x\in\mathbb{R}^{d}:\;\left\langle x,p\right\rangle\leq\left\langle v,p\right\rangle=1\right\}$
contains $L$.
Under translation by $z$, the half-space becomes
$H_{p}^{-}-z=\left\{x-z\in\mathbb{R}^{d}:\;\left\langle x,p\right\rangle-\left\langle z,p\right\rangle\leq 1-\left\langle z,p\right\rangle\right\}=\left\{y\in\mathbb{R}^{d}:\;\left\langle y,\frac{p}{1-\left\langle z,p\right\rangle}\right\rangle\leq 1\right\}.$
Thus, provided $\left\langle z,p\right\rangle<1$, the pair $(v-z,p^{\prime})$, where $p^{\prime}=\dfrac{p}{1-\left\langle z,p\right\rangle}$, is a contact pair of $K-z$ and $L-z$.
For every $i\in[m]$, define
$v_{i}(z)=v_{i}-z,\quad\text{and}\quad p_{i}(z)=\frac{p_{i}}{1-\left\langle z,p_{i}\right\rangle}.$
To satisfy the identity $\sum\limits_{i\in[m]}c_{i}(z)p_{i}(z)=\mathbf{0}$, we set
$c_{i}(z)=\upalpha_{i}\left(1-\left\langle z,p_{i}\right\rangle\right).$
Then,
$c_{i}(z)p_{i}(z)=\upalpha_{i}\left(1-\left\langle z,p_{i}\right\rangle\right)\frac{p_{i}}{1-\left\langle z,p_{i}\right\rangle}=\upalpha_{i}p_{i},$
which implies
$\sum_{i\in[m]}c_{i}(z)p_{i}(z)=\sum_{i\in[m]}\upalpha_{i}p_{i}=\mathbf{0}.$
Set $M=\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}.$ Next, we verify the identity $\sum\limits_{i\in[m]}c_{i}(z)v_{i}(z)\otimes p_{i}(z)=M$. Indeed,
$\sum_{i\in[m]}c_{i}(z)v_{i}(z)\otimes p_{i}(z)=\sum_{i\in[m]}\upalpha_{i}\left(1-\left\langle z,p_{i}\right\rangle\right)(v_{i}-z)\otimes\frac{p_{i}}{1-\left\langle z,p_{i}\right\rangle}=\sum_{i\in[m]}\upalpha_{i}(v_{i}-z)\otimes p_{i}.$
Simplifying, we get
[…]
Finally, we need to satisfy $P_{F}\left(\sum\limits_{i\in[m]}c_{i}(z)v_{i}(z)\right)=\mathbf{0}$. Computing, we have
$\sum_{i\in[m]}c_{i}(z)P_{F}v_{i}(z)$ $=\sum_{i\in[m]}\upalpha_{i}\left(1-\left\langle z,p_{i}\right\rangle\right)P_{F}(v_{i}-z)$
$=\sum_{i\in[m]}\upalpha_{i}P_{F}v_{i}-\sum_{i\in[m]}\upalpha_{i}\left\langle z,p_{i}\right\rangle P_{F}v_{i}-P_{F}z\sum_{i\in[m]}\upalpha_{i}+P_{F}z\sum_{i\in[m]}\upalpha_{i}\left\langle z,p_{i}\right\rangle.$
Using the identities $\sum\limits_{i\in[m]}\upalpha_{i}p_{i}=\mathbf{0}$ and $P_{F}\left(\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}\right)P_{F}=P_{F}$, we can simplify:
$\sum\limits_{i\in[m]}c_{i}(z)P_{F}v_{i}(z)=\sum\limits_{i\in[m]}\upalpha_{i}P_{F}v_{i}-P_{F}\left(\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}\right)P_{F}z-P_{F}z\sum\limits_{i\in[m]}\upalpha_{i}+P_{F}z\left\langle z,\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\right\rangle$
$=\sum\limits_{i\in[m]}\upalpha_{i}P_{F}v_{i}-P_{F}z-P_{F}z\left(\sum\limits_{i\in[m]}\upalpha_{i}\right)+P_{F}z\cdot 0=\left(\sum\limits_{i\in[m]}\upalpha_{i}P_{F}v_{i}\right)-(d+1)P_{F}z.$
Thus, setting
$z=\frac{\sum\limits_{i\in[m]}\upalpha_{i}v_{i}}{d+1}$
ensures that $\sum\limits_{i\in[m]}c_{i}(z)P_{F}v_{i}(z)=\mathbf{0}$, as desired. Since $\sum\limits_{i\in[m]}\upalpha_{i}=d$, and $\mathbf{0}\in\mathrm{int}K$, we have $z\in\mathrm{int}K$. ∎
Proof of Theorem 1.6.
Combining Theorem 3.2 with Lemma 3.5 yields our theorem. ∎
Proof of Theorem 1.4.
Combining Theorem 3.3 with Lemma 3.5 yields our theorem. ∎
3.4. The condition is sufficient for local extremality – in some cases
According to Theorem 3.3, the local maximality of the volume of $K$ implies certain relations on contact pairs of $K$ and $L$.
The converse is not true in general. If we consider a regular simplex centered at the origin as $K$, and $L=-d\cdot K$, then there are contact pairs satisfying equations (1.1)–(1.5), and yet, clearly, $K$ is not a locally largest volume $F$-position of $K$ in $L$, for $F=\mathbb{R}^{d}$.
However, as in the classical setting, the log-concavity of the determinant (Minkowski’s determinant inequality) yields the sufficiency of our conditions for positions $AK+a$ with positive definite $A$. Denote by
$\mathcal{A}_{F}^{+}(S)=\left\{AS+z:\;A\text{ is a positive definite }F\text{-operator, }z\in\mathbb{R}^{d}\right\},$
the positive $F$-positions of $S$.
Theorem 3.6.
Let $K$ and $L$ be convex bodies in $\mathbb{R}^{d}$ and $F$ a linear subspace in $\mathbb{R}^{d}$. Assume that $K\subseteq L$ and there are contact points of $K$ and $L$ and weights satisfying (1.1), (1.2) and (1.4). Then $K$ is of locally maximum volume among elements of $\mathcal{A}_{F}^{+}(K)$ contained in $L$.
4. Containment
Let $K_{\text{in}}$ and $K_{\text{out}}$ be convex bodies in $\mathbb{R}^{d}$ such that $K_{\text{in}}\subset K_{\text{out}}$ and $K_{\text{in}}$ contains the origin in its interior.
Assume there are contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $K_{\text{in}}$ and $K_{\text{out}},$ and weights $\upalpha_{1},\dots,\upalpha_{m}$ such that identities (1.1) – (1.4) hold.
Denote $L_{\text{in}}=\operatorname{conv}\left\{v_{1},\dots,v_{m}\right\}$ and $L_{\text{out}}=\bigcap\limits_{i\in[m]}\left\{x\in\mathbb{R}^{d}:\;\left\langle p_{i},x\right\rangle\leq 1\right\}.$
We call $L_{\text{in}}$ and $L_{\text{out}}$ the inner and outer contact polytope corresponding to the contact pairs $(v_{i},p_{i})$. Clearly,
$L_{\text{in}}\subset K_{\text{in}}\subset K_{\text{out}}\subset L_{\text{out}}.$
Lemma 4.1.
Under the above notation, we have
$L_{\mathrm{out}}\cap F\subset-d\cdot P_{F}L_{\mathrm{in}}.$
Proof.
Let $x$ be a point in $L_{\text{out}}\cap F$, then $\left\langle p_{i},x\right\rangle\leq 1$ for all $i\in[m].$
[…]
$\sum\limits_{i\in[m]}\upalpha_{i}(1-\left\langle p_{i},x\right\rangle)P_{F}v_{i}\in\sum\limits_{i\in[m]}\upalpha_{i}\cdot P_{F}L_{\text{in}}\cdot(1-\left\langle p_{i},x\right\rangle)\stackrel{{\scriptstyle\left\langle p_{i},x\right\rangle\leq 1}}{{=}}$
[…]
Clearly, this lemma yields Corollary 1.5 as well as (3) in Theorem 1.7.
4.1. No containment in the whole space
Contrasting Corollary 1.5, we show that there is no meaningful inclusion in ${F}^{\bot}$. Surprisingly, the example is provided by ellipsoids.
Lemma 4.2.
Fix $20\leq s\leq d-$ 2 and a positive number $\lambda$. Let $E$ be an origin-centered ellipsoid in $\mathbb{R}^{d}$ whose principal semi-axes have lengths $\lambda,\ldots,\lambda^{d}$. Let $E^{\prime}\in\mathcal{A}_{F}\!\left(\mathbf{B}^{d}\right)$ be the solution to Problem 2 for some fixed $s$-dimensional subspace $F$. Then
$\frac{\operatorname{vol}\nolimits_{d}E^{\prime}}{\operatorname{vol}\nolimits_{d}E}\leq\frac{1}{\lambda}\quad\text{and}\quad\forall x\in\mathbb{R}^{d},\quad E\cap{F}^{\bot}\not\subset x+\lambda\cdot\mathrm{int}P_{{F}^{\bot}}E^{\prime}.$
Proof.
We need the following observation:
Lemma 4.3.
Let $E_{1}\subseteq E_{2}\subset\mathbb{R}^{d}$ be two ellipsoids. Let the principal semi-axes of $E_{1}$ be $\lambda_{1}\leq\ldots\leq\lambda_{d}$, and those of $E_{2}$ be $\mu_{1}\leq\ldots\leq\mu_{d}$. Then $\lambda_{i}\leq\mu_{i}$ for all $i\in[d]$.
Proof.
By symmetry, we may assume that both ellipsoids $E_{1}$ and $E_{2}$ are centered at the origin. Let $L_{i}$ be the $i$-dimensional subspace spanned by the principal semi-axes of $E_{1}$ with lengths $\lambda_{d-i+1},\dots,\lambda_{d}$. The Courant–Fischer min-max theorem [HJ12, Theorem 4.2.6] yields $\lambda_{d-i+1}\leq\mu_{d-i+1}$, concluding the proof. ∎
It follows that the $i$-th (by length) principal semi-axis of $E^{\prime}$ is at most $\lambda^{i}$ for each $i\in[d]$. Moreover, at least $d-s$ of them must be equal. By direct computation, the maximal possible volume of $E^{\prime}$ is
${\operatorname{vol}\nolimits_{d}E}{\prod\limits_{j=0}^{d-s-1}\lambda^{-j}}.$
This bound is attained whenever $E^{\prime}$ is coaxial with $E$, and the lengths of some consecutive semi-axes of $E^{\prime}$ are equal to the corresponding $\lambda^{i}$ from $E$. However, this fixes the construction: ${F}^{\bot}$ must be the linear span of $d-s$ consecutive principal semi-axes of $E$. Thus, $E^{\prime}$ is centered at the origin. The desired statements follow since the ratio of the lengths of one collinear semi-axis of $E$ and $E^{\prime}$ is $\lambda$. ∎
We note that the above argument can be extended to any convex body $K$ in the following way.
Example 4.4.
Let $K$ be a convex body in $\mathbb{R}^{d}$, and let $A:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be the diagonal linear operator with diagonal entries, $\lambda,\ldots,\lambda^{d}$. Let $K^{\prime}=A^{\prime}(K)$ be a solution to Problem 4 with some linear operator $A^{\prime}$, for $K^{\prime\prime}=A(K)$, with any $d-s\geq 2$. Then, for any $\mu>0$ there is some value of $\lambda>0$ such that
  1. (a)
    for any $\mu^{\prime}<\mu$, no homothetic copy $x+\mu^{\prime}K^{\prime}$ contains $K^{\prime\prime}$,
  2. (b)
    $\frac{\operatorname{vol}\nolimits_{d}K^{\prime}}{\operatorname{vol}\nolimits_{d}K^{\prime\prime}}\leq\frac{1}{\mu}$.
Proof.
Let $B_{\mathrm{is}}$ be a largest ball contained in $K$, and let $B_{\mathrm{cs}}$ be a smallest ball containing $K$. Assume that $B_{\mathrm{cs}}=y+\nu B_{\mathrm{is}}$ for $\nu\geq 1$, $y\in\mathbb{R}^{d}$.
Let $\bar{\mu}>0$ be the smallest number such that a homothetic copy of $K^{\prime}$ of ratio $\bar{\mu}$ contains $K^{\prime\prime}$. Then there is a homothetic copy of $A^{\prime}(B_{\mathrm{cs}})$ of ratio $\bar{\mu}$ that contains $A(B_{\mathrm{is}})$. Consequently, there is a homothetic copy of $A^{\prime}(B_{\mathrm{is}})$ of ratio $\bar{\mu}\nu^{2}$ that contains $A(B_{\mathrm{cs}})$. Clearly, $A^{\prime}(B_{\mathrm{is}})\subseteq A(B_{\mathrm{cs}})$. Thus, by the previous example, $\bar{\mu}\nu^{2}\geq\lambda$, implying that $\bar{\mu}\geq\frac{\lambda}{\bar{\mu}^{2}}$, proving (a).
To prove (b), we may apply a similar consideration. ∎
5. Maximal volume ellipsoids of revolution
First, as immediate corollaries to Theorems 1.4 and 1.6, we obtain John type necessary conditions for an ellipsoid to be the largest volume ellipsoid of revolution in a given convex body. Next, as the main goal of the section we prove Theorem 1.7, and discuss the tightness of the results. Our proof strategy is to substitute $K$ with a suitable contact polytope; see Section 4.
5.1. Contact points on the sphere
In Theorems 1.4 and 1.6, we phrased our John type conditions in terms of the contact pairs of the two bodies. Since ellipsoids are affine images of the ball, we will, in this section, phrase our John type conditions for ellipsoids of revolution in terms of unit vectors that are pre-images of contact points under this affine transformation.
As an immediate corollary of Theorem 1.4, we obtain the following answer to Problem 1.
Corollary 5.1 (Necessary condition for John’s problem for ellipsoids with a fixed axis).
Let the origin-centered ellipsoid $E=A\mathbf{B}^{d}$ be a solution to the $F$-revolution John problem for a convex body $K$ in $\mathbb{R}^{d}$. Then there are unit vectors $u_{1},\dots,u_{m}$, and positive weights $\upalpha_{1},\dots,\upalpha_{m}$ such that $(Au_{1},{A}^{-1}u_{1}),\dots,(Au_{m},{A}^{-1}u_{m})$ are contact points of $E$ and $K$, and
$P_{F}\left(\sum\limits_{i\in[m]}\upalpha_{i}u_{i}\otimes u_{i}\right)P_{F}=P_{F};\quad\sum\limits_{i\in[m]}\upalpha_{i}u_{i}=\mathbf{0};\quad\sum\limits_{i\in[m]}\upalpha_{i}=d.$
Proof.
Observe that at a boundary point $Au\in\operatorname{bd}A\mathbf{B}^{d}$, the corresponding normal vector is $A^{-1}u$. The statement now follows from Theorem 1.4. ∎
Note that a standard argument using Carathéodory’s theorem shows that the number $m$ of points needed in Corollary 5.1 is at most $12d^{2}+d+$ 1.
As an immediate corollary of Theorem 1.6, we obtain the following answer to Problem 2.
Corollary 5.2 (Necessary condition for John’s problem for ellipsoids with any axis).
Assume that $E=A\mathbf{B}^{d}$ is a solution to the $F$-revolution John problem with any axis for a convex body $K$ in $\mathbb{R}^{d}$. Then there exist $m\leq 2d^{2}+d+1$, unit vectors $u_{1},\dots,u_{m}$, and positive weights $\upalpha_{1},\dots,\upalpha_{m}$ such that the conclusions of Corollary 5.1 hold as well as
$A\left(\sum\limits_{i\in[m]}\upalpha_{i}u_{i}\otimes u_{i}\right)A^{-1}={\left(A^{-1}\right)}^{\top}\left(\sum\limits_{i\in[m]}\upalpha_{i}u_{i}\otimes u_{i}\right){A}^{\top}.$
Definition 5.3.
We will call a position $K^{\prime}\in\mathcal{A}_{F}\!\left(K\right)$ with fixed axis $F$ of a convex body $K$ the John position with fixed axis $F$, if the standard unit ball is the solution to John’s general problem with fixed axis $F$ for $K^{\prime}$ and $\mathbf{B}^{d}.$
In the case $F$ is the whole space, the John position with fixed $F$ corresponds to the standard John position of a convex body.
Using the matrix $A$ from Corollary 5.1, we see that the position ${A}^{-1}K$ of $K$ is its John position with fixed axis $F$. Throughout this section, we fix the corresponding unit vectors $u_{1},\dots,u_{m}$ and positive weights $\upalpha_{1},\dots,\upalpha_{m}$. Considering the vectors $P_{F}u_{i}$ as lying in $F$, we obtain:
(5.1) $\sum\limits_{i\in[m]}\upalpha_{i}P_{F}u_{i}\otimes P_{F}u_{i}=\mathrm{Id}_{F};\quad\sum\limits_{i\in[m]}\upalpha_{i}P_{F}u_{i}=\mathbf{0};\quad\sum\limits_{i\in[m]}\upalpha_{i}=d.$
As in problems concerning convex bodies in John position, we will study the corresponding “outer contact polytopes”:
Denote
$K_{\text{J}}=\bigcap\limits_{i\in[m]}\left\{x\in\mathbb{R}^{d}:\;\left\langle x,u_{i}\right\rangle\leq 1\right\}.$
In the centrally symmetric case $K=-K$, define
$K_{\mathrm{J_{sym}}}=\bigcap\limits_{i\in[m]}\left\{x\in\mathbb{R}^{d}:\;|{\left\langle x,u_{i}\right\rangle}|\leq 1\right\}.$
5.2. Inradius in ${F}^{\bot}$
Lemma 5.4.
Let $v_{1},\dots,v_{m}$ be vectors in $\mathbf{B}^{n}.$ Assume there are positive weights $\beta_{1},\dots,\beta_{m}$ such that
$\sum\limits_{i\in[m]}\beta_{i}=1,\quad\sum\limits_{i\in[m]}\beta_{i}v_{i}=\mathbf{0},\quad\sum\limits_{i\in[m]}\beta_{i}\left|v_{i}\right|^{2}=\theta\text{ for some }\theta\in(0,1].$
Then the inradius of ${\left(\operatorname{conv}\{v_{1},\dots,v_{m}\}\right)}^{\circ}$ is at most $\frac{1}{\theta}.$
Proof.
Assume $r\mathbf{B}^{n}+c\subset{\left(\operatorname{conv}\{v_{1},\dots,v_{m}\}\right)}^{\circ}.$ This means that for every $i\in[m]$ with $v_{i}\neq 0,$ the point $c+r\frac{v_{i}}{\left|v_{i}\right|}$ is in ${\left(\operatorname{conv}\{v_{1},\dots,v_{m}\}\right)}^{\circ}.$ Hence,
$\left\langle v_{i},c+r\frac{v_{i}}{\left|v_{i}\right|}\right\rangle\leq 1\quad\Longleftrightarrow\quad\left\langle v_{i},c\right\rangle+r\left|v_{i}\right|\leq 1\quad\Longleftrightarrow\quad r\left|v_{i}\right|\leq 1-\left\langle v_{i},c\right\rangle.$
The latter inequality also holds if $v_{i}=\mathbf{0}.$ Multiplying it by $\beta_{i}$ and summing, one gets
$r\left(\sum\limits_{i\in[m]}\beta_{i}\left|v_{i}\right|\right)\leq\sum\limits_{i\in[m]}\beta_{i}-\left\langle\sum\limits_{i\in[m]}\beta_{i}v_{i},c\right\rangle.$
Since $\sum\limits_{i\in[m]}\beta_{i}v_{i}=\mathbf{0}$ and $\sum\limits_{i\in[m]}\beta_{i}=1,$ we get
$r\left(\sum\limits_{i\in[m]}\beta_{i}\left|v_{i}\right|\right)\leq 1\quad\Longleftrightarrow\quad r\leq\frac{1}{\sum\limits_{i\in[m]}\beta_{i}\left|v_{i}\right|}.$
Since $v_{i}\in\mathbf{B}^{n},$ we conclude that $\sum\limits_{i\in[m]}\beta_{i}\left|v_{i}\right|\geq\sum\limits_{i\in[m]}\beta_{i}{\left|v_{i}\right|^{2}}=\theta.$ The desired inequality follows. ∎
Lemma 5.5.
Under the notation of Subsection 5.1, with $s=\dim F$, the inradius of $K_{\text{J}}\cap{F}^{\bot}$ satisfies
$\operatorname{inrad}\!\left(K\cap{F}^{\bot}\right)\leq\frac{d}{d-s}\cdot\mathrm{radius}(E\cap{F}^{\bot}).$
Additionally, in the case when $K=-K$, the inradius of $K_{\mathrm{J_{sym}}}\cap{F}^{\bot}$ satisfies
$\operatorname{inrad}\!\left(K\cap{F}^{\bot}\right)\leq\sqrt{\frac{d}{d-s}}\cdot\mathrm{radius}(E\cap{F}^{\bot}).$
Proof.
The first identity in (5.1) yields $\sum\limits_{i\in[m]}\upalpha_{i}\left|P_{F}u_{i}\right|^{2}=s.$ Using here that $\sum\limits_{i\in[m]}\upalpha_{i}=d$, we conclude $\sum\limits_{i\in[m]}\upalpha_{i}\left|P_{{F}^{\bot}}u_{i}\right|^{2}=d-s.$
Since $\left|P_{{F}^{\bot}}u_{i}\right|\leq\left|u_{i}\right|=1$ for every $i\in[m]$, the vectors $P_{{F}^{\bot}}u_{1},\dots,P_{{F}^{\bot}}u_{m}$ satisfy the assumptions of Lemma 5.4 with $\beta_{1}=\frac{\upalpha_{1}}{d},\dots,\beta_{m}=\frac{\upalpha_{m}}{d},$ and $\theta=\frac{d-s}{d}.$ The general bound follows.
In the centrally symmetric case, we know that the origin is the center of an inball. Hence, the inradius cannot exceed the distance from the origin to a hyperplane $\left\{x\in{F}^{\bot}:\;\left\langle P_{{F}^{\bot}}u_{i},x\right\rangle\leq 1\right\}$ in ${F}^{\bot}$ for some non-zero $P_{{F}^{\bot}}u_{i}.$ This distance is exactly $\frac{1}{\left|P_{{F}^{\bot}}u_{i}\right|}.$ Identities (1.1) and (1.4) imply that $\left|P_{{F}^{\bot}}u_{j}\right|^{2}\geq\frac{d-s}{d}$ for some $j\in[m].$ The bound follows. ∎
5.3. Volume ratio in $F$
The proofs of our volume bounds follow very closely the proof of Theorem 1.1 of [AGB21], where the authors study upper bounds on the volume of a section of a convex body in John position. The authors of [AGB21] describe the corresponding section using the projections of the unit vectors that form the John decomposition in $\mathbb{R}^{d}$. In the rest of the proof in [AGB21], only algebraic identities (5.1) and Ball’s normalized version of the Brascamp–Lieb inequality [Bal89] are used.
Lemma 5.6 (Brascamp–Lieb inequality).
Let unit vectors $v_{1},\ldots,v_{m}\in\mathbb{R}^{n}$ and positive numbers $\beta_{1},\ldots,\beta_{m}$ satisfy
$\sum\limits_{i\in[m]}\beta_{i}v_{i}\otimes v_{i}=\mathrm{Id}_{\mathbb{R}^{n}}.$
Then
$\int_{\mathbb{R}^{n}}\prod_{i\in[m]}f_{i}(\left\langle x,v_{i}\right\rangle)^{\beta_{i}}\,\mathrm{d}x\leq\prod_{i\in[m]}\left(\int_{\mathbb{R}}f_{i}\right)^{\beta_{i}},$
holds for all integrable functions $f_{1},\ldots,f_{m}\colon\mathbb{R}\to[0,\infty).$
Lemma 5.7.
Under the notation of Subsection 5.1, with $s=\dim F$, the inequality
$\left(\frac{\operatorname{vol}\nolimits_{s}\left(K_{\mathrm{J_{sym}}}\cap F\right)}{\operatorname{vol}\nolimits_{s}\Box^{s}}\right)^{\frac{1}{s}}\leq\sqrt{\frac{d}{s}}$
holds in the centrally symmetric case, $K=-K$.
Proof.
Let $I\subset[m]$ be the set of those indices for which $P_{F}u_{i}\neq 0$. Set $v_{i}=\frac{P_{F}u_{i}}{\left|P_{F}u_{i}\right|}$ and $\beta_{i}={\upalpha_{i}}{\left|P_{F}u_{i}\right|^{2}}$ for all $i\in I.$ Then
$\sum_{i\in I}\beta_{i}v_{i}\otimes v_{i}=\sum_{i\in[m]}\upalpha_{i}P_{F}u_{i}\otimes P_{F}u_{i}=\mathrm{Id}_{F}.$
Let $f_{i}$ denote the indicator function of the interval $\left[-\frac{1}{\left|P_{F}u_{i}\right|},\frac{1}{\left|P_{F}u_{i}\right|}\right]$ for each $i\in I$.
We leave it to the reader to check that the integral on the left-hand side of the Brascamp–Lieb inequality (Lemma 5.6) is equal to the volume of $K_{\mathrm{J_{sym}}}\cap F$, as we identify $F$ with $\mathbb{R}^{n}$ (where $n=s$), and that the product of the integrals on the right-hand side of the Brascamp–Lieb inequality is equal to $2^{s}\prod\limits_{i\in I}\left|P_{F}u_{i}\right|^{-\upalpha_{i}\left|P_{F}u_{i}\right|^{2}}.$ The latter quantity is at most $2^{s}\left(\frac{d}{s}\right)^{\frac{s}{2}}$ by routine analysis, completing the proof of Lemma 5.7. ∎
Lemma 5.8.
Under the notation of Subsection 5.1, with $s=\dim F$, the inequality
$\left(\frac{\operatorname{vol}\nolimits_{s}\left(K_{\text{J}}\cap F\right)}{\operatorname{vol}\nolimits_{s}\Delta_{J}^{s}}\right)^{\frac{1}{s}}\leq\frac{1}{(s+1)^{\frac{d-s}{2s(d+1)}}}\sqrt{\frac{d(d+1)}{s(s+1)}}$
holds.
Proof.
Similarly to Ball’s argument [Bal91], we lift the picture to $\widetilde{F}=F\times\mathbb{R}$ and will write $(a,t)$ for a vector in $\widetilde{F}$ with $a\in F$ and $t\in\mathbb{R}.$
Let $I\subset[m]$ denote the subset of indices $i$ for which $P_{F}u_{i}\neq 0$. Denote $d^{\prime}=\sum\limits_{i\in[m]}\upalpha_{i};$ and set
$\widetilde{q}_{i}=\left(-\sqrt{\frac{d^{\prime}}{d^{\prime}+1}}P_{F}u_{i},\frac{1}{\sqrt{d^{\prime}+1}}\right),\quad\text{and }\quad\widetilde{\upalpha}_{i}=\frac{d^{\prime}+1}{d^{\prime}}\upalpha_{i}.$
Set $v_{i}=\frac{\widetilde{q}_{i}}{\left|\widetilde{q}_{i}\right|}$ and $\beta_{i}={\widetilde{\upalpha}_{i}}{\left|\widetilde{q}_{i}\right|^{2}}$ for all $i\in I.$ By straightforward computation,
$\sum\limits_{i\in I}\beta_{i}v_{i}\otimes v_{i}=\sum\limits_{i\in I}\widetilde{\upalpha}_{i}\widetilde{q}_{i}\otimes\widetilde{q}_{i}=\mathrm{Id}_{\widetilde{F}}.$
We will denote the function $t\mapsto e^{-\frac{t}{\left|\widetilde{q}_{i}\right|}}\mathbf{1}_{[0,+\infty)}$ by $f_{i}$ for each $i\in I,$ where $\mathbf{1}_{S}$ stands for the indicator function of the set $S.$ Then, the product of the integrals on the right-hand side of the Brascamp–Lieb inequality (Lemma 5.6) is equal to $\prod\limits_{i\in I}\left|P_{F}u_{i}\right|^{\upalpha_{i}\left|P_{F}u_{i}\right|^{2}}.$ The bound
$\prod\limits_{i\in I}\left|P_{F}u_{i}\right|^{\upalpha_{i}\left|P_{F}u_{i}\right|^{2}}\leq\frac{(s+1)^{\frac{s+1}{2(d^{\prime}+1)}}}{(d^{\prime}+1)^{\frac{d^{\prime}+1}{2(d^{\prime}+1)}}}$
follows from the standard majorization argument for the convex function $t\mapsto t\ln t$ on $(0,1)$. This bound was obtained in the proof of Theorem 1.1 in [AGB21]. For the sake of completeness, we provide the proof as a separate lemma in Appendix B.
We claim that the integral on the left-hand side of the Brascamp–Lieb inequality is
$\int_{\tilde{F}}\prod_{i\in I}f_{i}(\left\langle x,v_{i}\right\rangle)^{\beta_{i}}\,\mathrm{d}x=\frac{s!}{(d^{\prime}+1)^{\frac{s+1}{2}}{d^{\prime}}^{\frac{s}{2}}}\operatorname{vol}\nolimits_{s}\left(K_{\text{J}}\cap F\right).$
Indeed, set $f(x)=\prod\limits_{i\in I}f_{i}(\left\langle x,v_{i}\right\rangle)^{\beta_{i}}$ for $x\in\tilde{F}$. The support of $f$ is the cone $\mathrm{Pos}\ \!{\left\{\left(\frac{K_{\text{J}}\cap F}{\sqrt{d^{\prime}}},1\right)\right\}}$. Now, $f$ is constant $e^{-\tau{\sqrt{d^{\prime}+1}}}$ at the intersection of the hyperplane $H_{\tau}=\left\{(y,t)\in\widetilde{F}:\;t=\tau\right\}$ and this cone for any positive $\tau$. By monotonicity in $d^{\prime}$, the proof of Lemma 5.8 is complete. ∎
Remark 5.9.
The authors of [AGB21] used a different approach in the general case: they first lift the “contact vectors” to $\mathbb{R}^{d}\times\mathbb{R}$ and then project the configuration onto $F\times\mathbb{R}.$ This minor variation in the argument does not substantially affect the subsequent computations.
5.4. Proof of Theorem 1.7
A transformation of the space by an $F$-matrix changes both the volume form on $F$ and the distance in ${F}^{\bot}$ proportionally. Hence, under the notation of Subsection 5.1, it suffices to bound the volume of ${A}^{-1}K\cap F$ and the inradius of ${A}^{-1}K\cap{F}^{\bot}$ to obtain the corresponding bound.
By construction, ${A}^{-1}K\subset K_{\text{J}}$ and ${A}^{-1}K\subset K_{\mathrm{J_{sym}}}$ in the centrally symmetric case $K=-K.$ The corresponding volume bounds follow from Lemma 5.7 and Lemma 5.8; the inradius bound follows from Lemma 5.5.
The inclusion $E\cap F\subset K\cap F\subset d\left(E\cap F\right)$ follows from Corollary 1.5, since $z$ from the assertion of the corollary is zero and $P_{F}E=E\cap F.$
Let us now establish the inclusion
$E\cap F\subset K\cap F\subset\sqrt{d}\left(E\cap F\right)$
in the centrally symmetric case $K=-K$. Fix $x\in K_{\mathrm{J_{sym}}}\cap F.$ One has
[…]
$\sum\limits_{i\in[m]}\upalpha_{i}\left\langle x,u_{i}\right\rangle\left\langle u_{i},x\right\rangle=\sum\limits_{i\in[m]}\upalpha_{i}\left\langle x,u_{i}\right\rangle^{2}\leq\sum\limits_{i\in[m]}\upalpha_{i}=d.$
The desired bound follows. The proof of Theorem 1.7 is complete.
5.5. Tightness of the results
The inclusions of Theorem 1.7 are optimal since they are already optimal for the John ellipsoid. The volume bounds of Lemma 5.7 and Lemma 5.8 coincide with the corresponding bounds for an $s$-dimensional section of a $d$-dimensional convex body in John position (see Theorem 1.1 of [AGB21]). In the centrally symmetric case $K=-K,$ we conjecture the following:
Conjecture 5.10.
Let an origin-centered ellipsoid $E$ be a solution to Problem 1 for a centrally symmetric convex body $K$ and a linear subspace $F$ in $\mathbb{R}^{d}$ of dimension $s\in\{0,1,\dots,d\}$. Then
$\frac{\operatorname{vol}\nolimits_{s}\left(K\cap F\right)}{\operatorname{vol}\nolimits_{s}\left(E\cap F\right)}\leq C_{\Box}(d,s)\frac{\operatorname{vol}\nolimits_{s}\Box^{s}}{\operatorname{vol}\nolimits_{s}\mathbf{B}^{s}},$
where the constant $C_{\Box}(d,s)$ is given by
$C^{2}_{\Box}(d,s)=\left\lceil\frac{d}{s}\right\rceil^{d-s\lfloor d/s\rfloor}\left\lfloor\frac{d}{s}\right\rfloor^{s-(d-s\lfloor d/s\rfloor)}.$
The constant $C_{\Box}(d,s)$ in Conjecture 5.10 is attained on certain sections of the cube $\Box^{d}=[-1,1]^{d}$ by $s$-dimensional subspaces (see Lemma 2.2 of [IT21]). Peculiarly, in the case when $K$ is centrally symmetric and $s$ divides $d,$ both our volume ratio bound in $F$ and the inradius bound in ${F}^{\bot}$ are simultaneously tight: $K$ is the cube $[-1,1]^{d};$ $E$ is the John ellipsoid $\mathbf{B}^{d}$ of $K;$ $F$ is the solution set to the system of linear equations:
$\left\langle x,e_{\frac{d}{s}(i-1)+1}\right\rangle=\ldots=\left\langle x,e_{\frac{d}{s}i}\right\rangle,\quad\text{for all}\quad i\in[s].$
Moreover, the bound $\sqrt{\frac{d}{d-s}}$ on the inradius in ${F}^{\bot}$ is attained on certain sections of the cube $\Box^{d}.$ Indeed, denoting the standard basis of $\mathbb{R}^{d}$ by $u_{1},\dots,u_{d},$ the estimate $\sqrt{\frac{d}{d-s}}$ is attained whenever $\left|P_{{F}^{\bot}}u_{1}\right|^{2}=\dots=\left|P_{{F}^{\bot}}u_{d}\right|^{2}=\frac{d-s}{d}.$ The existence of subspaces with this property was proven in Lemma 2.4 of [Iva20], in which the set of all possible lengths of the projections was described.
The general case of a not necessarily symmetric body is more intricate. As shown in [Dir17], the bound in Lemma 5.8 is asymptotically tight for a regular simplex $\Delta_{J}^{d}$ in the John position. Consequently, it is asymptotically tight in our case. However, we believe that there is an intrinsic problem in obtaining the best possible bound using the Brascamp–Lieb inequality directly: if one chooses the parameters to maximize the functional in Lemma B.1, then the Brascamp–Lieb inequality is not optimal for the corresponding functions. That is, there are geometric obstructions to achieving the equality case in the Brascamp–Lieb inequality.
We also have an example showing that the bound Theorem 1.7 (1) on the inradius is the best possible. It is technical, and we state and prove it in Theorem A.5. Nevertheless, the corresponding vectors in $\mathbb{R}^{d-s}$ can be lifted to unit vectors satisfying the algebraic equations of Corollary 5.1. A more detailed analysis of Lemma 5.4 shows that the bound can be achieved, if certain vectors are close to zero and the others lie near the unit sphere, so it cannot be achieved via the projections of the unit vectors from the standard John condition for $F=\mathbb{R}^{d}$ (see [BP09]).
Finally, [IN22] by M. Naszódi and G. Ivanov may be interpreted as studying John’s general problem with fixed axis $F$ for $\mathbf{B}^{n+1}$ and $K$ being a certain not necessarily convex set in $\mathbb{R}^{n+1}$ associated with a log-concave function on $\mathbb{R}^{n}$, and $F$ being $\mathbb{R}^{n}.$ Moreover, the algebraic conditions for the “contact points” in that case coincide with those of Corollary 5.1; and the more general case considered in [IN23] fits within the framework of Theorem 1.4 for appropriately chosen parameter sets. The inclusions for ellipsoids and “contact-type” polytopes have been studied previously in the functional setting as well; see, for example, [Iva24, Lemma 4.1].
6. Uniqueness of the solution
We note that in general, the solution of John’s problem for ellipsoids with any axis of revolution is not unique: consider the convex body $K=E+\left[-v,v\right]$, where $E$ is an $F$-ellipsoid for some $s$-dimensional subspace $F$ and $v\in{F}^{\bot}$. On the other hand, we have the following theorem.
For a linear subspace $F$ and a subset $K$ of $\mathbb{R}^{d}$, we denote by $\mathcal{S}_{F}(K)$ the set of those elements of $\mathcal{A}_{F}\!\left(K\right)$, for which the corresponding linear transformation is positive definite.
Theorem 6.1.
For any convex bodies $K,L$ in $\mathbb{R}^{d}$, and for any elements $K_{1},K_{2}\in\mathcal{S}_{F}(K)$ of maximal volume satisfying $K_{1},K_{2}\subseteq L$, $K_{1}$ and $K_{2}$ are translates of each other.
Proof.
Consider the family of pairs $(S,x)$, where $F$ is a positive semidefinite $F$-operator, $x\in\mathbb{R}^{d}$ satisfying $x+S(K)\subseteq L$. We regard $(S,x)$ as a point of $\mathbb{R}^{d^{2}+d}$. It is easy to see that the family of pairs $(S,x)$ with the above condition is a compact, convex subset of $\mathbb{R}^{d^{2}+d}$, and thus, $\operatorname{vol}\nolimits_{d}\left(x+S(K)\right)=\det(S)\operatorname{vol}\nolimits_{d}K$ attains its maximum on it. Note also that if $x+S(K)$ maximizes $\operatorname{vol}\nolimits_{d}\left(x+S(K)\right)$ on this set, then $S$ is positive definite, as otherwise $\operatorname{vol}\nolimits_{d}\left(x+S(K)\right)=0$. Thus, there are some positive definite $F$-operators $S_{1},S_{2}$ and vectors $x_{1},x_{2}\in\mathbb{R}^{d}$ such that for $i=1,2$, $K_{i}=x_{i}+S_{i}(K)$.
By convexity, we have that $\bar{K}=\frac{x_{1}+x_{2}}{2}+\frac{1}{2}(S_{1}+S_{2})(K)\subseteq L$. On the other hand, it is well-known that for any two positive definite matrices $A,B$,
$\det\frac{A+B}{2}\geq\sqrt{\det A\cdot\det B}$
with equality if and only if $A=B$ (see e.g. [HJ12, Corollary 7.6.8]). Since $\frac{1}{2}(S_{1}+S_{2})$ is also a positive definite $F$-operator, it follows from the maximality of volume that $S_{1}=S_{2}$, implying that $K_{2}$ and $K_{1}$ are translates of each other. ∎
Remark 6.2.
We note that $\mathcal{S}_{F}(\mathbf{B}^{d})=\mathcal{A}_{F}\!\left(\mathbf{B}^{d}\right)$ for every linear subspace $F$, by the polar decomposition of matrices.
Remark 6.3.
We note that it is not true that if $K_{1},K_{2}\in\mathcal{S}_{F}(K)$ are of maximal volume satisfying $K_{1},K_{2}\subseteq L$, then the translation vector moving $K_{1}$ to $K_{2}$ is necessarily parallel to ${F}^{\bot}$, as shown by the following example. Set $d=2$ and $K=\mathbf{B}^{2}$. We define $L$ as an $\mathbf{0}$-symmetric rectangle circumscribed about $\mathbf{B}^{2}$ such that the sides of $L$ are parallel to the lines $y=\pm x$, and the only points of $\operatorname{bd}L$ on $\mathbf{B}^{2}$ are the points ±$\pm\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right)$. Then an elementary computation shows that the only $\mathbf{0}$-symmetric $F$-ellipse contained in $L$ is $\mathbf{B}^{2}$, and this is of maximal area. On the other hand, moving this ellipse along the line $y=-x$ yields other maximal area $F$-ellipses in $L$.
7. $F$-revolution Löwner ellipsoids
A dual construction to the largest volume ellipsoid contained in a convex body is the smallest volume ellipsoid containing the body, usually called the Löwner ellipsoid. Notably, the necessary and sufficient conditions for the Euclidean unit ball to be the Löwner ellipsoid coincide with the conditions in John’s characterization of the largest volume ellipsoid. The extension of the notion of the Löwner ellipsoid is straightforward – swap the bodies in John’s general problem with fixed axis $F$, that is, consider John’s general problem with fixed axis $F$ for $K$ and $\mathbf{B}^{d}.$
Definition 7.1.
We will call a position $K^{\prime}\in\mathcal{A}_{F}\!\left(K\right)$ with fixed axis $F$ of a convex body $K$ the Löwner position with fixed axis $F$ if $K^{\prime}$ is a solution to John’s general problem with fixed axis $F$ for $K$ and $\mathbf{B}^{d}.$
The following result is a dual version of Theorem 1.7:
Theorem 7.2.
Let a convex body $K$ be in the Löwner position with fixed axis $F$ of dimension $s\in\{0,1,\dots,d\}$. Then
  1. (1)
    the following outer volume bound holds:
    $\left(\frac{\operatorname{vol}\nolimits_{s}P_{F}K}{\operatorname{vol}\nolimits_{s}\Delta_{L}^{s}}\right)^{\frac{1}{s}}\geq\sqrt{\frac{s}{d}},$
    where $\Delta_{L}^{s}$ stands for the $s$-dimensional simplex inscribed in $\mathbf{B}^{s}.$
  2. (2)
    the following inclusion holds:
    $P_{F}K\subset P_{F}E\subset d\cdot P_{F}K.$
  3. (3)
    the circumradius of $P_{{F}^{\bot}}K$ is at least $\sqrt{\frac{d-s}{d}},$
If, additionally, $K$ is centrally symmetric, then
  1. (S1)
    the following outer volume bound holds:
    $\left(\frac{\operatorname{vol}\nolimits_{s}P_{F}K}{\operatorname{vol}\nolimits_{s}\Diamond^{s}}\right)^{\frac{1}{s}}\geq\sqrt{\frac{s}{d}},$
    where $\Diamond^{s}$ stands for the standard cross-polytope.
  2. (S2)
    the following inclusion holds:
    $P_{F}K\subset P_{F}E\subset\sqrt{d}\cdot P_{F}K.$
We will prove this theorem in the current section. The proof mostly consists of standard arguments dual to those used in the proof of Theorem 1.7. The only surprise is that the bound (3) on the circumradius does not require $K$ to be symmetric!
7.1. Notation
As a direct corollary of Theorem 1.4, we obtain:
Corollary 7.3.
Let a convex body $K$ be in Löwner position with fixed axis $F.$ Then there exist $m\leq 2d^{2}+d+1$, unit vectors $u_{1},\dots,u_{m}$, and positive weights $\upalpha_{1},\dots,\upalpha_{m}$ such that $(u_{1},u_{1}),\dots,(u_{m},u_{m})$ are contact pairs of $K$ and $\mathbf{B}^{d}$, and
$P_{F}\left(\sum\limits_{i\in[m]}\upalpha_{i}u_{i}\otimes u_{i}\right)P_{F}=P_{F};\quad\sum\limits_{i\in[m]}\upalpha_{i}u_{i}=\mathbf{0};\quad\sum\limits_{i\in[m]}\upalpha_{i}=d.$
In the case where $F$ is the whole space, the Löwner position with fixed axis $F$ coincides with the standard Löwner position of a convex body.
Throughout this section, we fix the corresponding unit vectors $u_{1},\dots,u_{m}$ and positive weights $\upalpha_{1},\dots,\upalpha_{m}$. Considering the vectors $P_{F}u_{i}$ as lying in $F$, we observe that they satisfy equations (5.1).
As in problems concerning convex bodies in the John position, we will study the corresponding “contact polytopes”:
Define
$K_{\text{L}}=\operatorname{conv}\left\{u_{1},\dots,u_{m}\right\}.$
In the centrally symmetric case $K=-K$, define
$K_{\mathrm{L_{sym}}}=\operatorname{conv}\left\{\pm u_{1},\dots,\pm u_{m}\right\}.$
7.2. Circumradius in ${F}^{\bot}$
Lemma 7.4.
Let $v_{1},\dots,v_{m}$ be vectors in $\mathbb{R}^{n}.$ Assume there are positive weights $\beta_{1},\dots,\beta_{m}$ such that
$\sum\limits_{i\in[m]}\beta_{i}=1,\quad\sum\limits_{i\in[m]}\beta_{i}v_{i}=\mathbf{0},\quad\sum\limits_{i\in[m]}\beta_{i}\left|v_{i}\right|^{2}=\theta.$
Then the circumradius of $\operatorname{conv}\{v_{1},\dots,v_{m}\}$ is at least $\sqrt{\theta}.$
Proof.
Assume $r\mathbf{B}^{n}+c\supset\operatorname{conv}\{v_{1},\dots,v_{m}\}.$ This means that $\left|c-v_{i}\right|^{2}\leq r^{2}$ for every $i\in[m].$ Multiplying both sides by $\beta_{i}$ and summing over $i$, we get:
$r^{2}\left(\sum\limits_{i\in[m]}\beta_{i}\right)\geq\sum\limits_{i\in[m]}\beta_{i}\left|c-v_{i}\right|^{2}=\left|c\right|^{2}\left(\sum\limits_{i\in[m]}\beta_{i}\right)-2\left\langle c,\sum\limits_{i\in[m]}\beta_{i}v_{i}\right\rangle+\sum\limits_{i\in[m]}\beta_{i}\left|v_{i}\right|^{2}.$
Since $\sum\limits_{i\in[m]}\beta_{i}v_{i}=\mathbf{0}$ and $\sum\limits_{i\in[m]}\beta_{i}=1$, we obtain
$r^{2}\geq\sum\limits_{i\in[m]}\beta_{i}\left|v_{i}\right|^{2}+\left|c\right|^{2}\geq\sum\limits_{i\in[m]}\beta_{i}\left|v_{i}\right|^{2}=\theta.$
The desired inequality follows. ∎
Combining (5.1) with the above lemma with $v_{i}=P_{F^{\perp}}u_{i}$ and $\beta_{i}=\upalpha_{i}/d$, and by identifying $F^{\perp}$ with $\mathbb{R}^{n}$ (where $n=d-s$), we conclude the following:
Lemma 7.5.
Under the notation of Subsection 7.1, the circumradius of $P_{{F}^{\bot}}K_{L}$ is at least $\sqrt{\frac{d-s}{d}}.$
7.3. Outer volume ratio in ${F}^{\bot}$
The proofs of our outer volume bounds follow from Theorem 1.2 of [AGB21], where the authors study lower bounds on the volume of projections of convex bodies in the Löwner position. Starting with the breakthrough work [Bar97], the main tool for obtaining outer volume ratio bounds is the following reverse form of the Brascamp–Lieb inequality:
Lemma 7.6 (Reverse Brascamp–Lieb inequality).
Let unit vectors $v_{1},\ldots,v_{m}\in\mathbb{R}^{n}$ and positive numbers $\beta_{1},\ldots,\beta_{m}$ satisfy $\sum\limits_{i\in[m]}\beta_{i}v_{i}\otimes v_{i}=\mathrm{Id}_{\mathbb{R}^{n}}.$ Define
$G(x)=\sup\left\{\prod\limits_{i\in[m]}g_{i}(\theta_{i})^{\beta_{i}}:\;x=\sum\limits_{i\in[m]}\beta_{i}\theta_{i}v_{i}\right\}$
for integrable functions $g_{1},\ldots,g_{m}\colon\mathbb{R}\to[0,\infty).$ Then,
$\int_{\mathbb{R}^{d}}G\geq\prod\limits_{i\in[m]}\left(\int g_{i}\right)^{\beta_{i}}.$
To simplify the reading, we provide a sketch of the proof, omitting technical details.
Lemma 7.7.
Under the notation of Subsection 7.1, the inequality
$\left(\frac{\operatorname{vol}\nolimits_{s}P_{F}K_{\mathrm{L_{sym}}}}{\operatorname{vol}\nolimits_{s}\Diamond^{s}}\right)^{\frac{1}{s}}\geq\sqrt{\frac{s}{d}}$
holds.
Proof.
Let $I\subset[m]$ be the set of those indices for which $P_{F}u_{i}\neq 0.$ Set $v_{i}=\frac{P_{F}u_{i}}{\left|P_{F}u_{i}\right|}$ and $\beta_{i}={\upalpha_{i}}{\left|P_{F}u_{i}\right|^{2}}$ for all $i\in I.$ Then
$\sum_{i\in I}\beta_{i}v_{i}\otimes v_{i}=\sum_{i\in[m]}\upalpha_{i}P_{F}u_{i}\otimes P_{F}u_{i}=\mathrm{Id}_{F}.$
Define
$g_{i}(t)=e^{-\frac{\left|t\right|}{\left|P_{F}u_{i}\right|}}$
for each $i\in I$.
We leave it to the reader to verify that the product of the integrals on the right-hand side of the reverse Brascamp–Lieb inequality (Lemma 7.6) equals $2^{s}\prod\limits_{i\in I}\left|P_{F}u_{i}\right|^{\upalpha_{i}\left|P_{F}u_{i}\right|^{2}}$. This is at least $2^{s}\left(\frac{d}{s}\right)^{\frac{s}{2}}$ by routine analysis.
On the other hand, the corresponding function $G$ equals $e^{-\left\|\cdot\right\|_{K_{\mathrm{L_{sym}}}}},$ where $\left\|\cdot\right\|_{K_{\mathrm{L_{sym}}}}$ is the gauge function of $K_{\mathrm{L_{sym}}}:$
$\left\|x\right\|_{K_{\mathrm{L_{sym}}}}=\inf\left\{t>0:\;x\in tK_{\mathrm{L_{sym}}}\right\}.$
Hence, by standard computation, the integral of $G$ is
$\int_{\mathbb{R}^{s}}G=s!\operatorname{vol}\nolimits_{s}P_{F}K_{\mathrm{L_{sym}}}.$
The desired bound follows, since $\operatorname{vol}\nolimits_{s}\Diamond^{s}=\frac{2^{s}}{s!}.$
Lemma 7.8.
Under the notation of Subsection 7.1, the inequality
$\left(\frac{\operatorname{vol}\nolimits_{s}P_{F}K_{L}}{\operatorname{vol}\nolimits_{s}\Delta_{L}^{s}}\right)^{\frac{1}{s}}\geq\sqrt{\frac{s}{d}}$
holds.
Proof.
Following Barthe’s argument [Bar97], we lift the picture to $\widetilde{F}=F\times\mathbb{R}$, and write $(a,t)$ for a vector in $\widetilde{F}$ with $a\in F$ and $t\in\mathbb{R}.$ Let $I\subset[m]$ be the index set of those $i$ such that $P_{F}u_{i}\neq 0.$ Let $d^{\prime}=\sum\limits_{i\in[m]}\upalpha_{i}$, and define
$\widetilde{q}_{i}=\left(-\sqrt{\frac{d^{\prime}}{d^{\prime}+1}}P_{F}u_{i},\frac{1}{\sqrt{d^{\prime}+1}}\right),\quad\widetilde{\upalpha}_{i}=\frac{d^{\prime}+1}{d^{\prime}}\upalpha_{i}.$
Set $v_{i}=\frac{\widetilde{q}_{i}}{\left|\widetilde{q}_{i}\right|}$ and $\beta_{i}=\widetilde{\upalpha}_{i}\left|\widetilde{q}_{i}\right|^{2}$ for all $i\in I.$ Then,
$\sum\limits_{i\in I}\beta_{i}v_{i}\otimes v_{i}=\sum\limits_{i\in I}\widetilde{\upalpha}_{i}\widetilde{q}_{i}\otimes\widetilde{q}_{i}=\mathrm{Id}_{\widetilde{F}}.$
Define a gauge-like function $C$ on $\widetilde{F}$ by
$C(x)=\inf\left\{\sum_{i\in I}\lambda_{i}\colon\lambda_{i}\geq 0\text{ and }x=\sum_{i\in I}\lambda_{i}\widetilde{q}_{i}\right\},$
where we set $C(x)=+\infty$ if such a decomposition does not exist. Let $G(x)=e^{-C(x)}$.
By convexity, the support of $G$ is the cone $\mathrm{Pos}\ \!{\left(P_{F}K_{\text{L}},\frac{1}{\sqrt{d^{\prime}}}\right)}$; the function $G$ is constant and equal to $e^{-\tau\sqrt{d^{\prime}+1}}$ on the intersection of this cone with the hyperplane $H_{\tau}=\left\{(x,t)\in\widetilde{F}:\;t=\tau\right\}$ for any $\tau>0$. Thus,
$\int_{\widetilde{F}}G=\operatorname{vol}\nolimits_{s}P_{F}K_{\text{L}}\cdot\int_{0}^{\infty}e^{-r\sqrt{d^{\prime}+1}}(r\sqrt{d^{\prime}})^{s}\,\mathrm{d}r=\frac{s!(d^{\prime})^{s/2}}{(d^{\prime}+1)^{(s+1)/2}}\operatorname{vol}\nolimits_{s}P_{F}K_{\text{L}}.$
Now define
$g_{i}(t)=e^{-t/\left|\widetilde{q}_{i}\right|}\cdot\mathbf{1}_{[0,+\infty)}(t)$
for each $i\in I$. Then, for any $x$ in the support of $G,$
$\sup\left\{\prod\limits_{i\in I}g_{i}(\theta_{i})^{\beta_{i}}:\;x=\sum\limits_{i\in I}\beta_{i}\theta_{i}v_{i}\right\}\stackrel{{\scriptstyle\lambda_{i}=\widetilde{\upalpha}_{i}\left|\widetilde{q}_{i}\right|\theta_{i}}}{{=}}e^{-\inf\left\{\sum\limits_{i\in I}\lambda_{i}:\;x=\sum\limits_{i\in I}\lambda_{i}\widetilde{q}_{i}\right\}}=G(x).$
Thus, by the reverse Brascamp–Lieb inequality (Lemma 7.6),
$\int_{\mathbb{R}^{d}}G\geq\prod\limits_{i\in I}\left|\widetilde{q}_{i}\right|^{\widetilde{\upalpha}_{i}\left|\widetilde{q}_{i}\right|^{2}}.$
Applying Jensen’s inequality with $\sum\limits_{i\in I}\widetilde{\upalpha}_{i}\left|\widetilde{q}_{i}\right|^{2}=s+1$ and $\sum\limits_{i\in I}\widetilde{\upalpha}_{i}=d^{\prime}$, we obtain
$\prod\limits_{i\in I}\left|\widetilde{q}_{i}\right|^{\widetilde{\upalpha}_{i}\left|\widetilde{q}_{i}\right|^{2}}\geq\left(\frac{s+1}{d^{\prime}+1}\right)^{(s+1)/2}.$
Substituting into the expression for $\int_{\widetilde{F}}G$, we find
$\operatorname{vol}\nolimits_{s}P_{F}K_{\text{L}}\geq\left(\frac{s+1}{d^{\prime}+1}\right)^{(s+1)/2}\frac{(d^{\prime}+1)^{(s+1)/2}}{s!(d^{\prime})^{s/2}}=\left(\frac{s}{d^{\prime}}\right)^{s/2}\operatorname{vol}\nolimits_{s}\Delta_{L}^{s}.$
The desired bound follows since $d^{\prime}\leq d.$
Remark 7.9.
The bound is optimal whenever $s+1$ divides $d+1$. It is attained when $K=\Delta_{L}^{d}$ and $F$ is the subspace spanned by the centroids of pairwise disjoint $\left(\frac{d+1}{s+1}-1\right)$-dimensional faces of $\Delta_{L}^{d}.$
7.4. Proof of Theorem 7.2
Proof of Theorem 7.2.
We use the notation introduced in Subsections 5.1 and 7.1. By duality, $K_{\text{L}}={\left(K_{\text{J}}\right)}^{\circ}$ and $K_{\mathrm{L_{sym}}}={\left(K_{\mathrm{J_{sym}}}\right)}^{\circ}.$ Hence, the inclusions follow from the corresponding inclusions of Theorem 1.7. The bound on the circumradius was obtained in Lemma 7.5. The volume bounds were proven in Lemma 7.8 in the general case and in Lemma 7.7 in the centrally symmetric case $K=-K.$
Appendix A Tightness of the inradius bound
The goal of this section is to state and prove Theorem A.5 according to which the bound Theorem 1.7 (1) on the inradius is the best possible.
Define
$a=\frac{1}{\sqrt{n}}(1,\dots,1)\in\mathbb{R}^{n},\quad n\geq 2.$
We choose vectors $v_{1},\dots,v_{n+1}$ that form a directionally dilated regular simplex along the direction of $a$.
For each $i\in[n]$, define the vector $p_{i}\in\mathbb{R}^{n}$ by
$p_{i}[j]=\begin{cases}\frac{n-1}{n},&\text{if }j=i,\\ -\frac{1}{n},&\text{if }j\neq i.\end{cases}$
Then $\left|p_{i}\right|^{2}=\frac{n-1}{n}.$ Define $q_{i}=\frac{p_{i}}{\left|p_{i}\right|}\in\mathbb{R}^{n}.$ Note that for each $i\in[n]$, the vectors $p_{i}$ and $q_{i}$ belong to ${a}^{\bot}$.
Define the vectors
$v_{i}=tq_{i}+\sqrt{1-t^{2}}\cdot a,\quad\text{for }i\in[n]$
Now, fix a constant $\gamma$ satisfying
(A.1) $n\gamma<1.$
We observe that
$\sum_{i\in[n]}p_{i}=\mathbf{0}=\sum_{i\in[n]}q_{i},$
$\sum_{i\in[n]}v_{i}=t\sum_{i\in[n]}q_{i}+n\sqrt{1-t^{2}}\cdot a=n\sqrt{1-t^{2}}\cdot a.$
Define $v_{n+1}$ by the identity
$\gamma\sum_{i\in[n]}v_{i}+(1-n\gamma)v_{n+1}=\mathbf{0}.$
So,
$v_{n+1}=-\frac{\gamma n}{1-\gamma n}\sqrt{1-t^{2}}\cdot a.$
Our second restriction is
(A.2) $\frac{\gamma n}{1-\gamma n}\sqrt{1-t^{2}}<1.$
Denote $\Delta=\operatorname{conv}\left\{v_{1},\dots,v_{n+1}\right\}$. Let us estimate the inradius of ${\Delta}^{\circ}.$
Lemma A.1.
The vertices of the simplex ${\Delta}^{\circ}$ are
$x_{n+1}=\frac{1}{\sqrt{1-t^{2}}}a,$
and
$x_{i}=-\frac{1}{n\gamma}\frac{(n-1)}{t}q_{i}-\frac{1-n\gamma}{n\gamma}\cdot\frac{1}{\sqrt{1-t^{2}}}a,\quad\text{for }i\in[n].$
Proof.
The vertex of ${\Delta}^{\circ}$ corresponding to the facet $\operatorname{conv}\{v_{1},\dots,v_{n}\}$ of $\Delta$ is
$x_{n+1}=\frac{1}{\sqrt{1-t^{2}}}a.$
Clearly, $\left\langle x_{n+1},v_{n+1}\right\rangle<0$ and $\left\langle x_{n+1},v_{i}\right\rangle=1$ for all $i\in[n]$.
By direct computation, for each $i\in[n]$,
$\left\langle x_{i},v_{i}\right\rangle=-\frac{1}{n\gamma}\cdot\frac{n-1}{t}\cdot t\left\langle q_{i},q_{i}\right\rangle-\frac{1-n\gamma}{n\gamma}\cdot\frac{1}{\sqrt{1-t^{2}}}\cdot\sqrt{1-t^{2}}\left\langle a,a\right\rangle=-\frac{n-n\gamma}{n\gamma},$
where we used the fact that $\left\langle q_{i},a\right\rangle=0$. Note that if inequality (A.1) holds, then $\left\langle x_{i},v_{i}\right\rangle<0$ for every $i\in[n]$.
Since $\left\langle q_{i},q_{j}\right\rangle=-\frac{1}{n-1}$ whenever $j\in[n]$ and $i\neq j$, we conclude that
$\left\langle x_{i},v_{j}\right\rangle=-\frac{1}{n\gamma}\cdot\frac{n-1}{t}\cdot t\left\langle q_{i},q_{j}\right\rangle-\frac{1-n\gamma}{n\gamma}\cdot\frac{1}{\sqrt{1-t^{2}}}\cdot\sqrt{1-t^{2}}\left\langle a,a\right\rangle=\frac{1}{n\gamma}-\frac{1-n\gamma}{n\gamma}=1$
for all $j\neq i$, as required for duality. Clearly, $\left\langle x_{i},v_{n+1}\right\rangle=1$ for all $i\in[n]$. The lemma follows. ∎
Remark A.2.
There is no mystery in the explicit formulas for the vertices of ${\Delta}^{\circ}$. By symmetry, the facet of ${\Delta}^{\circ}$ opposite to $x_{n+1}$ must be parallel to ${a}^{\bot}$. Therefore, the $a$-component can be found directly from the equation $\left\langle x_{i},v_{n+1}\right\rangle=1$. On the other hand, this facet must be a translation of the simplex $-\tau\operatorname{conv}\{q_{1},\dots,q_{n}\}$ for some positive $\tau$. Indeed, this facet is a homothet of ${\Delta}^{\circ}\cap{a}^{\bot}$ with center at $x_{n+1}$. The set ${\Delta}^{\circ}\cap{a}^{\bot}$ is the regular simplex
$\operatorname{conv}\left\{\frac{1-n}{t}q_{1},\dots,\frac{1-n}{t}q_{n}\right\},$
since ${\Delta}^{\circ}\cap{a}^{\bot}$ is the polar in ${a}^{\bot}$ of the projection of $\Delta$ onto ${a}^{\bot}$. This projection is the regular simplex $\operatorname{conv}\left\{tq_{1},\dots,tq_{n}\right\}$.
Lemma A.3.
The section of ${\Delta}^{\circ}$ by the plane $L$ spanned by the vectors $x_{1}$ and $x_{n+1}$ is the triangle $T$ with vertices $x_{1}$, $x_{n+1}$, and $y$, given by
$y=\frac{1}{n\gamma}\cdot\frac{q_{1}}{t}-\frac{1-n\gamma}{n\gamma}\cdot\frac{a}{\sqrt{1-t^{2}}}.$
Proof.
Clearly,
$L=\operatorname{span}\left\{v_{1},v_{n+1}\right\}=\operatorname{span}\left\{q_{1},a\right\},$
and $q_{1}$ and $a$ form an orthonormal basis of $L$.
By duality, the polar in $L$ of ${\Delta}^{\circ}\cap L$ is the orthogonal projection of $\Delta$ onto $L$. We compute this projection. First, $v_{1}$ and $v_{n+1}$ already lie in $L$. Second, for each $i\in[n]\setminus\{1\}$,
$P_{L}v_{i}=\left\langle v_{i},q_{1}\right\rangle q_{1}+\left\langle v_{i},a\right\rangle a=t\left\langle q_{1},q_{i}\right\rangle q_{1}+\sqrt{1-t^{2}}\cdot a=-\frac{t}{n-1}q_{1}+\sqrt{1-t^{2}}\cdot a,$
where $P_{L}$ denotes the orthogonal projection onto $L$. In particular, all $v_{2},\dots,v_{n}$ project onto the same point.
Thus, $P_{L}\Delta$ is the triangle with vertices $v_{1}$, $v_{n+1}$, and $P_{L}v_{2}$. Hence, ${\Delta}^{\circ}\cap L$ is a triangle, say $T$. Let us compute its vertices. Since $x_{1}$ and $x_{n+1}$ lie in $L$, they are two of the vertices of $T$. The remaining vertex, say $y$, can be found using duality in $L$:
$y=\frac{1}{n\gamma}\cdot\frac{q_{1}}{t}-\frac{1-n\gamma}{n\gamma}\cdot\frac{a}{\sqrt{1-t^{2}}}.$
Indeed,
$\left\langle y,v_{n+1}\right\rangle=\frac{1-n\gamma}{n\gamma}\cdot\frac{1}{\sqrt{1-t^{2}}}\left\langle a,v_{n+1}\right\rangle=1,$
and
$\left\langle y,v_{1}\right\rangle=\frac{1}{n\gamma}\cdot\left\langle\frac{q_{1}}{t},tq_{1}\right\rangle-\frac{1-n\gamma}{n\gamma}\cdot\left\langle\frac{a}{\sqrt{1-t^{2}}},\sqrt{1-t^{2}}a\right\rangle=\frac{1}{n\gamma}-\frac{1-n\gamma}{n\gamma}=1.$
The lemma follows. ∎
Lemma A.4.
Let $h$ be the foot of the altitude from the vertex $x_{n+1}$ of the triangle $T$, and let $z$ be the intersection point of the altitude $x_{n+1}h$ and the angle bisector at vertex $y$. Then the inradius $R(t)$ of ${\Delta}^{\circ}$ equals $\left|h-z\right|$. Moreover,
$R(t)=\frac{1}{n\gamma}\cdot\frac{1}{1+\sqrt{1-t^{2}}}.$
Proof.
Consider a regular simplex $K$ centered at the origin with a facet $G$ parallel to ${a}^{\bot}$. Let $g$ be the vertex opposite to $G$, and let $v$ be any vertex of $G$. By symmetry, the origin is the incenter of $K$. In particular, the altitude from $v$ passes through the origin and lies in the plane $L=\operatorname{span}\{g,v\}$.
After applying a directional dilation along the direction of $a$ that maps $K$ to a simplex $K^{\prime}$, the incenter of $K^{\prime}$ remains on the line with direction vector $a$. Consequently, the incenter and its orthogonal projection onto the facet opposite to $v$ remain in the plane $L$.
In our setting, this shows that the incenter of ${\Delta}^{\circ}$ lies in $T$ and is the intersection point of the altitude $x_{n+1}h$ and the angle bisector from vertex $y$.
Let us now compute $R(t)$. We use identities from Lemma A.1 and Lemma A.3 for $x_{1}$, $x_{n+1}$, and $y$.
By construction, the line passing through $y$ and $x_{1}$ is orthogonal to $a$, thus
$h=-\frac{1-n\gamma}{n\gamma}\cdot\frac{a}{\sqrt{1-t^{2}}}.$
Hence,
$\left|y-h\right|=\left|\frac{1}{n\gamma}\cdot\frac{q_{1}}{t}\right|=\frac{1}{n\gamma}\cdot\frac{1}{t}$
and
$\left|h-x_{n+1}\right|=\frac{1}{\sqrt{1-t^{2}}}\left(1+\frac{1-n\gamma}{n\gamma}\right)=\frac{1}{n\gamma}\cdot\frac{1}{\sqrt{1-t^{2}}}.$
From the right-angle triangle $x_{n+1}hy$,
$\left|y-x_{n+1}\right|=\sqrt{\left|y-h\right|^{2}+\left|h-x_{n+1}\right|^{2}}=\frac{1}{n\gamma}\cdot\sqrt{\frac{1}{1-t^{2}}+\frac{1}{t^{2}}}.$
Since $yz$ is the angle bisector of triangle $yhx_{n+1}$, we have:
$\frac{\left|h-z\right|}{\left|h-y\right|}=\frac{\left|z-x_{n+1}\right|}{\left|y-x_{n+1}\right|}.$
Therefore, the inradius of the simplex $x_{1},\dots,x_{n+1}$ is
$R(t)=\left|h-z\right|=\frac{\left|y-h\right|}{\left|y-h\right|+\left|y-x_{n+1}\right|}\cdot\left|h-x_{n+1}\right|.$
Substituting the computed values, we obtain:
$R(t)=\frac{1}{n\gamma}\cdot\frac{1}{\sqrt{1-t^{2}}}\cdot\frac{\frac{1}{t}}{\frac{1}{t}+\sqrt{\frac{1}{1-t^{2}}+\frac{1}{t^{2}}}}=\frac{1}{n\gamma}\cdot\frac{1}{\sqrt{1-t^{2}}+\sqrt{t^{2}+1-t^{2}}}=\frac{1}{n\gamma}\cdot\frac{1}{1+\sqrt{1-t^{2}}}.$
Denote $\sigma(t)=\sum\limits_{i\in[n]}\gamma\left|v_{i}\right|^{2}+(1-n\gamma)\left|v_{n+1}\right|^{2}.$ Using the obtained expressions:
$\sigma(t)=n\gamma+\frac{\left(n\gamma\right)^{2}}{1-n\gamma}\left(1-t^{2}\right)=n\gamma\left(1+\frac{n\gamma}{1-n\gamma}\left(1-t^{2}\right)\right).$
Theorem A.5.
Fix a positive integer $d$ such that $d>n\geq 2,$ and set $\gamma=\frac{1}{d}.$ Then, inequality (A.1) holds and the following hold as $t$ tends to $1^{-}$:
  1. (1)
    the left-hand side of inequality (A.2) tends to zero;
  2. (2)
    $\sigma(t)$ is monotonically decreasing to $\frac{n}{d};$
  3. (3)
    $R(t)\to\frac{d}{n}.$
In particular, for any positive integer $m>d-s$ and positive $\varepsilon,$ there is a configuration of $m$ vectors $b_{1},\dots,b_{m}$ of $\mathbb{R}^{d-s}$ and positive numbers $\beta_{1},\dots,\beta_{m}$ satisfying
$\sum\limits_{i\in[m]}\beta_{i}=1,\quad\sum\limits_{i\in[m]}\beta_{i}b_{i}=\mathbf{0},\quad\sum\limits_{i\in[m]}\beta_{i}\left|b_{i}\right|^{2}=\frac{d-s}{d},$
and such that the inradius of the polar of the convex hull of $\{b_{1}\dots,b_{m}\}$ is at least $(1-\varepsilon)\frac{d}{d-s}.$
Proof.
The computations of limits are straightforward.
Set $n=d-s.$ For each $i\in[n],$ set $b_{i}=\sqrt{\frac{\theta}{\sigma(t)}}v_{i}$ with $\theta=\frac{d-s}{d}.$ Set $b_{n+1}=\ldots=b_{m}=\sqrt{\frac{\theta}{\sigma(t)}}v_{n+1}.$ Then,
${\gamma\sum_{i\in[n]}b_{i}+\frac{(1-n\gamma)}{m-n}\sum_{i\in[m]\setminus[n]}b_{i}}=\sqrt{\frac{\theta}{\sigma(t)}}\left(\gamma\sum_{i\in[n]}v_{i}+(1-n\gamma)v_{n+1}\right)=\mathbf{0};$
and
$\gamma\sum_{i\in[n]}\left|b_{i}\right|^{2}+\frac{(1-n\gamma)}{m-n}\sum_{i\in[m]\setminus[n]}\left|b_{i}\right|^{2}=\sigma(t).$
The inradius of the polar of $\operatorname{conv}\{b_{1}\dots,b_{m}\}$ is greater than that of the simplex $v_{1}\dots,v_{d-s+1}$ for $t$ sufficiently close to $1$ On the other hand, there is $t$ such that the latter is at least $(1-\varepsilon)\frac{d}{d-s}.$
Appendix B Majorization inequality for the volume bound
Lemma B.1.
The functional $W(x_{1},\dots,x_{m},\delta_{1},\dots,\delta_{m})=\prod\limits_{i\in[m]}x_{i}^{\delta_{i}x_{i}^{2}}$ under the constraints
  • $\frac{1}{d^{\prime}+1}\leq x_{i}\leq 1\quad\forall\,i\in[m],$
  • $\sum\limits_{i\in[m]}\delta_{i}x_{i}=s+1,$
  • $\sum\limits_{i\in[m]}\delta_{i}=d+1,$
  • $0\leq\delta_{i}\leq 1$
on the variables satisfies the inequality
$W(x_{1},\dots,x_{m},\delta_{1},\dots,\delta_{m})\leq\frac{(s+1)^{\frac{s+1}{2(d+1)}}}{(d+1)^{\frac{d+1}{2(d+1)}}}.$
Proof.
Set $\Lambda=\frac{s+1}{d+1}.$ We will study the logarithm of our functional. We have to show that
$\ln W(x_{1},\dots,x_{m},\delta_{1},\dots,\delta_{m})\leq\frac{1}{2}\Lambda\ln\Lambda-\frac{d-s}{d+1}\ln(d+1).$
Note that the function $\ln W$ is continuous on the set in $\mathbb{R}^{2m}$ defined by our constraints. Hence, it attains its maximum. For every $x=(x_{1},\ldots,x_{m})$ with $\frac{1}{d+1}\leq x_{i}\leq 1$ for all $i\in[m]$, let $\ln W_{x}(\delta_{1},\dots,\delta_{m})$ be the function
$\ln W_{x}(\delta_{1},\dots,\delta_{m})=\frac{1}{2}\sum\limits_{i\in[m]}\delta_{i}x_{i}\ln x_{i}.$
Notice that $\ln W_{x}$ is a convex function of $\delta=(\delta_{1},\dots,\delta_{m}).$ Since the set
[…]
is a compact convex set, $\ln W_{x}$ attains its maximum on some extreme point of $A_{x}$. These are the points of intersection of the 2-dimensional faces of the cube
[…]
with the $(m-2)$-dimensional affine subspace
$\left\{\delta\in\mathbb{R}^{m}:\sum\limits_{i\in[m]}\delta_{i}x_{i}=s+1,\quad\sum\limits_{i\in[m]}\delta_{i}=d+1\right\}.$
Therefore, a maximizer of the function $\ln W_{x}$ has to be a point of the form
[…]
for some $\frac{1}{2}\leq\lambda\leq 1$ (or a permutation of it), such that $\sum\limits_{i\in[m]}\delta_{i}x_{i}=s+1$ is satisfied. For every $\delta_{\lambda}$ with $\frac{1}{2}\leq\lambda\leq 1$, we will find the maximizer of the function
$\ln W_{\delta_{\lambda}}(x)=\frac{1}{2}\sum\limits_{i\in[m]}\delta_{\lambda,i}x_{i}\ln x_{i}$
on the compact convex set
$B_{\lambda}=\left\{x\in\mathbb{R}^{m}:\sum\limits_{i\in[m]}\delta_{\lambda,i}x_{i}=s+1,\quad\frac{1}{d+1}\leq x_{i}\leq 1\quad\forall\,i\in[m]\right\}.$
If $\delta_{\lambda}^{*}$ is the decreasing rearrangement of $\delta_{\lambda}$, we can assume without loss of generality that $\delta_{\lambda}=\delta_{\lambda}^{*}$. Set
$\tilde{x}=\left(\underbrace{1,1,\ldots,1}_{s},\ \Lambda,\ \underbrace{\frac{1}{d+1},\ldots,\frac{1}{d+1}}_{m-s-1}\right).$
We check that $\frac{1}{d+1}\leq\Lambda\leq 1$ and
$\sum\limits_{i\in[m]}\delta_{\lambda,i}\tilde{x}_{i}=s+\Lambda+\frac{d-s}{d+1}=s+1.$
It is easy to see that $\tilde{x}$ majorizes any $x=(x_{1},\dots,x_{m})\in B_{\lambda}.$ Thus, by the weighted Karamata inequality (see [Fuc47]), $\ln W_{\delta_{\lambda}}(x)\leq\ln W_{\delta_{\lambda}}\left(\tilde{x}\right)$ for all $x\in B_{\lambda}.$ By the definition of $\delta_{\lambda},$ we conclude that the maximum under our constraints is equal to
$\max\limits_{\lambda\in[1/2,1]}\ln W_{\delta_{\lambda}}(x)=\max\limits_{\lambda\in[1/2,1]}\ln W_{\delta_{\lambda}}=\frac{1}{2}\Lambda\ln\Lambda-\frac{d-s}{d+1}\ln(d+1).$
The lemma follows. ∎
Acknowledgements
We thank Alexander Litvak for the fruitful conversations. We are also grateful to Susanna Dann, who organized the ’Analysis and Convex Geometry Week at UniAndes’ in Bogotá in 2025, where the initial steps of this work were made.
G. I. is supported by Projeto Paz and by CAPES (Coordenação de Aperfeiçoamento de Pessoal de Nível Superior) - Brasil, grant number 23038.015548/2016-06. Z.L. and M.N. is partially supported by the NRDI research grant K147544. M.N. is also supported by the NRDI research grant K131529 as well as the ELTE TKP 2021-NKTA-62 funding scheme. Á. S. is funded by the grant NKFIH 150151, a project that has been implemented with the support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the ADVANCED_24 funding scheme.