### v2

# 2cm2cm1.5cm1.5cm

Abstract.
Finding a largest Euclidean ball in a given convex body $K\subset\mathbb{R}^{d}$ and finding a largest volume ellipsoid in $K$ are two problems of fundamentally different nature. The first is a purely Euclidean problem, where we consider scaled copies of the origin-centered closed unit ball, whereas in the second problem, we search among all affine copies of the unit ball.
In this paper, we interpolate between these two classical problems by considering ellipsoids of revolution. More generally, we study pairs of convex bodies $K$ and $L$, and seek a largest-volume affine image of $K$ contained within $L$, subject to certain restrictions on the allowed affine transformations. We derive first-order necessary conditions for optimality, generalizing known conditions from the unrestricted affine setting.
Using these conditions, we show that an extremal ellipsoid of revolution exhibits properties analogous to those of either the largest-volume ellipsoid or the largest Euclidean ball, depending on whether the ellipsoid is considered along its axis of revolution or along the orthogonal complement of that axis.
Key words and phrases:
John ellipsoid, body of revolution, convexity, affine image
2020 Mathematics Subject Classification:
Primary 52A23; Secondary 52A40, 46T99
1. Introduction
Finding the largest Euclidean ball contained in a convex body $K\subset\mathbb{R}^{d}$ and finding the largest volume ellipsoid contained in $K$ are problems of fundamentally different nature. The former is a Euclidean problem involving scaled copies of the origin-centered unit ball $\mathbf{B}^{d}$, while the latter involves all affine images of $\mathbf{B}^{d}$. In this paper, we study a natural interpolation between these two settings by considering ellipsoids of revolution. More generally, we investigate pairs of convex bodies $K$ and $L$ and aim to determine the largest volume affine image of $K$ contained in $L$, subject to specific restrictions on the class of admissible affine transformations. Our starting point is John’s classical theorem [Joh48, Joh14] characterizing the maximal volume ellipsoid in a convex body and its generalizations to pairs of convex bodies [Lew79, TJ89, BR02, GLMP04]. Our goal is to initiate the development of a similar theory for ellipsoids of revolution and for constrained affine images in this more general setting.
1.1. Ellipsoids of revolution
In order to introduce ellipsoids of revolution, we start with defining a special class of linear operators.
Definition 1.1.
For a linear subspace $F$ in $\mathbb{R}^{d}$, we call an invertible linear operator $A$ on $\mathbb{R}^{d}$ an $F$-operator, if both $F$ and ${F}^{\bot}$ are invariant subspaces of $A$, and the restriction $A|_{{F}^{\bot}}$ of $A$ on the orthogonal complement ${F}^{\bot}$ of $F$ is a (nonzero) multiple of the identity operator $\mathrm{Id}_{{F}^{\bot}}$ on ${F}^{\bot}$.
Note that if we take an orthonormal basis of $F$, and extend it to an orthonormal basis of $\mathbb{R}^{d}$, then the matrix of an $F$ operator takes the form
$M=\begin{pmatrix}{M_{1}}&\mathbf{0}_{s\times(d-s)}\\ \mathbf{0}_{(d-s)\times s}&\mu\mathrm{Id}_{d-s}\end{pmatrix},$
where $M_{1}\in\mathbb{R}^{s\times s},$ $\mu\in\mathbb{R}\setminus\{0\},$ $\mathbf{0}_{s\times(d-s)}$ and $\mathbf{0}_{(d-s)\times s}$ are zero matrices of the appropriate sizes, and $\mathrm{Id}_{d-s}$ denotes the identity matrix of size $d-s$.
Definition 1.2.
Let $F$ be a linear subspace in $\mathbb{R}^{d}$ and $S$ be any subset of $\mathbb{R}^{d}$. We will refer to the family of affine transformations
$\mathcal{A}_{F}=\left\{x\mapsto Ax+z:\;A\text{ is an }F\text{-operator, }z\in\mathbb{R}^{d}\right\}$
as the affine transformations with axis $F$, and the family of sets
$\mathcal{A}_{F}\!\left(S\right)=\left\{AS+z:\;A\text{ is an }F\text{-operator, }z\in\mathbb{R}^{d}\right\}$
as the positions of $S$ with axis $F$.
In particular, we call members of $\mathcal{A}_{F}\!\left(\mathbf{B}^{d}\right)$ ellipsoids of revolution with axis $F$.
The naming is justified, since if $E\in\mathcal{A}_{F}\!\left(\mathbf{B}^{d}\right)$, then $E$ is an ellipsoid all of whose non-degenerate sections by translates of ${F}^{\bot}$ are $(d-\dim{F})$-dimensional balls. Our first problem is the following.
Problem 1 (John’s problem for ellipsoids with a fixed axis of revolution).
For a given convex body (that is, a closed, convex set with non-empty interior) $K$ in $\mathbb{R}^{d}$ and a linear subspace $F$ in $\mathbb{R}^{d}$, find a largest volume ellipsoid of revolution with axis $F$ contained in $K$.
Clearly, when $F=\{\mathbf{0}\}$ is the 0-dimensional subspace, then this problem is equivalent to finding a largest volume Euclidean ball in $K$. On the other hand, when $F=\mathbb{R}^{d}$, then it is exactly John’s problem of finding the largest volume ellipsoid contained in $K$.
In John’s problem for ellipsoids with a fixed axis of revolution, the subspace $F$ is fixed. It is natural to allow for the variation of $F$, which leads us to our second problem.
Problem 2 (John’s problem for ellipsoids with any axis of revolution).
Let $K$ be a convex body in $\mathbb{R}^{d}$ and $s\in\{0,1,\dots,d\}$. Find a largest volume ellipsoid of revolution in $K$ with any $s$-dimensional axis.
1.2. A more general problem
The problems above are motivated by John’s fundamental theorem [Joh48, Joh14], where a necessary condition is given for the ball $\mathbf{B}^{d}$ to be the maximal volume ellipsoid contained in a convex body $K$. The sufficiency of the same condition was noted by K. Ball [Bal97].
A natural generalization of this question was studied by V. Milman [TJ89, Theorem 14.5], Lewis [Lew79], Giannopoulos, Perissinaki and Tsolomitis [GPT01], Bastero and Romance [BR02], and Gordon, Litvak, Meyer and Pajor [GLMP04]: given convex bodies $K$ and $L$ in $\mathbb{R}^{d}$, find a maximal volume affine image of $K$ that is contained in $L$. In this work, we study the same problem under additional restrictions on the affine transformation.
Problem 3 (John’s general problem with fixed axis).
For given convex bodies $K$ and $L$, and a linear subspace $F$ in $\mathbb{R}^{d}$, find a largest volume member of $\mathcal{A}_{F}\!\left(K\right)$ contained in $L$.
Clearly, when $F=\{\mathbf{0}\}$ is the 0-dimensional subspace, then this problem is equivalent to finding a largest homothet of $K$ in $L$. On the other hand, when $F=\mathbb{R}^{d}$, then it is exactly the problem described in the paragraph preceding Problem 3.
Definition 1.3.
Let $F$ be a linear subspace in $\mathbb{R}^{d}$ and $S$ be any subset of $\mathbb{R}^{d}$. We will refer to the family of affine transformations
$\mathcal{R}_{F}=\left\{x\mapsto RAx+z:\;R\in O(d),A\text{ is an }F\text{-operator, }z\in\mathbb{R}^{d}\right\},$
where $O(d)$ denotes the orthogonal group on $\mathbb{R}^{d}$, as the rotated affine transformations with axis $F$, and the family of sets
$\mathcal{R}_{F}\!\left(S\right)=\left\{RAS+z:\;R\in O(d),A\text{ is an }F\text{-operator, }z\in\mathbb{R}^{d}\right\},$
as the rotated positions of $S$ with axis $F$.
Problem 4 (John’s general problem with rotated axis).
Let $K$ and $L$ be convex bodies in $\mathbb{R}^{d},$ and fix a linear subspace $F$ in $\mathbb{R}^{d}$. Find a largest volume rotated position of $K$ with axis $F$ that is contained in $L$.
Note a subtle but very important difference between this general case and ellipsoids. To define ellipsoids of revolution with an $s$-dimensional axis, we only need to specify the dimension $s$ of the axis, but not the axis $F$. In the general case introduced here, we fix $F$, apply an $F$-operator $A$ and then allow a rotation $R$. The advantage of this is that in this way, we can specify what ‘part’ of $K$ we allow to be changed by an affine transformation (the $F$-component), and what part (the ${F}^{\bot}$ component) is rigid, that is, it only allows similarities. A simple illustration is the task of finding the largest area isosceles triangle contained in a planar set $L$. In our framework, we simply take $K$ to be a regular triangle, and $F$ to be one of its altitude lines. Then $\mathcal{R}_{F}\!\left(K\right)$ is exactly the set of isosceles triangles.
When $F=\{\mathbf{0}\}$, then Problem 4 is the problem discussed at the beginning of this subsection considered by Lewis, Milman and others. On the other hand, when $F=\mathbb{R}^{d}$, then we are looking for the largest similar copy of $K$ in $L$.
1.3. Results
Our first goal is to find conditions of optimality in Problems 1, 2, 3 and 4 similar to John’s condition in the case of ellipsoids and the condition given by Gordon, Litvak, Meyer and Pajor [GLMP04] in the general case. We start with Problems 3 and 4, as the optimality conditions in the case of ellipsoids of revolution will be obtained as corollaries to these general results.
[…]
Our necessary condition of optimality for Problem 3 is the following.
Theorem 1.4 (Condition for John’s general problem with fixed axis).
Let $K,L\subset\mathbb{R}^{d}$ be convex bodies containing the origin in their interiors and such that $K\subseteq L$, and let $F$ be a linear subspace in $\mathbb{R}^{d}$. Assume that $K$ is a solution to John’s general problem with fixed axis $F$ for $K$ and $L$, that is, Problem 3. Then there are
  1. (1)
    contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $L$ and $K$,
  2. (2)
    positive weights $\upalpha_{1},\dots,\upalpha_{m}$,
such that
(1.1) $P_{F}\left(\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\otimes v_{i}\right)P_{F}=P_{F};$
(1.2) $\sum\limits_{i\in[m]}\upalpha_{i}p_{i}=\mathbf{0};$
(1.3) $\operatorname{tr}\left(\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\otimes v_{i}\right)=\sum\limits_{i\in[m]}\upalpha_{i}=d.$
As a corollary, we will obtain the following.
Theorem 1.5 (Containment for John’s general problem with fixed axis).
In addition to the assumptions of Theorem 1.4, suppose that $P_{F}K\subset L\cap F$. Then there exists a translation vector $z$ in the relative interior of $P_{F}K$ such that
$P_{F}(K-z)\subset\left(L-z\right)\cap F\subset-d\cdot P_{F}(K-z).$
In particular, such a translation vector $z$ exists whenever $K$ is symmetric with respect to $F$.
In the case $F=\mathbb{R}^{d}$, Theorem 1.4 and Theorem 1.5 were obtained in [GLMP04].
Our necessary condition of optimality for Problem 4 is the following.
Theorem 1.6 (Condition for John’s general problem with rotated axis).
With the notation of Theorem 1.4, assume that $K$ contains the origin in its interior and is a solution to John’s general problem with rotated axis for $K$, $L$ and the subspace $F$. Then there are contact pairs of $L$ and $K$, and weights as in Theorem 1.4 such that in addition to identities (1.1) – (1.3), we also have
(1.4) $\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\otimes v_{i}=\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}.$
We will state our analogous results for ellipsoids of revolution, that is for Problems 1 and 2 in Section 5. Here we state other geometric properties of maximal volume ellipsoids of revolution.
We define the inradius $\operatorname{inrad}A$ of a compact, convex set $A$ as the radius of a largest radius ball inside the affine hull of $A$ contained in $A$.
Theorem 1.7 (Properties of solutions of John’s problem for ellipsoids with a fixed axis of revolution).
Let an origin-centered ellipsoid $E$ be a solution to Problem 1 for a convex body $K$ and a linear subspace $F$ in $\mathbb{R}^{d}$ of dimension $s\in\{0,1,\dots,d-1\}$. Then
  1. (1)
    the inradius of $K\cap{F}^{\bot}$ satisfies
    $\operatorname{inrad}\left(K\cap{F}^{\bot}\right)\leq\frac{d}{d-s}\cdot\mathrm{radius}(E\cap{F}^{\bot});$
  2. (2)
    the following volume bound holds:
    $\left(\frac{\operatorname{vol}\nolimits_{s}\left(K\cap F\right)}{\operatorname{vol}\nolimits_{s}\left(E\cap F\right)}\right)^{\frac{1}{s}}\leq\frac{1}{(s+1)^{\frac{d-s}{2s(d+1)}}}\sqrt{\frac{d(d+1)}{s(s+1)}}\left(\frac{\operatorname{vol}\nolimits_{s}\Delta^{s}_{J}}{\operatorname{vol}\nolimits_{s}\mathbf{B}^{s}}\right)^{\frac{1}{s}},$
    where $\Delta_{J}^{s}$ stands for the $s$-dimensional regular simplex circumscribed around $\mathbf{B}^{s}$;
  3. (3)
    the following inclusion holds:
    $E\cap F\subset K\cap F\subset d\left(E\cap F\right).$
If additionally, $K$ is centrally symmetric, then
  1. (S1)
    the inradius of $K\cap{F}^{\bot}$ satisfies
    $\operatorname{inrad}\left(K\cap{F}^{\bot}\right)\leq\sqrt{\frac{d}{d-s}}\cdot\mathrm{radius}(E\cap{F}^{\bot});$
  2. (S2)
    the following volume bound holds:
    $\left(\frac{\operatorname{vol}\nolimits_{s}\left(K\cap F\right)}{\operatorname{vol}\nolimits_{s}\left(E\cap F\right)}\right)^{\frac{1}{s}}\leq\sqrt{\frac{d}{s}}\left(\frac{\operatorname{vol}\nolimits_{s}\Box^{s}}{\operatorname{vol}\nolimits_{s}\mathbf{B}^{s}}\right)^{\frac{1}{s}},$
    where $\Box^{s}$ stands for the standard cube $[-1,1]^{s}$;
  3. (S3)
    the following inclusion holds:
    $E\cap F\subset K\cap F\subset\sqrt{d}\left(E\cap F\right).$
Note that (3) and (S3) are sharp: $\mathbf{B}^{d}$ is the John ellipsoid of $\Delta^{d}$ (resp., of $\Box^{d}$), and hence, it is the solution of John’s problem for ellipsoids with a fixed axis of revolution for any $F$. If we now choose $F$ properly, then we have equality in these two inequalities.
On the other hand, (2) and (S2) are identical to the best bounds that are known for sections of $K$ and its (unrestricted) John ellipsoid. See our discussion of optimality in Subsection 5.5.
There is hardly any difference between the proofs of (2), (3) (and (S2), (S3)) and the corresponding statements for John ellipsoids. Everything reduces to the algebraic conditions on the projections of the contact vectors onto $F$. The proof of (1) (and (S1)) however, is more intricate.
We prove Theorems 1.4 and 1.6 in Sections 2 and 3. In Subsection 3.3, we state a certain converse of Theorems 1.4 and 1.6, that is, that the equations in this theorem ensure that $K$ is of locally maximum volume among certain affine images of $K$ under some additional conditions.
In Section 4, we consider the problem of finding a homothety ratio such that enlarging the inner body in some sort of John position by this ratio, it will contain the outer body. In particular, we prove (3) in Theorem 1.7, and also show that no similar containment holds in the whole space $\mathbb{R}^{d}$.
We may refer to the notion of an isosceles triangle as a ‘semi-Euclidean’ notion, since it is neither an affine notion nor purely Euclidean – that is, a notion invariant under similarity alone. In Subsection 4.4, we show an application of our results to isosceles triangles, and more generally, to right cones to illustrate how our language is capable of addressing semi-Euclidean questions.
In Section 5, we discuss the problem of finding largest volume ellipsoids of revolution in a convex body, and in particular, we prove Theorem 1.7. We discuss the uniqueness of the solution to these John type problems in Section 6, and finally, in Section 7, we outline how some of the results change for the dual question (often referred to as Löwner’s ellipsoid), where the smallest volume ellipsoid containing a convex body is considered. In Appendix A, we show that (1) of Theorem 1.7 is tight. Appendix B contains the proof of a technical fact that we use in Section 5.
1.4. Notation
We denote by $\mathbf{0}$ the origin of a vector space, and sometimes in a subscript, we specify the space, e.g. $\mathbf{0}_{\mathcal{L}}$ in a space $\mathcal{L}$ of matrices. We use $\left\langle u,v\right\rangle$ to denote the standard inner product of the vectors $u,v\in\mathbb{R}^{d}$, and $\mathbf{B}^{d}$ for the closed Euclidean unit ball in $\mathbb{R}^{d}$ centered at the origin. We denote the positive hull of a set $S$ in $\mathbb{R}^{n}$ by $\operatorname{pos}\left(S\right)=\left\{\lambda x:\;\lambda\geq 0,x\in S\right\}.$ The polar of a set $S$ is defined as $S^{\circ}=\left\{p\in\mathbb{R}^{n}:\;\left\langle x,p\right\rangle\leq 1\,\text{ for all }x\in S\right\}$. Note that for a positive cone $C$ (that is, the positive hull of some set), we have $C^{\circ}=\left\{p\in\mathbb{R}^{n}:\;\left\langle x,p\right\rangle\leq 0\,\text{ for all }x\in C\right\}$. The group of orthogonal transformations on $\mathbb{R}^{d}$ is $O(d)$.
Recall that the polar decomposition of a square matrix $A\in\mathbb{R}^{d\times d}$ is a factorization of $A$ in the form $A=RM,$ where $R$ is an orthogonal matrix and $M$ is positive semi-definite. Moreover, if $A$ is invertible, then this representation is unique ($M=\sqrt{{A}^{\top}A}$ and $U=A{M}^{-1}$), where ${A}^{\top}$ denotes the transpose of $A$.
2. First-order necessary conditions
In this section, we reformulate Theorems 1.4 and 1.6 in terms of the (non-)existence of a certain curve in a vector space. The difference between the proofs of the two theorems lies in the choice of the vector space in which the curve (non-)exists. This condition is then translated into a statement about the separation of a point and a convex cone in the corresponding space. This is a standard technique in proving John-type theorems; cf. [Bal97].
First, we recall basic facts describing convex sets locally, and next, we introduce the vector space which parametrizes local affine transformations. In the final part of this section, we rephrase the existence of a curve in this vector space along which an ‘improved’ position of $K$ in $L$ can be found in terms of separation of cones.
2.1. The normal and the tangent cone
Definition 2.1.
Let $C$ be a convex body in $\mathbb{R}^{n}$ and $a\in\operatorname{bd}C$ be a boundary point. The normal cone $N_{K}\!\left(a\right)$ is the set consisting of the origin and all the outward normal vectors to $C$ at $a$. That is,
$N_{C}\!\left(a\right)=\left\{p\in\mathbb{R}^{n}:\;\left\langle p,z-a\right\rangle\leq 0\quad\forall z\in C\right\}.$
The tangent cone $\mathrm{T}_{C}\!\left(a\right)$ at $a$ is the closure of the cone $\mathrm{Pos}\ \!{\left(C-a\right)}$.
The following facts are well-known, and follow from standard separation properties of convex sets [HUL04, Section 5.2]:
Lemma 2.2.
Consider a convex set $C$ in $\mathbb{R}^{n}$ and a point $a\in\operatorname{cl}\left(C\right)$. Then
${\mathrm{T}_{C}\!\left(a\right)}={\left(N_{C}\!\left(a\right)\right)}^{\circ}\quad\text{and}\quad{\left(\mathrm{T}_{C}\!\left(a\right)\right)}^{\circ}={N_{C}\!\left(a\right)}.$
In particular, $h\in\mathrm{int}\mathrm{T}_{C}\!\left(a\right)$ if and only if
$\left\langle h,p\right\rangle<0$
for all non-zero $p\in N_{C}\!\left(a\right)$.
Lemma 2.3.
Let $C$ be a convex body in $\mathbb{R}^{n}$, and $a\in C.$ Let $\gamma_{t}$ where $t\in[0,1]$ be a curve in $\mathbb{R}^{n}$ with $\gamma_{0}=a$ and
$\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}\gamma_{t}=h.$
Consider the following assertions:
  1. (i)
    $h\in\mathrm{int}\mathrm{T}_{C}\!\left(a\right);$
  2. (ii)
    $\gamma_{t}\in C$ for all $t\in[0,\tau]$ for some $\tau>0;$
  3. (iii)
    $h\in\mathrm{T}_{C}\!\left(a\right)$.
Then (i) implies (ii), and (ii) implies (iii).
2.2. Operator spaces
We use $\mathcal{M}^{d}$ and $\mathcal{M}_{\mathrm{Sym}}^{d}$ to denote the space of operators and the space of self-adjoint operators from $\mathbb{R}^{d}$ to $\mathbb{R}^{d}$, respectively. The inner product on $\mathcal{M}^{d}$ is defined as $\left\langle A,B\right\rangle_{\mathcal{M}^{d}}=\operatorname{tr}\left(A{B}^{\top}\right)$.
Define
=$\mathcal{R}=\left\{(M,S,g):\;M,S\in\mathcal{M}^{d},\ g\in\mathbb{R}^{d}\right\}.$
We endow $\mathcal{R}$ with an inner product as follows:
$\left\langle(M_{1},S_{1},g_{1}),(M_{2},S_{2},g_{2})\right\rangle_{\mathcal{R}}=\left\langle M_{1},M_{2}\right\rangle_{\mathcal{M}^{d}}+\left\langle S_{1},S_{2}\right\rangle_{\mathcal{M}^{d}}+\left\langle g_{1},g_{2}\right\rangle.$
2.3. A curve in $\mathcal{R}$
First, we describe certain one parameter families of affine transformations of $\mathbb{R}^{d}$. More specifically, we consider a differentiable path $\gamma:[0,t_{0})\longrightarrow\mathcal{R}$ for some $t_{0}>0$ defined by
(2.1) $\gamma_{t}=(\mathrm{Id}_{\mathbb{R}^{d}}+tH,R_{t},h_{t}),$
where $h_{0}=\mathbf{0}$, $R_{t}\in O(d)$ and $R_{0}=\mathrm{Id}_{\mathbb{R}^{d}}$.
It is well-known [Hal13, Proposition 3.24.] that by setting
(2.2) $W=\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}R_{t},$
we have that $W$ is a skew-symmetric matrix, that is, $W=-{W}^{\top}$, and conversely, for every skew-symmetric matrix $W$, there is a path $R_{t}$ in the orthogonal group satisfying (2.2).
Denoting $\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}h_{t}$ by $z$, we have
(2.3) $\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}\gamma_{t}=(H,W,z).$
For any $t$, we will regard $\gamma_{t}$ as a representation of the affine transformation
(2.4) $A_{t}=R_{t}(\mathrm{Id}_{\mathbb{R}^{d}}+tH)+h_{t}$
on $\mathbb{R}^{d}$.
Throughout the section, $R_{t},A_{t},h_{t},H,W,z$ denote the quantities defined so far, all derived from a differentiable path $\gamma_{t}$ described in (2.1).
Let $K$ and $L$ be convex bodies in $\mathbb{R}^{d}$ with $K\subset L$. We are interested in two questions: first, how does the volume of $K$ change when we replace $K$ by $A_{t}(K)$? Second, under what conditions is $A_{t}(K)$ contained in $L$?
The first question is easy. We differentiate the determinant of the linear operator part of $A_{t}$:
(2.5) $\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}\det(R_{t}(\mathrm{Id}_{\mathbb{R}^{d}}+tH))=\operatorname{tr}H=$
$\left\langle\left(\mathrm{Id}_{\mathbb{R}^{d}},\mathbf{0}_{\mathcal{M}^{d}},\mathbf{0}\right),\left(H,W,z\right)\right\rangle_{\mathcal{R}}=\left\langle\left(\mathrm{Id}_{\mathbb{R}^{d}},\mathbf{0}_{\mathcal{M}^{d}},\mathbf{0}\right),\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}\gamma_{t}\right\rangle_{\mathcal{R}}.$
Setting
$\mathcal{V}_{\mathcal{R}}=\mathrm{Pos}\ \!{\left(\mathrm{Id}_{\mathbb{R}^{d}},\mathbf{0}_{\mathcal{M}^{d}},\mathbf{0}\right)},$
which is a ray in $\mathcal{R}$, we conclude the following.
Lemma 2.4.
If
$\left(H,W,z\right)\in-\mathrm{int}\left({\left(\mathcal{V}_{\mathcal{R}}\right)}^{\circ}\right),$
then the volume of $A_{t}(K)$ is strictly increasing for sufficiently small $t>0$.
Second, Lemmas 2.2 and 2.3 provide a nice tool for characterization of local inclusions along curves.
Lemma 2.5.
Let $L$ be a convex body in $\mathbb{R}^{d}$ containing the origin in its interior, and let $v$ be a boundary point of $L$. Define the perturbation $v_{t}$ of $v$ by
$v_{t}=A_{t}(v).$
Consider the following statements:
  1. (1)
    $\left\langle(H,W,z),(p\otimes v,p\otimes v,p)\right\rangle_{\mathcal{R}}<0;$
    for all $p\in{L}^{\circ}$ with $\left\langle p,v\right\rangle=1$.
  2. (2)
    $v_{t}\in L$ for all $t\in(0,\tau]$ and some $\tau>0$.
  3. (3)
    $\left\langle(H,W,z),(p\otimes v,p\otimes v,p)\right\rangle_{\mathcal{R}}\leq 0;$
    for all $p\in{L}^{\circ}$ with $\left\langle p,v\right\rangle=1$.
Then (1) implies (2), and (2) implies (3).
Proof.
By direct computation,
$\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}v_{t}=\left(H+W\right)v+z.$
Thus, for any $p\in\mathbb{R}^{d}$, we have
$\left\langle\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}v_{t},p\right\rangle=\left\langle\left(H+W\right)v+z,p\right\rangle=\left\langle Hv,p\right\rangle+\left\langle Wv,p\right\rangle+\left\langle z,p\right\rangle=$
$\left\langle(H,W,z),(p\otimes v,p\otimes v,p)\right\rangle_{\mathcal{R}}.$
The result follows from the descriptions of the tangent cone in Lemma 2.2 and Lemma 2.3. ∎
We denote the set of contact points of $K$ and $L$ by
$\mathrm{Cont}_{\mathbb{R}^{d}}=\operatorname{bd}K\cap\operatorname{bd}L,$
and, from this point onward, assume that $K$ contains the origin in its interior.
Corollary 2.6.
If
(2.6) $\left\langle(H,W,z),(p\otimes v,p\otimes v,p)\right\rangle_{\mathcal{R}}<0$
for all $v\in\mathrm{Cont}_{\mathbb{R}^{d}}$ and for all $p\in{L}^{\circ}$ with $\left\langle p,v\right\rangle=1$, then for a sufficiently small $\varepsilon>0$, the set $A_{t}K+h_{t}$ is contained in $\mathrm{int}L$ for all $t\in(0,\varepsilon)$.
Proof.
By Lemma 2.5, for any $v\in\mathrm{Cont}_{\mathbb{R}^{d}}$, $A_{t}(v)$ is in $\mathrm{int}K$ for all $t\in(0,\varepsilon_{v})$ with some $\varepsilon_{v}>0$ depending on $v$. A standard compactness argument completes the proof. ∎
Define the set of contact operators by
$\mathrm{Cont}_{\mathcal{R}}=\left\{\left(p\otimes v,p\otimes v,p\right):\;v\in\mathrm{Cont}_{\mathbb{R}^{d}},p\in{L}^{\circ},\left\langle v,p\right\rangle=1\right\},$
and the contact operator cone by
$\mathcal{C}_{\mathcal{R}}=\mathrm{Pos}\ \!{\mathrm{Cont}_{\mathcal{R}}}.$
With this notation, Corollary 2.6 can be restated as follows.
Corollary 2.7.
If
(2.7) $(H,W,z)\in\mathrm{int}\left({\mathcal{C}_{\mathcal{R}}}^{\circ}\right),$
then for a sufficiently small $\varepsilon>0$, the set $A_{t}K+h_{t}$ is contained in $\mathrm{int}L$ for all $t\in(0,\varepsilon)$.
Definition 2.8.
Let $K\subseteq L$ be convex bodies in $\mathbb{R}^{d}$. We say that $K$ is of locally maximum volume among elements of $\mathcal{R}_{F}\!\left(K\right)$ (resp., $\mathcal{A}_{F}\!\left(K\right)$) contained in $L$, if there is no $t_{0}>0$ and differentiable curve $A_{t}:[0,t_{0})\longrightarrow\mathcal{R}_{F}$ (resp., $A_{t}:[0,t_{0})\longrightarrow\mathcal{A}_{F}$) such that $A_{0}=\mathrm{Id}_{\mathbb{R}^{d}}$, $\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}\operatorname{vol}\nolimits_{d}A_{t}(K)>0$ and $A_{t}(K)\subset\mathrm{int}L$ for all $t\in(0,t_{0})$.
For some fixed subspace $F$ of $\mathbb{R}^{d}$, we set
$\mathcal{D}=\bigg{\{}(H,W,z)\in\mathcal{R}:\;H\ \text{ is an }F\text{-operator, }W=-{W}^{\top},\ z\in\mathbb{R}^{d}\bigg{\}}.$
Lemma 2.9 (Local optimality in $\mathcal{R}_{F}\!\left(K\right)$).
If $\mathbf{0}\in\mathrm{int}K$ and $K$ is of locally maximum volume among elements of $\mathcal{R}_{F}\!\left(K\right)$ contained in $L$, then
$\left(\mathcal{V}_{\mathcal{R}}+{\mathcal{D}}^{\bot}\right)\cap\mathcal{C}_{\mathcal{R}}\neq\{\mathbf{0}_{\mathcal{R}}\}.$
Proof.
By (2.2) and (2.3), we have that if $(H,W,z)$ is in $\mathcal{D}$, then there is a curve $\gamma_{t}$ in $\mathcal{R}$ with $\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}\gamma_{t}=(H,W,z)$, and $\gamma_{0}=(\mathrm{Id}_{\mathbb{R}^{d}},\mathbf{0}_{\mathcal{M}^{d}},\mathbf{0})$ such that $A_{t}$ defined by (2.4) is a perturbation of $K$ with $A_{t}(K)\in\mathcal{R}_{F}\!\left(K\right).$
Combining this observation with Lemma 2.4 and Corollary 2.7, we obtain the following.
If $K\subseteq L$ are convex bodies in $\mathbb{R}^{d}$ with $\mathbf{0}\in\mathrm{int}L$, and
$\mathcal{D}\cap\mathrm{int}\left({\mathcal{C}_{\mathcal{R}}}^{\circ}\right)\cap\mathrm{int}\left(-{\left(\mathcal{V}_{\mathcal{R}}\right)}^{\circ}\right)\neq\{\mathbf{0}_{\mathcal{R}}\},$
then there is a curve $\gamma$ showing that $K$ is not of locally maximum volume among elements of $\mathcal{R}_{F}\!\left(K\right)$ contained in $L$.
Equivalently, if $K$ is of locally maximum volume among elements of $\mathcal{R}_{F}\!\left(K\right)$ contained in $L$, then
(2.8) $\mathcal{D}\cap\mathrm{int}\left(-{\left(\mathcal{V}_{\mathcal{R}}\right)}^{\circ}\right)\cap\mathrm{int}\left({\mathcal{C}_{\mathcal{R}}}^{\circ}\right)=\{\mathbf{0}_{\mathcal{R}}\}.$
Notice that $\mathcal{D}$ is a linear subspace in $\mathcal{R}$, and the other two sets on the left-hand side are open cones in $\mathcal{R}$. By basic separation properties of cones in Euclidean space, the proof of Lemma 2.9 is complete. ∎
What changes if we are not allowed to use a rotation in the setting of Theorem 1.4? In this case, the curve $A_{t}$ is in the subspace $\mathcal{A}_{F}$, or, equivalently, $\gamma_{t}$ is such that $R_{t}=\mathrm{Id}_{\mathbb{R}^{d}}$ for all $t$, which is equivalent to the condition that $\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}\gamma_{t}$ takes all values in
$\mathcal{D}_{0}=\bigg{\{}(H,\mathbf{0}_{\mathcal{M}^{d}},z)\in\mathcal{R}:\;H\ \text{ is an }F\text{-operator, }\ z\in\mathbb{R}^{d}\bigg{\}}.$
Thus, (2.8) changes to the following.
If $K$ is of locally maximum volume among elements of $\mathcal{A}_{F}\!\left(K\right)$ contained in $L$, then
$\mathcal{D}_{0}\cap\mathrm{int}\left(-{\left(\mathcal{V}_{\mathcal{R}}\right)}^{\circ}\right)\cap\mathrm{int}\left({\mathcal{C}_{\mathcal{R}}}^{\circ}\right)=\{\mathbf{0}_{\mathcal{R}}\}.$
This yields the following.
Lemma 2.10 (Local optimality in $\mathcal{A}_{F}\!\left(K\right)$).
If $\mathbf{0}\in\mathrm{int}K$ and $K$ is of locally maximum volume among elements of $\mathcal{A}_{F}\!\left(K\right)$ contained in $L$, then
$\left(\mathcal{V}_{\mathcal{R}}+{\mathcal{D}_{0}}^{\bot}\right)\cap\mathcal{C}_{\mathcal{R}}\neq\{\mathbf{0}_{\mathcal{R}}\}.$
In the next two sections, we prove Theorems 1.6 and 1.4, using Lemmas 2.9 and 2.10, respectively.
3. Proof of Theorems 1.6 and 1.4
We fix a non-trivial subspace $F$ of $\mathbb{R}^{d}$ of dimension $s.$ We use $P_{F}$ to denote the orthogonal projection onto $F$. Set
$\mathcal{Q}_{F}=\left\{Q\in\operatorname{Lin}\left(\mathbb{R}^{d}\right):\;P_{F}QP_{F}=\mathbf{0}_{F},\operatorname{rank}\left(P_{{F}^{\bot}}QP_{{F}^{\bot}}\right)=d-s,\operatorname{tr}\left(P_{{F}^{\bot}}QP_{{F}^{\bot}}\right)=0\right\},$
where $\mathbf{0}_{F}$ is the zero operator on $F$. In coordinates, if we consider an orthonormal basis of $F$ and extend it to an orthonormal basis of $\mathbb{R}^{d}$, then the matrix of $Q\in\mathcal{Q}_{F}$ in this basis is of the form
$Q=\begin{pmatrix}\mathbf{0}_{s\times s}&Q_{1}\\ Q_{2}&Q_{3}\end{pmatrix},$
where $\mathbf{0}_{s\times s}$ is the $s\times s$ zero matrix, $Q_{1}$ and ${Q_{2}}^{\top}$ are $s\times(d-s)$ matrices, and $Q_{3}$ is a full rank $(d-s)\times(d-s)$ matrix of zero trace.
Lemma 3.1.
The following identities hold:
${\mathcal{D}}^{\bot}=\mathcal{Q}_{F}\times\mathcal{M}_{\mathrm{Sym}}^{d}\times\{\mathbf{0}\},\quad\text{and}\quad{\mathcal{D}_{0}}^{\bot}=\mathcal{Q}_{F}\times\mathcal{M}^{d}\times\{\mathbf{0}\}.$
Proof.
Let $\left(M,S,g\right)\in\mathcal{R}$. If $\left(M,S,g\right)\in{\mathcal{D}}^{\bot}$, then
$\left\langle\left(M,S,g\right),\left(\mathbf{0}_{\mathcal{M}^{d}},\mathbf{0}_{\mathcal{M}^{d}},z\right)\right\rangle_{\mathcal{R}}=\left\langle g,z\right\rangle=0$
for all $z\in\mathbb{R}^{d}$. Hence, $g=\mathbf{0}$.
The orthogonal complement of the space of skew-symmetric matrices in $\mathcal{M}^{d}$ is $\mathcal{M}_{\mathrm{Sym}}^{d}$. Thus, $S\in\mathcal{M}_{\mathrm{Sym}}^{d}$.
Finally, let $M$ and an $F$-matrix $H$ be of the form:
$H=\begin{pmatrix}{H_{1}}&\mathbf{0}_{s\times(d-s)}\\ \mathbf{0}_{(d-s)\times s}&\mu\mathrm{Id}_{d-s}\end{pmatrix}\quad\text{and}\quad M=\begin{pmatrix}{M_{1}}&{M_{2}}\\ {M_{3}}&{M_{4}}\end{pmatrix}$
for some non-zero constant $\mu$, where $H_{1}$ and $M_{1}$ are $s\times s$ matrices, $M_{4}$ is a $(d-s)\times(d-s)$ matrix, and $M_{2}$ and ${M_{3}}^{\top}$ are $s\times(d-s)$ matrices.
Then
$\left\langle\left(M,S,\mathbf{0}\right),\left(H,\mathbf{0}_{\mathcal{M}^{d}},z\right)\right\rangle_{\mathcal{R}}=\operatorname{tr}\left(M{H}^{\top}\right)=\operatorname{tr}\left(M_{1}{H_{1}}^{\top}\right)+\mu\operatorname{tr}M_{4}.$
As $H_{1}$ and $\mu$ can be chosen arbitrarily,
$\left\langle\left(M,S,\mathbf{0}\right),\left(H,\mathbf{0}_{\mathcal{M}^{d}},z\right)\right\rangle_{\mathcal{R}}=0$
if and only if $M_{1}=\mathbf{0}_{s\times s}$ and $\operatorname{tr}M_{4}=0$, while there are no constraints on $M_{2}$ and $M_{3}$. This is equivalent to $M\in\mathcal{Q}_{F}$.
Thus,
${\mathcal{D}}^{\bot}=\mathcal{Q}_{F}\times\mathcal{M}_{\mathrm{Sym}}^{d}\times\{\mathbf{0}\}.$
The second statement of the lemma follows analogously. ∎
3.1. Proofs of Theorem 1.6 and 1.4
Proof of Theorem 1.6.
Lemmas 2.9 and 3.1 yield that
$\left(\mathrm{Pos}\ \!{\{}(\mathrm{Id}_{\mathbb{R}^{d}},\mathbf{0}_{\mathcal{M}^{d}},\mathbf{0})\}+\mathcal{Q}_{F}\times\mathcal{M}_{\mathrm{Sym}}^{d}\times\{\mathbf{0}\}\right)\cap\mathcal{C}_{\mathcal{R}}\neq\{\mathbf{0}_{\mathcal{R}}\}.$
Thus, there are matrices $Q\in\mathcal{Q}_{F}$ and $S\in\mathcal{M}_{\mathrm{Sym}}^{d}$ such that $\left(\mathrm{Id}_{\mathbb{R}^{d}}+Q,S,\mathbf{0}\right)\in\mathcal{C}_{\mathcal{R}}$. That is, there are contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $K$ and $L,$ and weights $\upalpha_{1},\dots,\upalpha_{m}>0$ such that $\sum\limits_{i\in[m]}\upalpha_{i}p_{i}=\mathbf{0}$ and the matrix $M=\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\otimes v_{i}$ satisfies the conditions
  • $P_{F}MP_{F}=P_{F}(\mathrm{Id}_{\mathbb{R}^{d}}+Q)P_{F}=P_{F}\mathrm{Id}_{\mathbb{R}^{d}}P_{F}=P_{F}$,
  • $M={M}^{\top}$,
  • $\operatorname{tr}M=\operatorname{tr}\left(\mathrm{Id}_{\mathbb{R}^{d}}+Q\right)=\operatorname{tr}\mathrm{Id}_{\mathbb{R}^{d}}=d.$
The proof of Theorem 1.6 is complete. ∎
The proof of Theorem 1.4 is essentially identical to the previous one.
Proof of Theorem 1.4.
Lemmas 2.10 and 3.1 yield that
$\left(\mathrm{Pos}\ \!{\{}(\mathrm{Id}_{\mathbb{R}^{d}},\mathbf{0}_{\mathcal{M}^{d}},\mathbf{0})\}+\mathcal{Q}_{F}\times\mathcal{M}^{d}\times\{\mathbf{0}\}\right)\cap\mathcal{C}_{\mathcal{R}}\neq\{\mathbf{0}_{\mathcal{R}}\}.$
Thus, there are matrices $Q\in\mathcal{Q}_{F}$ and $S\in\mathcal{M}^{d}$ such that $\left(\mathrm{Id}_{\mathbb{R}^{d}}+Q,S,\mathbf{0}\right)\in\mathcal{C}_{\mathcal{R}}$. That is, there are contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $K$ and $L,$ and weights $\upalpha_{1},\dots,\upalpha_{m}>0$ such that $\sum\limits_{i\in[m]}\upalpha_{i}p_{i}=\mathbf{0}$ and the matrix $M=\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\otimes v_{i}$ satisfies the conditions
  • $P_{F}MP_{F}=P_{F}(\mathrm{Id}_{\mathbb{R}^{d}}+Q)P_{F}=P_{F}\mathrm{Id}_{\mathbb{R}^{d}}P_{F}=P_{F}$, and
  • $\operatorname{tr}M=\operatorname{tr}\left(\mathrm{Id}_{\mathbb{R}^{d}}+Q\right)=\operatorname{tr}\mathrm{Id}_{\mathbb{R}^{d}}=d.$
The proof of Theorem 1.4 is complete. ∎
3.2. Rotations
This subsection is a small detour in which we clarify the geometry hidden behind the symmetry condition (1.4). The proofs of our other results do not rely on it.
Theorem 3.2.
Let $K$ and $L$ be convex bodies in $\mathbb{R}^{d}$ such that the origin is in the interior of $L\cap K,$ and there are no contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $L$ and $K$ and corresponding positive weights $\upalpha_{1},\dots,\upalpha_{m}$ satisfying identity (1.4). Then for some $\varepsilon>0$, there is a differentiable path $R_{t},$ $t\in(0,\varepsilon]$ of orthogonal transformations such that $R_{t}K\subset\mathrm{int}L$ for all $t\in(0,\varepsilon].$
Proof.
There is nothing to prove if the set of contact pairs is empty. We assume that the cone $\operatorname{Cone}_{\mathrm{op}}=\mathrm{Pos}\ \!{\left\{p\otimes v:\;(v,p)\text{ is a contact pair of }K\text{ and }L\right\}}$ is non-empty. Clearly, it is a closed set.
Then the negation of identity (1.4) is equivalent to
$\mathcal{M}_{\mathrm{Sym}}^{d}\cap\operatorname{Cone}_{\mathrm{op}}=\left\{\mathbf{0}_{\mathcal{M}^{d}}\right\}.$
Thus, there is a non-zero matrix $W$ separating the closed convex cone $\operatorname{Cone}_{\mathrm{op}}$ from $\mathcal{M}_{\mathrm{Sym}}^{d}$ in the following sense: for every nonzero $M_{1}\in\mathcal{M}_{\mathrm{Sym}}^{d}$ and $M_{2}\in\operatorname{Cone}_{\mathrm{op}}$, we have $\left\langle M_{1},W\right\rangle_{\mathcal{M}^{d}}\leq 0<\left\langle M_{2},W\right\rangle_{\mathcal{M}^{d}}$. Since $\mathcal{M}_{\mathrm{Sym}}^{d}$ is a subspace, $W\in{\left(\mathcal{M}_{\mathrm{Sym}}^{d}\right)}^{\bot},$ that is, $W$ is skew-symmetric.
Set $R_{t}=e^{tW}$ and $H=0$ in (2.1). Clearly, $R_{t},t\in[0,1]$ is a differentiable path of orthogonal matrices in $\mathcal{M}^{d}$ satisfying (2.2). By direct computation,
$\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}\left(R_{t}v\right)=Wv.$
The result follows from the descriptions of the tangent cone in Lemma 2.2 and Lemma 2.3, and standard compactness arguments.
3.3. The condition is sufficient for local extremality – in some cases
According to Theorem 1.4, the local maximality of the volume of $K$ implies certain relations on contact pairs of $K$ and $L$.
The converse is not true in general. If we consider a regular simplex centered at the origin as $K$, and $L=-d\cdot K$, then there are contact pairs satisfying equations (1.1)–(1.4), and yet, clearly, $K$ is not a locally largest volume $F$-position of $K$ in $L$, for $F=\mathbb{R}^{d}$.
However, as in the classical setting, the log-concavity of the determinant (Minkowski’s determinant inequality) yields the sufficiency of our conditions for positions $AK+a$ with positive definite $A$. Denote by
$\mathcal{A}_{F}^{+}(S)=\left\{AS+z:\;A\text{ is a positive definite }F\text{-operator, }z\in\mathbb{R}^{d}\right\},$
the positive $F$-positions of $S$.
Theorem 3.3.
Let $K$ and $L$ be convex bodies in $\mathbb{R}^{d}$ and $F$ a linear subspace in $\mathbb{R}^{d}$. Assume that $K\subseteq L$ and there are contact points of $K$ and $L$ and weights satisfying (1.1), (1.2) and (1.3). Then $K$ is of locally maximum volume among elements of $\mathcal{A}_{F}^{+}(K)$ contained in $L$.
We leave the proof to the reader.
4. Containment
Let us recall how the authors of [GLMP04] derived Theorem 1.5 from Theorem 1.4 in the case $F=\mathbb{R}^{d}$. Note that the identities (1.1), (1.2), and (1.3) in Theorem 1.4 can be rewritten with respect to an arbitrary point in the interior of $K$; in other words, we may choose the origin arbitrarily within the interior of $K$.
In [GLMP04] it is shown that in the case $F=\mathbb{R}^{d}$, one can choose the origin so that, in addition to the identities (1.1), (1.2), and (1.3), the following identity also holds
$\sum\limits_{i\in[m]}\upalpha_{i}v_{i}=0.$
As we will show, in our setting, the following relation suffices to prove the containment in Theorem 1.5:
(4.1) $\sum\limits_{i\in[m]}\upalpha_{i}P_{F}v_{i}=\mathbf{0}.$
4.1. Inclusion for contact polytopes
Let $K$ and $L$ be convex bodies in $\mathbb{R}^{d}$ such that $K\subset L$ and $K$ contains the origin in its interior. Let $F$ be a non-trivial subspace of $\mathbb{R}^{d}.$
Assume that there are contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $K$ and $L,$ and weights $\upalpha_{1},\dots,\upalpha_{m}$ such that identities (1.1) – (1.3) hold, and moreover, identity (4.1) holds.
Denote $K_{\text{in}}=\operatorname{conv}\left\{v_{1},\dots,v_{m}\right\}$ and $L_{\text{out}}=\bigcap\limits_{i\in[m]}\left\{x\in\mathbb{R}^{d}:\;\left\langle p_{i},x\right\rangle\leq 1\right\}.$
We call $K_{\text{in}}$ and $L_{\text{out}}$ the inner and outer contact polytope, respectively, corresponding to the contact pairs $(v_{i},p_{i})$. Clearly,
$K_{\text{in}}\subset K\subset L\subset L_{\text{out}}.$
Lemma 4.1.
With the above notation and assumptions, we have
$L_{\mathrm{out}}\cap F\subset-d\cdot P_{F}K_{\mathrm{in}}.$
Proof.
Equation (1.1) is equivalent to
(4.2) $P_{F}{\left(\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}\right)}P_{F}=P_{F}{\left(\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\otimes v_{i}\right)}^{\top}P_{F}=P_{F}.$
Let $x$ be a point in $L_{\text{out}}\cap F$, then $\left\langle p_{i},x\right\rangle\leq 1$ for all $i\in[m].$
[…]
$\sum\limits_{i\in[m]}\upalpha_{i}(1-\left\langle p_{i},x\right\rangle)P_{F}v_{i}\in\sum\limits_{i\in[m]}\upalpha_{i}\cdot P_{F}K_{\text{in}}\cdot(1-\left\langle p_{i},x\right\rangle)\stackrel{{\scriptstyle\left\langle p_{i},x\right\rangle\leq 1}}{{=}}$
[…]
4.2. Search for a good center
Now, we show that the containment $P_{F}K\subset L\cap F$ guarantees the existence of a translation, after which (4.1) holds.
Lemma 4.2.
Let $K\subseteq L$ be convex bodies in $\mathbb{R}^{d}$ such that the origin is in the interior of $K$, and there are
  1. (1)
    an $s$-dimensional subspace $F;$
  2. (2)
    contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $L$ and $K;$
  3. (3)
    positive weights $\upalpha_{1},\dots,\upalpha_{m}$
satisfying identities (1.1), (1.2), and (1.3). Additionally, assume $P_{F}K\subset L\cap F.$
Then there exists a point $z$ in the relative interior of $L\cap F$, contact pairs $\{(v_{1}^{\prime},p_{1}^{\prime}),\dots,(v_{m}^{\prime},p_{m}^{\prime})\}$ of $K-z$ and $L-z$, and weights $c_{1},\dots,c_{m}>0$ such that
${\sum_{i\in[m]}c_{i}p_{i}^{\prime}\otimes v_{i}^{\prime}}=\sum_{i\in[m]}\upalpha_{i}p_{i}\otimes v_{i},\quad\text{and}\quad\sum_{i\in[m]}c_{i}P_{F}v_{i}^{\prime}=\sum_{i\in[m]}c_{i}p_{i}^{\prime}=\mathbf{0}.$
Proof.
This is a purely linear algebraic argument.
We aim at finding a point $z$ in the relative interior of $L\cap F$ such that when we shift the origin to $z$, the shifted contact pairs $(v_{i}-z,p_{i}^{\prime})$ satisfy the required identities for $K-z$ and $L-z$ with appropriate normal vectors $p_{i}^{\prime}$.
First, observe how a contact pair transforms under translation. A pair $(v,p)$ is a contact pair of $K$ and $L$ if and only if, $v$ belongs to the intersection of the boundaries of $K$ and $L$, and the half-space
$H_{p}^{-}=\left\{x\in\mathbb{R}^{d}:\;\left\langle x,p\right\rangle\leq\left\langle v,p\right\rangle=1\right\}$
contains $L$.
Under translation by $z$, the half-space becomes
$H_{p}^{-}-z=\left\{x-z\in\mathbb{R}^{d}:\;\left\langle x,p\right\rangle-\left\langle z,p\right\rangle\leq 1-\left\langle z,p\right\rangle\right\}=\left\{y\in\mathbb{R}^{d}:\;\left\langle y,\frac{p}{1-\left\langle z,p\right\rangle}\right\rangle\leq 1\right\}.$
Thus, provided $\left\langle z,p\right\rangle<1$, the pair $(v-z,p^{\prime})$, where $p^{\prime}=\dfrac{p}{1-\left\langle z,p\right\rangle}$, is a contact pair of $K-z$ and $L-z$.
For every $i\in[m]$, define
$v_{i}(z)=v_{i}-z,\quad\text{and}\quad p_{i}(z)=\frac{p_{i}}{1-\left\langle z,p_{i}\right\rangle}.$
To satisfy the identity $\sum\limits_{i\in[m]}c_{i}(z)p_{i}(z)=\mathbf{0}$, we set
$c_{i}(z)=\upalpha_{i}\left(1-\left\langle z,p_{i}\right\rangle\right).$
Then,
$c_{i}(z)p_{i}(z)=\upalpha_{i}\left(1-\left\langle z,p_{i}\right\rangle\right)\frac{p_{i}}{1-\left\langle z,p_{i}\right\rangle}=\upalpha_{i}p_{i},$
which implies
$\sum_{i\in[m]}c_{i}(z)p_{i}(z)=\sum_{i\in[m]}\upalpha_{i}p_{i}=\mathbf{0}.$
Set $M=\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\otimes v_{i}.$ Next, we verify the identity $\sum\limits_{i\in[m]}c_{i}(z)p_{i}(z)\otimes v_{i}(z)=M$. Indeed,
$\sum_{i\in[m]}c_{i}(z)p_{i}(z)\otimes v_{i}(z)=\sum_{i\in[m]}\upalpha_{i}\left(1-\left\langle z,p_{i}\right\rangle\right)\frac{p_{i}}{1-\left\langle z,p_{i}\right\rangle}\otimes(v_{i}-z)=\sum_{i\in[m]}\upalpha_{i}p_{i}\otimes(v_{i}-z)=$
[…]
Finally, we need to satisfy $P_{F}\left(\sum\limits_{i\in[m]}c_{i}(z)v_{i}(z)\right)=\mathbf{0}$. Computing, we have
$P_{F}\left(\sum_{i\in[m]}c_{i}(z)v_{i}(z)\right)$ $=\sum_{i\in[m]}\upalpha_{i}\left(1-\left\langle z,p_{i}\right\rangle\right)P_{F}(v_{i}-z)$
$=\sum_{i\in[m]}\upalpha_{i}P_{F}v_{i}-\sum_{i\in[m]}\upalpha_{i}\left\langle z,p_{i}\right\rangle P_{F}v_{i}-P_{F}z\sum_{i\in[m]}\upalpha_{i}+P_{F}z\sum_{i\in[m]}\upalpha_{i}\left\langle z,p_{i}\right\rangle.$
Using the identities $\sum\limits_{i\in[m]}\upalpha_{i}p_{i}=\mathbf{0}$ and $M=\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\otimes v_{i}$, we can simplify:
$\sum\limits_{i\in[m]}c_{i}(z)P_{F}v_{i}(z)=\sum\limits_{i\in[m]}\upalpha_{i}P_{F}v_{i}-P_{F}\left(\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}\right)z-P_{F}z\sum\limits_{i\in[m]}\upalpha_{i}+P_{F}z\left\langle z,\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\right\rangle$
$=\sum\limits_{i\in[m]}\upalpha_{i}P_{F}v_{i}-P_{F}{M}^{\top}z-P_{F}z\left(\sum\limits_{i\in[m]}\upalpha_{i}\right)+P_{F}z\cdot 0=\left(\sum\limits_{i\in[m]}\upalpha_{i}P_{F}v_{i}\right)-P_{F}{M}^{\top}z-dP_{F}z.$
Set
$z=P_{F}\left(\frac{\sum\limits_{i\in[m]}\upalpha_{i}v_{i}}{d+1}\right).$
Since by our assumption $P_{F}K\subset L\cap F,$ and since $\sum\limits_{i\in[m]}\upalpha_{i}=d$, and $\mathbf{0}\in\mathrm{int}K$, we conclude that $z$ is in the relative interior of ${L\cap F}$. Thus, $\sum\limits_{i\in[m]}c_{i}(z)P_{F}v_{i}(z)=\mathbf{0}$, as desired. ∎
4.3. Proof of Theorem 1.5
We use Lemma 4.2 to find a “good center” $z.$ The inclusion follows from Lemma 4.1 and the inclusion $K_{\text{in}}\subset K\subset L\subset L_{\text{out}},$ where $K_{\text{in}}$ and $L_{\text{out}}$ are as defined in Subsection 4.1. In the case when $K$ is symmetric with respect to $F,$ the inclusion $P_{F}K\subset L\cap F$ holds.
4.4. Application: right cones
[…]
As an application of Lemma 4.1, we obtain the following result, whose proof is essentially the same as that of Theorem 1.5:
Proposition 4.3.
[…]
$(K-z)\cap\ell\subseteq d(C_{0}-z)\cap\ell.$
In particular, if $K$ is a convex body on the plane, and $\Delta$ is the largest area isosceles triangle in $K$ with a given direction of its axis of symmetry $\ell$, then with the right choice of $z\in\Delta\cap\ell$, we have $(K-z)\cap\ell\subseteq 2(\Delta-z)\cap\ell$. Note that the factor 2 in this formula is optimal, as shown by the example when $K$ is a square and $\ell$ is parallel to one of its diagonals.
4.5. No containment in the whole space
Contrasting Theorem 1.5, we show that there is no meaningful inclusion in ${F}^{\bot}$. Surprisingly, the example is provided by ellipsoids.
Proposition 4.4.
Fix an integer $0\leq s\leq d-2$ and a positive real $\lambda$. Let $E$ be an origin-centered ellipsoid in $\mathbb{R}^{d}$ whose principal semi-axes have lengths $\lambda,\lambda^{2},\ldots,\lambda^{d}$. Let $E^{\prime}\in\mathcal{A}_{F}\!\left(\mathbf{B}^{d}\right)$ be the solution to Problem 2, and let the $s$-dimensional subspace $F$ be the axis of $E^{\prime}$. Then
$\frac{\operatorname{vol}\nolimits_{d}E^{\prime}}{\operatorname{vol}\nolimits_{d}E}\leq\frac{1}{\lambda}\quad\text{and}\quad\forall x\in\mathbb{R}^{d},\quad E\cap{F}^{\bot}\not\subset x+\lambda\cdot\mathrm{int}P_{{F}^{\bot}}E^{\prime}.$
We need the following observation:
Lemma 4.5.
Let $E_{1}\subseteq E_{2}\subset\mathbb{R}^{d}$ be two ellipsoids. Let the principal semi-axes of $E_{1}$ be $\lambda_{1}\leq\ldots\leq\lambda_{d}$, and those of $E_{2}$ be $\mu_{1}\leq\ldots\leq\mu_{d}$. Then $\lambda_{i}\leq\mu_{i}$ for all $i\in[d]$.
Proof.
By symmetry, we may assume that both ellipsoids $E_{1}$ and $E_{2}$ are centered at the origin. Let $L_{i}$ be the $i$-dimensional subspace spanned by the principal semi-axes of $E_{1}$ with lengths $\lambda_{d-i+1},\dots,\lambda_{d}$. The Courant–Fischer min-max theorem [HJ12, Theorem 4.2.6] yields $\lambda_{d-i+1}\leq\mu_{d-i+1}$, concluding the proof. ∎
Proof of Proposition 4.4.
It follows that the $i$-th (by length) principal semi-axis of $E^{\prime}$ is at most $\lambda^{i}$ for each $i\in[d]$. Moreover, at least $d-s$ of them must be equal. By direct computation, the maximal possible volume of $E^{\prime}$ is
${\operatorname{vol}\nolimits_{d}E}{\prod\limits_{j=0}^{d-s-1}\lambda^{-j}}.$
This bound is attained whenever $E^{\prime}$ is coaxial with $E$, and the lengths of some consecutive semi-axes of $E^{\prime}$ are equal to the corresponding $\lambda^{i}$ from $E$. However, this fixes the construction: ${F}^{\bot}$ must be the linear span of $d-s$ consecutive principal semi-axes of $E$. Thus, $E^{\prime}$ is centered at the origin. The desired statements follow since the ratio of the lengths of one collinear semi-axis of $E$ and $E^{\prime}$ is $\lambda$. ∎
We note that the above argument can be extended to any convex body $K$ in the following way.
Proposition 4.6.
Let $K$ be a convex body in $\mathbb{R}^{d}$, and let $A:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be the diagonal linear operator with diagonal entries $\lambda,\lambda^{2},\ldots,\lambda^{d}$. Let $K^{\prime}=A^{\prime}K$ be a solution to Problem 4 with some linear operator $A^{\prime}$, for $K$ and $L=AK$, with any $d-s\geq 2$. Then, for any $\mu>0$ there is some value of $\lambda>0$ such that
  1. (a)
    for any $\mu^{\prime}<\mu$, no homothetic copy $x+\mu^{\prime}K^{\prime}$ of $K^{\prime}$ contains $L$,
  2. (b)
    $\frac{\operatorname{vol}\nolimits_{d}K^{\prime}}{\operatorname{vol}\nolimits_{d}L}\leq\frac{1}{\mu}$.
Proof.
Let $B_{\mathrm{is}}$ be a largest ball contained in $K$, and let $B_{\mathrm{cs}}$ be the smallest ball containing $K$. Assume that $B_{\mathrm{cs}}=y+\nu B_{\mathrm{is}}$ for $\nu\geq 1$, $y\in\mathbb{R}^{d}$.
Let $\bar{\mu}>0$ be the smallest number such that a homothetic copy of $K^{\prime}$ of ratio $\bar{\mu}$ contains $L$. Then there is a homothetic copy of $A^{\prime}B_{\mathrm{cs}}$ of ratio $\bar{\mu}$ that contains $AB_{\mathrm{is}}$. Consequently, there is a homothetic copy of $A^{\prime}B_{\mathrm{is}}$ of ratio $\bar{\mu}\nu^{2}$ that contains $AB_{\mathrm{cs}}$. Clearly, $A^{\prime}B_{\mathrm{is}}\subseteq L\subseteq AB_{\mathrm{cs}}$. Thus, by the previous example, $\bar{\mu}\nu^{2}\geq\lambda$, implying that $\bar{\mu}\geq\frac{\lambda}{\nu^{2}}$, proving (a).
To prove (b), we may apply a similar consideration. ∎
5. Maximal volume ellipsoids of revolution
First, as immediate corollaries to Theorems 1.4 and 1.6, we obtain John type necessary conditions for an ellipsoid to be the largest volume ellipsoid of revolution in a given convex body. Next, as the main goal of the section we prove Theorem 1.7, and discuss the tightness of the results. Our proof strategy is to substitute $K$ with a suitable contact polytope; see Section 4.
5.1. Contact points on the sphere
In Theorems 1.4 and 1.6, we phrased our John type conditions in terms of the contact pairs of the two bodies. Since ellipsoids are affine images of the ball, we will, in this section, phrase our John type conditions for ellipsoids of revolution in terms of unit vectors that are pre-images of contact points under this affine transformation.
As an immediate corollary of Theorem 1.4, we obtain the following answer to Problem 1.
Corollary 5.1 (Necessary condition for John’s problem for ellipsoids with a fixed axis).
Let the origin-centered ellipsoid $E=A\mathbf{B}^{d}$ be a solution to Problem 1, that is John’s problem for ellipsoids with a fixed axis of revolution for a convex body $K$ in $\mathbb{R}^{d}$. Then there are unit vectors $u_{1},\dots,u_{m}$, and positive weights $\upalpha_{1},\dots,\upalpha_{m}$ such that $(Au_{1},{A}^{-1}u_{1}),\dots,(Au_{m},{A}^{-1}u_{m})$ are contact pairs of $E$ and $K$, and
$P_{F}\left(\sum\limits_{i\in[m]}\upalpha_{i}u_{i}\otimes u_{i}\right)P_{F}=P_{F};\quad\sum\limits_{i\in[m]}\upalpha_{i}u_{i}=\mathbf{0};\quad\sum\limits_{i\in[m]}\upalpha_{i}=d.$
Proof.
Observe that at a boundary point $Au\in\operatorname{bd}A\mathbf{B}^{d}$, the corresponding normal vector is $A^{-1}u$. The statement now follows from Theorem 1.4. ∎
Note that a standard argument using Carathéodory’s theorem for the space $\mathcal{R}$ shows that the number $m$ of points needed in Corollary 5.1 is at most $2d^{2}+d+1$.
As an immediate corollary of Theorem 1.6, we obtain the following answer to Problem 2.
Corollary 5.2 (Necessary condition for John’s problem for ellipsoids with any axis).
Assume that $E=A\mathbf{B}^{d}$ is a solution to Problem 2, that is John’s problem for ellipsoids with any axis of revolution for a convex body $K$ in $\mathbb{R}^{d}$. Then there exist $m\leq 2d^{2}+d+1$ unit vectors $u_{1},\dots,u_{m}$ and positive weights $\upalpha_{1},\dots,\upalpha_{m}$ such that the conclusions of Corollary 5.1 hold as well as
$A\left(\sum\limits_{i\in[m]}\upalpha_{i}u_{i}\otimes u_{i}\right)A^{-1}={\left(A^{-1}\right)}^{\top}\left(\sum\limits_{i\in[m]}\upalpha_{i}u_{i}\otimes u_{i}\right){A}^{\top}.$
Definition 5.3.
We will call a position $K^{\prime}\in\mathcal{A}_{F}\!\left(K\right)$ with fixed axis $F$ of a convex body $K$ the John position with fixed axis $F$, if the standard unit ball is a solution to John’s general problem with fixed axis $F$ for $K^{\prime}$ and $\mathbf{B}^{d}.$
In the case $F$ is the whole space, the John position with fixed $F$ corresponds to the standard John position of a convex body.
Using the matrix $A$ from Corollary 5.1, we see that the position ${A}^{-1}K$ of $K$ is its John position with fixed axis $F$. Throughout this section, we fix the corresponding unit vectors $u_{1},\dots,u_{m}$ and positive weights $\upalpha_{1},\dots,\upalpha_{m}$. Considering the vectors $P_{F}u_{i}$ as lying in $F$, we obtain:
(5.1) $\sum\limits_{i\in[m]}\upalpha_{i}P_{F}u_{i}\otimes P_{F}u_{i}=\mathrm{Id}_{F};\quad\sum\limits_{i\in[m]}\upalpha_{i}P_{F}u_{i}=\mathbf{0};\quad\sum\limits_{i\in[m]}\upalpha_{i}=d.$
As in problems concerning convex bodies in John position, we will study the corresponding “outer contact polytopes”:
Denote
$K_{\text{J}}=\bigcap\limits_{i\in[m]}\left\{x\in\mathbb{R}^{d}:\;\left\langle x,u_{i}\right\rangle\leq 1\right\}.$
In the centrally symmetric case $K=-K$, define
$K_{\mathrm{J_{sym}}}=\bigcap\limits_{i\in[m]}\left\{x\in\mathbb{R}^{d}:\;|{\left\langle x,u_{i}\right\rangle}|\leq 1\right\}.$
5.2. Inradius in ${F}^{\bot}$
Lemma 5.4.
Let $v_{1},\dots,v_{m}$ be vectors in $\mathbf{B}^{n}.$ Assume there are positive weights $\beta_{1},\dots,\beta_{m}$ such that
$\sum\limits_{i\in[m]}\beta_{i}=1,\quad\sum\limits_{i\in[m]}\beta_{i}v_{i}=\mathbf{0},\quad\sum\limits_{i\in[m]}\beta_{i}\left|v_{i}\right|^{2}=\theta\text{ for some }\theta\in(0,1].$
Then the inradius of ${\left(\operatorname{conv}\{v_{1},\dots,v_{m}\}\right)}^{\circ}$ is at most $\frac{1}{\theta}.$
Proof.
Assume $r\mathbf{B}^{n}+c\subset{\left(\operatorname{conv}\{v_{1},\dots,v_{m}\}\right)}^{\circ}.$ This means that for every $i\in[m]$ with $v_{i}\neq 0,$ the point $c+r\frac{v_{i}}{\left|v_{i}\right|}$ is in ${\left(\operatorname{conv}\{v_{1},\dots,v_{m}\}\right)}^{\circ}.$ Hence,
$\left\langle v_{i},c+r\frac{v_{i}}{\left|v_{i}\right|}\right\rangle\leq 1\quad\Longleftrightarrow\quad\left\langle v_{i},c\right\rangle+r\left|v_{i}\right|\leq 1\quad\Longleftrightarrow\quad r\left|v_{i}\right|\leq 1-\left\langle v_{i},c\right\rangle.$
The latter inequality also holds if $v_{i}=\mathbf{0}.$ Multiplying it by $\beta_{i}$ and summing, one gets
$r\left(\sum\limits_{i\in[m]}\beta_{i}\left|v_{i}\right|\right)\leq\sum\limits_{i\in[m]}\beta_{i}-\left\langle\sum\limits_{i\in[m]}\beta_{i}v_{i},c\right\rangle.$
Since $\sum\limits_{i\in[m]}\beta_{i}v_{i}=\mathbf{0}$ and $\sum\limits_{i\in[m]}\beta_{i}=1,$ we get
$r\left(\sum\limits_{i\in[m]}\beta_{i}\left|v_{i}\right|\right)\leq 1\quad\Longleftrightarrow\quad r\leq\frac{1}{\sum\limits_{i\in[m]}\beta_{i}\left|v_{i}\right|}.$
Since $v_{i}\in\mathbf{B}^{n},$ we conclude that $\sum\limits_{i\in[m]}\beta_{i}\left|v_{i}\right|\geq\sum\limits_{i\in[m]}\beta_{i}{\left|v_{i}\right|^{2}}=\theta.$ The desired inequality follows. ∎
Lemma 5.5.
Under the notation of Subsection 5.1, with $s=\dim F$, the inradius of $K_{\text{J}}\cap{F}^{\bot}$ satisfies
$\operatorname{inrad}\!\left(K\cap{F}^{\bot}\right)\leq\frac{d}{d-s}\cdot\mathrm{radius}(E\cap{F}^{\bot}).$
Additionally, in the case when $K=-K$, the inradius of $K_{\mathrm{J_{sym}}}\cap{F}^{\bot}$ satisfies
$\operatorname{inrad}\!\left(K\cap{F}^{\bot}\right)\leq\sqrt{\frac{d}{d-s}}\cdot\mathrm{radius}(E\cap{F}^{\bot}).$
Proof.
The first identity in (5.1) yields $\sum\limits_{i\in[m]}\upalpha_{i}\left|P_{F}u_{i}\right|^{2}=s.$ Using here that $\sum\limits_{i\in[m]}\upalpha_{i}=d$, we conclude $\sum\limits_{i\in[m]}\upalpha_{i}\left|P_{{F}^{\bot}}u_{i}\right|^{2}=d-s.$
Since $\left|P_{{F}^{\bot}}u_{i}\right|\leq\left|u_{i}\right|=1$ for every $i\in[m]$, the vectors $P_{{F}^{\bot}}u_{1},\dots,P_{{F}^{\bot}}u_{m}$ satisfy the assumptions of Lemma 5.4 with $\beta_{1}=\frac{\upalpha_{1}}{d},\dots,\beta_{m}=\frac{\upalpha_{m}}{d},$ and $\theta=\frac{d-s}{d}.$ The general bound follows.
In the centrally symmetric case, we know that the origin is the center of an inball. Hence, the inradius cannot exceed the distance from the origin to a hyperplane $\left\{x\in{F}^{\bot}:\;\left\langle P_{{F}^{\bot}}u_{i},x\right\rangle=1\right\}$ in ${F}^{\bot}$ for some non-zero $P_{{F}^{\bot}}u_{i}.$ This distance is exactly $\frac{1}{\left|P_{{F}^{\bot}}u_{i}\right|}.$ Identities (1.1) and (1.3) imply that $\left|P_{{F}^{\bot}}u_{j}\right|^{2}\geq\frac{d-s}{d}$ for some $j\in[m].$ The bound follows. ∎
5.3. Volume ratio in $F$
The proofs of our volume bounds follow very closely the proof of Theorem 1.1 of [AGB21], where the authors study upper bounds on the volume of a section of a convex body in John position. The authors of [AGB21] describe the corresponding section using the projections of the unit vectors that form the John decomposition in $\mathbb{R}^{d}$. In the rest of the proof in [AGB21], only algebraic identities (5.1) and Ball’s normalized version of the Brascamp–Lieb inequality [Bal89] are used.
Lemma 5.6 (Brascamp–Lieb inequality).
Let unit vectors $v_{1},\ldots,v_{m}\in\mathbb{R}^{n}$ and positive numbers $\beta_{1},\ldots,\beta_{m}$ satisfy
$\sum\limits_{i\in[m]}\beta_{i}v_{i}\otimes v_{i}=\mathrm{Id}_{\mathbb{R}^{n}}.$
Then
(5.2) $\int_{\mathbb{R}^{n}}\prod_{i\in[m]}f_{i}(\left\langle x,v_{i}\right\rangle)^{\beta_{i}}\,\mathrm{d}x\leq\prod_{i\in[m]}\left(\int_{\mathbb{R}}f_{i}\right)^{\beta_{i}},$
holds for all integrable functions $f_{1},\ldots,f_{m}\colon\mathbb{R}\to[0,\infty).$
Lemma 5.7.
Under the notation of Subsection 5.1, with $s=\dim F$, the inequality
$\left(\frac{\operatorname{vol}\nolimits_{s}\left(K_{\mathrm{J_{sym}}}\cap F\right)}{\operatorname{vol}\nolimits_{s}\Box^{s}}\right)^{\frac{1}{s}}\leq\sqrt{\frac{d}{s}}$
holds in the centrally symmetric case, $K=-K$.
Proof.
Let $I\subset[m]$ be the set of those indices $i\in[m]$ for which $P_{F}u_{i}\neq 0$. Set $v_{i}=\frac{P_{F}u_{i}}{\left|P_{F}u_{i}\right|}$ and $\beta_{i}={\upalpha_{i}}{\left|P_{F}u_{i}\right|^{2}}$ for all $i\in I.$ Then
$\sum_{i\in I}\beta_{i}v_{i}\otimes v_{i}=\sum_{i\in[m]}\upalpha_{i}P_{F}u_{i}\otimes P_{F}u_{i}=\mathrm{Id}_{F}.$
Let $f_{i}$ denote the indicator function of the interval $\left[-\frac{1}{\left|P_{F}u_{i}\right|},\frac{1}{\left|P_{F}u_{i}\right|}\right]$ for each $i\in I$.
We leave it to the reader to check that the integral on the left-hand side of (5.2) is equal to the volume of $K_{\mathrm{J_{sym}}}\cap F$, as we identify $F$ with $\mathbb{R}^{n}$ (where $n=s$), and that the product of the integrals on the right-hand side of (5.2) is equal to $2^{s}\prod\limits_{i\in I}\left|P_{F}u_{i}\right|^{-\upalpha_{i}\left|P_{F}u_{i}\right|^{2}}.$ The latter quantity is at most $2^{s}\left(\frac{d}{s}\right)^{\frac{s}{2}}$ by Jensen’s inequality, completing the proof of Lemma 5.7. ∎
Lemma 5.8.
Under the notation of Subsection 5.1, with $s=\dim F$, the inequality
$\left(\frac{\operatorname{vol}\nolimits_{s}\left(K_{\text{J}}\cap F\right)}{\operatorname{vol}\nolimits_{s}\Delta_{J}^{s}}\right)^{\frac{1}{s}}\leq\frac{1}{(s+1)^{\frac{d-s}{2s(d+1)}}}\sqrt{\frac{d(d+1)}{s(s+1)}}$
holds.
Proof.
[…]
Let $I\subset[m]$ denote the subset of indices $i\in[m]$ for which $P_{F}u_{i}\neq 0$. Denote $d^{\prime}=\sum\limits_{i\in I}\upalpha_{i};$ and set
$\widetilde{q}_{i}=\left(-\sqrt{\frac{d^{\prime}}{d^{\prime}+1}}P_{F}u_{i},\frac{1}{\sqrt{d^{\prime}+1}}\right),\quad\text{and }\quad\widetilde{\upalpha}_{i}=\frac{d^{\prime}+1}{d^{\prime}}\upalpha_{i}.$
Set $v_{i}=\frac{\widetilde{q}_{i}}{\left|\widetilde{q}_{i}\right|}$ and $\beta_{i}={\widetilde{\upalpha}_{i}}{\left|\widetilde{q}_{i}\right|^{2}}$ for all $i\in I.$ By straightforward computation,
$\sum\limits_{i\in I}\beta_{i}v_{i}\otimes v_{i}=\sum\limits_{i\in I}\widetilde{\upalpha}_{i}\widetilde{q}_{i}\otimes\widetilde{q}_{i}=\mathrm{Id}_{\widetilde{F}}.$
We will denote the function $t\mapsto e^{-\frac{t}{\left|\widetilde{q}_{i}\right|}}\mathbf{1}_{[0,+\infty)}$ by $f_{i}$ for each $i\in I,$ where $\mathbf{1}_{S}$ stands for the indicator function of the set $S.$ Then, the product of the integrals on the right-hand side of (5.2) is equal to $\prod\limits_{i\in I}\left|P_{F}u_{i}\right|^{\upalpha_{i}\left|P_{F}u_{i}\right|^{2}}.$ The bound
(5.3) $\prod\limits_{i\in I}\left|P_{F}u_{i}\right|^{\upalpha_{i}\left|P_{F}u_{i}\right|^{2}}\leq\frac{(s+1)^{\frac{s+1}{2(d^{\prime}+1)}}}{(d^{\prime}+1)^{\frac{d^{\prime}+1}{2(d^{\prime}+1)}}}$
follows from the standard majorization argument for the convex function $t\mapsto t\ln t$ on (0,1)(0,1)( 0 , 1 ). This bound was obtained in the proof of Theorem 1.1 in [AGB21]. For the sake of completeness, we provide the proof as a separate lemma in Appendix B.
We claim that the integral on the left-hand side of (5.2) is
$\int_{\tilde{F}}\prod_{i\in I}f_{i}(\left\langle x,v_{i}\right\rangle)^{\beta_{i}}\,\mathrm{d}x=\frac{s!}{(d^{\prime}+1)^{\frac{s+1}{2}}{d^{\prime}}^{\frac{s}{2}}}\operatorname{vol}\nolimits_{s}\left(K_{\text{J}}\cap F\right).$
Indeed, set $f(x)=\prod\limits_{i\in I}f_{i}(\left\langle x,v_{i}\right\rangle)^{\beta_{i}}$ for $x\in\tilde{F}$. The support of $f$ is the cone $\mathrm{Pos}\ \!{\left\{\left(\frac{K_{\text{J}}\cap F}{\sqrt{d^{\prime}}},1\right)\right\}}$. Now, $f$ is constant $e^{-\tau{\sqrt{d^{\prime}+1}}}$ at the intersection of the hyperplane $H_{\tau}=\left\{(y,t)\in\widetilde{F}:\;t=\tau\right\}$ and this cone for any positive $\tau$. By monotonicity in $d^{\prime}$, the proof of Lemma 5.8 is complete. ∎
Remark 5.9.
The authors of [AGB21] used a different approach in the general case: they first lift the “contact vectors” to $\mathbb{R}^{d}\times\mathbb{R}$ and then project the configuration onto $F\times\mathbb{R}.$ This minor variation in the argument does not substantially affect the subsequent computations.
5.4. Proof of Theorem 1.7
A transformation of the space by an $F$-matrix changes both the volume form on $F$ and the distance in ${F}^{\bot}$ proportionally. Hence, under the notation of Subsection 5.1, it suffices to bound the volume of ${A}^{-1}K\cap F$ and the inradius of ${A}^{-1}K\cap{F}^{\bot}$ to obtain the corresponding bound.
By construction, ${A}^{-1}K\subset K_{\text{J}}$ and ${A}^{-1}K\subset K_{\mathrm{J_{sym}}}$ in the centrally symmetric case $K=-K.$ The corresponding volume bounds follow from Lemma 5.7 and Lemma 5.8; the inradius bound follows from Lemma 5.5.
The inclusion $E\cap F\subset K\cap F\subset d\left(E\cap F\right)$ follows from Theorem 1.5 with $z=\mathbf{0}$ and $P_{F}E=E\cap F\subset K\cap F.$
Let us now establish the inclusion
$E\cap F\subset K\cap F\subset\sqrt{d}\left(E\cap F\right)$
in the centrally symmetric case $K=-K$. Fix $x\in K_{\mathrm{J_{sym}}}\cap F.$ One has
[…]
$\sum\limits_{i\in[m]}\upalpha_{i}\left\langle x,u_{i}\right\rangle\left\langle u_{i},x\right\rangle=\sum\limits_{i\in[m]}\upalpha_{i}\left\langle x,u_{i}\right\rangle^{2}\leq\sum\limits_{i\in[m]}\upalpha_{i}=d.$
The desired bound follows. The proof of Theorem 1.7 is complete.
5.5. Tightness of the results
The inclusions of Theorem 1.7 are optimal since they are already optimal for the John ellipsoid as shown by the simplex and the cube. The volume bounds of Lemma 5.7 and Lemma 5.8 coincide with the corresponding bounds for an $s$-dimensional section of a $d$-dimensional convex body in John position (see Theorem 1.1 of [AGB21]). In the centrally symmetric case $K=-K,$ we conjecture the following:
Conjecture 5.10.
Let an origin-centered ellipsoid $E$ be a solution to Problem 1 for a centrally symmetric convex body $K$ and a linear subspace $F$ in $\mathbb{R}^{d}$ of dimension $s\in[d]$. Then
$\frac{\operatorname{vol}\nolimits_{s}\left(K\cap F\right)}{\operatorname{vol}\nolimits_{s}\left(E\cap F\right)}\leq C_{\Box}(d,s)\frac{\operatorname{vol}\nolimits_{s}\Box^{s}}{\operatorname{vol}\nolimits_{s}\mathbf{B}^{s}},$
where the constant $C_{\Box}(d,s)$ is given by
$C^{2}_{\Box}(d,s)=\left\lceil\frac{d}{s}\right\rceil^{d-s\lfloor d/s\rfloor}\left\lfloor\frac{d}{s}\right\rfloor^{s-(d-s\lfloor d/s\rfloor)}.$
The constant $C_{\Box}(d,s)$ in Conjecture 5.10 is attained on certain sections of the cube $\Box^{d}=[-1,1]^{d}$ by $s$-dimensional subspaces (see Lemma 2.2 of [IT21]). Peculiarly, in the case when $K$ is centrally symmetric and $s$ divides $d,$ both our volume ratio bound in $F$ and the inradius bound in ${F}^{\bot}$ are simultaneously tight: $K$ is the cube $[-1,1]^{d};$ $E$ is the John ellipsoid $\mathbf{B}^{d}$ of $K;$ $F$ is the solution set to the system of linear equations:
$\left\langle x,e_{\frac{d}{s}(i-1)+1}\right\rangle=\ldots=\left\langle x,e_{\frac{d}{s}i}\right\rangle,\quad\text{for all}\quad i\in[s].$
Moreover, the bound $\sqrt{\frac{d}{d-s}}$ on the inradius in ${F}^{\bot}$ is attained on certain sections of the cube $\Box^{d}.$ Indeed, denoting the standard basis of $\mathbb{R}^{d}$ by $u_{1},\dots,u_{d},$ the estimate $\sqrt{\frac{d}{d-s}}$ is attained whenever $\left|P_{{F}^{\bot}}u_{1}\right|^{2}=\dots=\left|P_{{F}^{\bot}}u_{d}\right|^{2}=\frac{d-s}{d}.$ The existence of subspaces with this property was proven in Lemma 2.4 of [Iva20], in which the set of all possible lengths of the projections was described.
The general case of a not necessarily symmetric body is more intricate. As shown in [Dir17], the bound in Lemma 5.8 is asymptotically tight for a regular simplex $\Delta_{J}^{d}$ in the John position. Consequently, it is asymptotically tight in our case. However, we believe that there is an intrinsic problem in obtaining the best possible bound using the Brascamp–Lieb inequality directly: if one chooses the parameters to maximize the functional in Lemma B.1, then the Brascamp–Lieb inequality is not optimal for the corresponding functions. That is, there are geometric obstructions to achieving the equality case in the Brascamp–Lieb inequality.
We also have an example showing that the bound Theorem 1.7 (1) on the inradius is the best possible. It is technical, and we state and prove it in Theorem A.4. Nevertheless, the corresponding vectors in $\mathbb{R}^{d-s}$ can be lifted to unit vectors satisfying the algebraic equations of Corollary 5.1. A more detailed analysis of Lemma 5.4 shows that the bound can be achieved, if certain vectors are close to zero and the others lie near the unit sphere, so it cannot be achieved via the projections of the unit vectors from the standard John condition for $F=\mathbb{R}^{d}$ (see [BP09]).
Finally, [IN22] by M. Naszódi and G. Ivanov may be interpreted as studying John’s general problem with fixed axis $F$ for $\mathbf{B}^{n+1}$ and $K$ being a certain not necessarily convex set in $\mathbb{R}^{n+1}$ associated with a log-concave function on $\mathbb{R}^{n}$, and $F$ being $\mathbb{R}^{n}.$ Moreover, the algebraic conditions for the “contact points” in that case coincide with those of Corollary 5.1; and the more general case considered in [IN23] fits within the framework of Theorem 1.4 for appropriately chosen parameter sets. The inclusions for ellipsoids and “contact-type” polytopes have been studied previously in the functional setting as well; see, for example, [Iva24, Lemma 4.1].
6. Uniqueness of the solution
We note that in general, the solution of John’s problem for ellipsoids with any axis of revolution is not unique: consider the convex body $K=E+\left[-v,v\right]$, where $E$ is an $F$-ellipsoid for some $s$-dimensional subspace $F$ and $v\in{F}^{\bot}$. On the other hand, we have the following theorem.
For a linear subspace $F$ and a subset $K$ of $\mathbb{R}^{d}$, we denote by $\mathcal{S}_{F}(K)$ the set of those elements of $\mathcal{A}_{F}\!\left(K\right)$, for which the corresponding linear transformation is positive definite.
Theorem 6.1.
For any convex bodies $K,L$ in $\mathbb{R}^{d}$, and for any elements $K_{1},K_{2}\in\mathcal{S}_{F}(K)$ of maximal volume satisfying $K_{1},K_{2}\subseteq L$, $K_{1}$ and $K_{2}$ are translates of each other.
Proof.
[…]
Consider some positive definite $F$-operators $S_{1},S_{2}$ and vectors $x_{1},x_{2}\in\mathbb{R}^{d}$ such that for $i=1,2$, $K_{i}=x_{i}+S_{i}(K)$ maximize volume. By convexity, we have that $\bar{K}=\frac{x_{1}+x_{2}}{2}+\frac{1}{2}(S_{1}+S_{2})(K)\subseteq L$. On the other hand, it is well-known that for any two positive definite matrices $A,B$,
$\det\frac{A+B}{2}\geq\sqrt{\det A\cdot\det B}$
with equality if and only if $A=B$ (see e.g. [HJ12, Corollary 7.6.8]). Since $\frac{1}{2}(S_{1}+S_{2})$ is also a positive definite $F$-operator, it follows from the maximality of volume that $S_{1}=S_{2}$, implying that $K_{1}$ and $K_{2}$ are translates of each other. ∎
Remark 6.2.
We note that $\mathcal{S}_{F}(\mathbf{B}^{d})=\mathcal{A}_{F}\!\left(\mathbf{B}^{d}\right)$ for every linear subspace $F$, by the polar decomposition of matrices.
Remark 6.3.
It is not true that if $K_{1},K_{2}\in\mathcal{S}_{F}(K)$ are of maximal volume satisfying $K_{1},K_{2}\subseteq L$, then the translation vector moving $K_{1}$ to $K_{2}$ is necessarily parallel to ${F}^{\bot}$, as shown by the following example. Set $d=2$ and $K=\mathbf{B}^{2}$. We define $L$ as an $\mathbf{0}$-symmetric rectangle containing $\mathbf{B}^{2}$ such that the sides of $L$ are parallel to the lines $y=\pm x$, and the only points of $\operatorname{bd}L$ on $\mathbf{B}^{2}$ are the points ±$\pm\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right)$. Let $F$ be the $y$-axis. Then an elementary computation shows that the only $\mathbf{0}$-symmetric $F$-ellipse inscribed in $L$ is $\mathbf{B}^{2}$, and thus, this is of maximal area. On the other hand, moving this ellipse along the line $y=-x$ yields other maximal area $F$-ellipses in $L$.
7. $F$-revolution Löwner ellipsoids
A dual construction to the largest volume ellipsoid contained in a convex body is the smallest volume ellipsoid containing the body, usually called the Löwner ellipsoid. Notably, the necessary and sufficient conditions for the Euclidean unit ball to be the Löwner ellipsoid coincide with the conditions in John’s characterization of the largest volume ellipsoid. The extension of the notion of the Löwner ellipsoid is straightforward – swap the bodies in John’s general problem with fixed axis $F$, that is, consider John’s general problem with fixed axis $F$ for $K$ and $\mathbf{B}^{d}.$
Definition 7.1.
We will call a position $K^{\prime}\in\mathcal{A}_{F}\!\left(K\right)$ with fixed axis $F$ of a convex body $K$ the Löwner position with fixed axis $F$ if $K^{\prime}$ is a solution to John’s general problem with fixed axis $F$ for $K$ and $\mathbf{B}^{d}.$
The following result is a dual version of Theorem 1.7:
Theorem 7.2.
Let a convex body $K$ be in the Löwner position with fixed axis $F$ of dimension $s\in\{0,1,\dots,d\}$. Then
  1. (1)
    the following outer volume bound holds:
    $\left(\frac{\operatorname{vol}\nolimits_{s}P_{F}K}{\operatorname{vol}\nolimits_{s}\Delta_{L}^{s}}\right)^{\frac{1}{s}}\geq\sqrt{\frac{s}{d}},$
    where $\Delta_{L}^{s}$ stands for the regular $s$-dimensional simplex inscribed in $\mathbf{B}^{s}$;
  2. (2)
    the following inclusion holds:
    $P_{F}K\subset P_{F}\mathbf{B}^{d}\subset d\cdot P_{F}K;$
  3. (3)
    the circumradius of $P_{{F}^{\bot}}K$ is at least $\sqrt{\frac{d-s}{d}}.$
If, additionally, $K$ is centrally symmetric, then
  1. (S1)
    the following outer volume bound holds:
    $\left(\frac{\operatorname{vol}\nolimits_{s}P_{F}K}{\operatorname{vol}\nolimits_{s}\Diamond^{s}}\right)^{\frac{1}{s}}\geq\sqrt{\frac{s}{d}},$
    where $\Diamond^{s}$ stands for the standard cross-polytope;
  2. (S2)
    the following inclusion holds:
    $P_{F}K\subset P_{F}\mathbf{B}^{d}\subset\sqrt{d}\cdot P_{F}K.$
We will prove this theorem in the current section. The proof mostly consists of standard arguments dual to those used in the proof of Theorem 1.7. The only surprise is that the bound (3) on the circumradius does not require $K$ to be symmetric!
7.1. Notation
As a direct corollary of Theorem 1.4, we obtain:
Corollary 7.3.
Let a convex body $K$ be in Löwner position with fixed axis $F.$ Then there exist $m\leq 2d^{2}+d+1$, unit vectors $u_{1},\dots,u_{m}$, and positive weights $\upalpha_{1},\dots,\upalpha_{m}$ such that $(u_{1},u_{1}),\dots,(u_{m},u_{m})$ are contact pairs of $K$ and $\mathbf{B}^{d}$, and
$P_{F}\left(\sum\limits_{i\in[m]}\upalpha_{i}u_{i}\otimes u_{i}\right)P_{F}=P_{F};\quad\sum\limits_{i\in[m]}\upalpha_{i}u_{i}=\mathbf{0};\quad\sum\limits_{i\in[m]}\upalpha_{i}=d.$
In the case where $F$ is the whole space, the Löwner position with fixed axis $F$ coincides with the standard Löwner position of a convex body.
Throughout this section, we fix the corresponding unit vectors $u_{1},\dots,u_{m}$ and positive weights $\upalpha_{1},\dots,\upalpha_{m}$. Considering the vectors $P_{F}u_{i}$ as lying in $F$, we observe that they satisfy equations (5.1).
As in problems concerning convex bodies in the John position, we will study the corresponding “contact polytopes”:
Define
$K_{\text{L}}=\operatorname{conv}\left\{u_{1},\dots,u_{m}\right\}.$
In the centrally symmetric case $K=-K$, define
$K_{\mathrm{L_{sym}}}=\operatorname{conv}\left\{\pm u_{1},\dots,\pm u_{m}\right\}.$
7.2. Circumradius in ${F}^{\bot}$
Lemma 7.4.
Let $v_{1},\dots,v_{m}$ be vectors in $\mathbb{R}^{n}.$ Assume that there are positive weights $\beta_{1},\dots,\beta_{m}$ such that
$\sum\limits_{i\in[m]}\beta_{i}=1,\quad\sum\limits_{i\in[m]}\beta_{i}v_{i}=\mathbf{0},\quad\sum\limits_{i\in[m]}\beta_{i}\left|v_{i}\right|^{2}=\theta.$
Then the circumradius of $\operatorname{conv}\{v_{1},\dots,v_{m}\}$ is at least $\sqrt{\theta}.$
Proof.
Assume $r\mathbf{B}^{n}+c\supset\operatorname{conv}\{v_{1},\dots,v_{m}\}.$ This means that $\left|c-v_{i}\right|^{2}\leq r^{2}$ for every $i\in[m].$ Multiplying both sides by $\beta_{i}$ and summing over $i$, we get:
$r^{2}\left(\sum\limits_{i\in[m]}\beta_{i}\right)\geq\sum\limits_{i\in[m]}\beta_{i}\left|c-v_{i}\right|^{2}=\left|c\right|^{2}\left(\sum\limits_{i\in[m]}\beta_{i}\right)-2\left\langle c,\sum\limits_{i\in[m]}\beta_{i}v_{i}\right\rangle+\sum\limits_{i\in[m]}\beta_{i}\left|v_{i}\right|^{2}.$
Since $\sum\limits_{i\in[m]}\beta_{i}v_{i}=\mathbf{0}$ and $\sum\limits_{i\in[m]}\beta_{i}=1$, we obtain
$r^{2}\geq\sum\limits_{i\in[m]}\beta_{i}\left|v_{i}\right|^{2}+\left|c\right|^{2}\geq\sum\limits_{i\in[m]}\beta_{i}\left|v_{i}\right|^{2}=\theta.$
The desired inequality follows. ∎
Combining (5.1) with the above lemma with $v_{i}=P_{F^{\perp}}u_{i}$ and $\beta_{i}=\upalpha_{i}/d$, and by identifying $F^{\perp}$ with $\mathbb{R}^{n}$ (where $n=d-s$), we conclude the following:
Lemma 7.5.
Under the notation of Subsection 7.1, the circumradius of $P_{{F}^{\bot}}K_{L}$ is at least $\sqrt{\frac{d-s}{d}}.$
7.3. Outer volume ratio in $F$
The proofs of our outer volume bounds follow from Theorem 1.2 of [AGB21], where the authors study lower bounds on the volume of projections of convex bodies in the Löwner position. Starting with the breakthrough work [Bar97] of Barthe, the main tool for obtaining outer volume ratio bounds is the following reverse form of the Brascamp–Lieb inequality:
Lemma 7.6 (Reverse Brascamp–Lieb inequality).
Let unit vectors $v_{1},\ldots,v_{m}\in\mathbb{R}^{n}$ and positive numbers $\beta_{1},\ldots,\beta_{m}$ satisfy $\sum\limits_{i\in[m]}\beta_{i}v_{i}\otimes v_{i}=\mathrm{Id}_{\mathbb{R}^{n}}$, and $g_{1},\ldots,g_{m}\colon\mathbb{R}\to[0,\infty)$ be integrable functions. Define
$G(x)=\sup\left\{\prod\limits_{i\in[m]}g_{i}(\theta_{i})^{\beta_{i}}:\;\theta_{i}\in\mathbb{R}\text{ and }x=\sum\limits_{i\in[m]}\beta_{i}\theta_{i}v_{i}\right\}.$
Then,
(7.1) $\int_{\mathbb{R}^{n}}G\geq\prod\limits_{i\in[m]}\left(\int g_{i}\right)^{\beta_{i}}.$
To simplify the reading, we provide a sketch of the proof of the following lemma, omitting technical details.
Lemma 7.7.
Under the notation of Subsection 7.1, the inequality
$\left(\frac{\operatorname{vol}\nolimits_{s}P_{F}K_{\mathrm{L_{sym}}}}{\operatorname{vol}\nolimits_{s}\Diamond^{s}}\right)^{\frac{1}{s}}\geq\sqrt{\frac{s}{d}}$
holds.
Proof.
Let $I\subset[m]$ be the set of those indices $i\in[m]$ for which $P_{F}u_{i}\neq 0.$ Set $v_{i}=\frac{P_{F}u_{i}}{\left|P_{F}u_{i}\right|}$ and $\beta_{i}={\upalpha_{i}}{\left|P_{F}u_{i}\right|^{2}}$ for all $i\in I.$ Then
$\sum_{i\in I}\beta_{i}v_{i}\otimes v_{i}=\sum_{i\in[m]}\upalpha_{i}P_{F}u_{i}\otimes P_{F}u_{i}=\mathrm{Id}_{F}.$
Define
$g_{i}(t)=e^{-\frac{\left|t\right|}{\left|P_{F}u_{i}\right|}}$
for each $i\in I$.
We leave it to the reader to verify that the product of the integrals on the right-hand side of (7.1) equals $2^{s}\prod\limits_{i\in I}\left|P_{F}u_{i}\right|^{\upalpha_{i}\left|P_{F}u_{i}\right|^{2}}$. This is at least $2^{s}\left(\frac{d}{s}\right)^{\frac{s}{2}}$ by Jensen’s inequality.
On the other hand, the corresponding function $G$ equals $e^{-\left\|\,\cdot\,\right\|_{K_{\mathrm{L_{sym}}}}},$ where $\left\|\,\cdot\,\right\|_{K_{\mathrm{L_{sym}}}}$ is the gauge function of $K_{\mathrm{L_{sym}}}:$
$\left\|x\right\|_{K_{\mathrm{L_{sym}}}}=\inf\left\{t>0:\;x\in tK_{\mathrm{L_{sym}}}\right\}.$
Hence, by standard computation, the integral of $G$ is
$\int_{\mathbb{R}^{s}}G=s!\operatorname{vol}\nolimits_{s}P_{F}K_{\mathrm{L_{sym}}}.$
The desired bound follows, since $\operatorname{vol}\nolimits_{s}\Diamond^{s}=\frac{2^{s}}{s!}.$
Lemma 7.8.
Under the notation of Subsection 7.1, the inequality
$\left(\frac{\operatorname{vol}\nolimits_{s}P_{F}K_{L}}{\operatorname{vol}\nolimits_{s}\Delta_{L}^{s}}\right)^{\frac{1}{s}}\geq\sqrt{\frac{s}{d}}$
holds.
Proof.
[…]
$\widetilde{q}_{i}=\left(-\sqrt{\frac{d^{\prime}}{d^{\prime}+1}}P_{F}u_{i},\frac{1}{\sqrt{d^{\prime}+1}}\right),\quad\widetilde{\upalpha}_{i}=\frac{d^{\prime}+1}{d^{\prime}}\upalpha_{i}.$
Set $v_{i}=\frac{\widetilde{q}_{i}}{\left|\widetilde{q}_{i}\right|}$ and $\beta_{i}=\widetilde{\upalpha}_{i}\left|\widetilde{q}_{i}\right|^{2}$ for all $i\in I.$ Then,
$\sum\limits_{i\in I}\beta_{i}v_{i}\otimes v_{i}=\sum\limits_{i\in I}\widetilde{\upalpha}_{i}\widetilde{q}_{i}\otimes\widetilde{q}_{i}=\mathrm{Id}_{\widetilde{F}}.$
Define a gauge-like function $C$ on $\widetilde{F}$ by
$C(x)=\inf\left\{\sum_{i\in I}\lambda_{i}\colon\lambda_{i}\geq 0\text{ and }x=\sum_{i\in I}\lambda_{i}\widetilde{q}_{i}\right\},$
where we set $C(x)=+\infty$, if such a decomposition does not exist. Let $G(x)=e^{-C(x)}$.
By convexity, the support of $G$ is the cone $\mathrm{Pos}\ \!{\left(P_{F}K_{\text{L}},\frac{1}{\sqrt{d^{\prime}}}\right)}$; the function $G$ is constant and equal to $e^{-\tau\sqrt{d^{\prime}+1}}$ on the intersection of this cone with the hyperplane $H_{\tau}=\left\{(x,t)\in\widetilde{F}:\;t=\tau\right\}$ for any $\tau>0$. Thus,
$\int_{\widetilde{F}}G=\operatorname{vol}\nolimits_{s}P_{F}K_{\text{L}}\cdot\int_{0}^{\infty}e^{-r\sqrt{d^{\prime}+1}}(r\sqrt{d^{\prime}})^{s}\,\mathrm{d}r=\frac{s!(d^{\prime})^{s/2}}{(d^{\prime}+1)^{(s+1)/2}}\operatorname{vol}\nolimits_{s}P_{F}K_{\text{L}}.$
Now define
$g_{i}(t)=e^{-t/\left|\widetilde{q}_{i}\right|}\cdot\mathbf{1}_{[0,+\infty)}(t)$
for each $i\in I$. Then, for any $x$ in the support of $G,$
$\sup\left\{\prod\limits_{i\in I}g_{i}(\theta_{i})^{\beta_{i}}:\;x=\sum\limits_{i\in I}\beta_{i}\theta_{i}v_{i}\right\}\stackrel{{\scriptstyle\lambda_{i}=\widetilde{\upalpha}_{i}\left|\widetilde{q}_{i}\right|\theta_{i}}}{{=}}e^{-\inf\left\{\sum\limits_{i\in I}\lambda_{i}:\;x=\sum\limits_{i\in I}\lambda_{i}\widetilde{q}_{i}\right\}}=G(x).$
Thus, by the reverse Brascamp–Lieb inequality (Lemma 7.6),
$\int_{\mathbb{R}^{d}}G\geq\prod\limits_{i\in I}\left|\widetilde{q}_{i}\right|^{\widetilde{\upalpha}_{i}\left|\widetilde{q}_{i}\right|^{2}}.$
Applying Jensen’s inequality with $\sum\limits_{i\in I}\widetilde{\upalpha}_{i}\left|\widetilde{q}_{i}\right|^{2}=s+1$ and $\sum\limits_{i\in I}\widetilde{\upalpha}_{i}=d^{\prime}+1$, we obtain
$\prod\limits_{i\in I}\left|\widetilde{q}_{i}\right|^{\widetilde{\upalpha}_{i}\left|\widetilde{q}_{i}\right|^{2}}\geq\left(\frac{s+1}{d^{\prime}+1}\right)^{(s+1)/2}.$
Substituting into the expression for $\int_{\widetilde{F}}G$, we find
$\operatorname{vol}\nolimits_{s}P_{F}K_{\text{L}}\geq\left(\frac{s+1}{d^{\prime}+1}\right)^{(s+1)/2}\frac{(d^{\prime}+1)^{(s+1)/2}}{s!(d^{\prime})^{s/2}}=\left(\frac{s}{d^{\prime}}\right)^{s/2}\operatorname{vol}\nolimits_{s}\Delta_{L}^{s}.$
The desired bound follows since $d^{\prime}\leq d.$
Remark 7.9.
The bound is optimal whenever $s+1$ divides $d+1$. It is attained when $K=\Delta_{L}^{d}$ and $F$ is the subspace spanned by the centroids of $s+1$ pairwise disjoint $\left(\frac{d+1}{s+1}-1\right)$-dimensional faces of $\Delta_{L}^{d}.$
7.4. Proof of Theorem 7.2
Proof of Theorem 7.2.
We use the notation introduced in Subsections 5.1 and 7.1. By duality, $K_{\text{L}}={\left(K_{\text{J}}\right)}^{\circ}$ and $K_{\mathrm{L_{sym}}}={\left(K_{\mathrm{J_{sym}}}\right)}^{\circ}.$ Hence, the inclusions follow from the corresponding inclusions of Theorem 1.7. The bound on the circumradius was obtained in Lemma 7.5. The volume bounds were proven in Lemma 7.8 in the general case and in Lemma 7.7 in the centrally symmetric case $K=-K.$
Appendix A Tightness of the inradius bound
The goal of this section is to state and prove Theorem A.4 according to which the bound given in (1) of Theorem 1.7 on the inradius is the best possible.
Set
$a=\frac{1}{\sqrt{n}}(1,\dots,1)\in\mathbb{R}^{n},\quad n\geq 2.$
We will define vectors $v_{1},\dots,v_{n+1}$ that form a directionally dilated regular simplex along the direction of $a$.
For each $i\in[n]$, define the vector $p_{i}\in\mathbb{R}^{n}$ whose $j$-th coordinate is given by
$p_{i}[j]=\begin{cases}\frac{n-1}{n},&\text{if }j=i,\\ -\frac{1}{n},&\text{if }j\neq i.\end{cases}$
Then $\left|p_{i}\right|^{2}=\frac{n-1}{n}.$ Define $q_{i}=\frac{p_{i}}{\left|p_{i}\right|}\in\mathbb{R}^{n}.$ Note that for each $i\in[n]$, the vectors $p_{i}$ and $q_{i}$ belong to ${a}^{\bot}$, and that $\left\langle q_{i},q_{j}\right\rangle=-\frac{1}{n-1}$ for any $1\leq i<j\leq n$.
For any $t\in(0,1)$, define the vectors
$v_{i}=tq_{i}+\sqrt{1-t^{2}}\cdot a,\quad\text{for }i\in[n]$
Now, fix a positive constant $\gamma$ satisfying
(A.1) $n\gamma<1.$
We observe that
$\sum_{i\in[n]}p_{i}=\mathbf{0}=\sum_{i\in[n]}q_{i},$
$\sum_{i\in[n]}v_{i}=t\sum_{i\in[n]}q_{i}+n\sqrt{1-t^{2}}\cdot a=n\sqrt{1-t^{2}}\cdot a.$
Define $v_{n+1}$ by the identity
(A.2) $\gamma\sum_{i\in[n]}v_{i}+(1-n\gamma)v_{n+1}=\mathbf{0}.$
So,
$v_{n+1}=-\frac{\gamma n}{1-\gamma n}\sqrt{1-t^{2}}\cdot a.$
Our second restriction is
(A.3) $\frac{\gamma n}{1-\gamma n}\sqrt{1-t^{2}}<1.$
Denote $\Delta=\operatorname{conv}\left\{v_{1},\dots,v_{n+1}\right\}$. Let us estimate the inradius of the polar ${\Delta}^{\circ}$ of $\Delta$. A direct computation yields the following:
Lemma A.1.
The vertices of the simplex ${\Delta}^{\circ}$ are
$x_{n+1}=\frac{1}{\sqrt{1-t^{2}}}a,$
and
$x_{i}=-\frac{1}{n\gamma}\frac{(n-1)}{t}q_{i}-\frac{1-n\gamma}{n\gamma}\cdot\frac{1}{\sqrt{1-t^{2}}}a,\quad\text{for }i\in[n].$
Lemma A.2.
The section of ${\Delta}^{\circ}$ by the plane $L$ spanned by the vectors $x_{1}$ and $x_{n+1}$ is the triangle $T$ with vertices $x_{1}$, $x_{n+1}$, and $y$, given by
$y=\frac{1}{n\gamma}\cdot\frac{q_{1}}{t}-\frac{1-n\gamma}{n\gamma}\cdot\frac{a}{\sqrt{1-t^{2}}}.$
Proof.
Clearly,
$L=\operatorname{span}\left\{v_{1},v_{n+1}\right\}=\operatorname{span}\left\{q_{1},a\right\},$
and $q_{1}$ and $a$ form an orthonormal basis of $L$.
By duality, the polar in $L$ of ${\Delta}^{\circ}\cap L$ is the orthogonal projection of $\Delta$ onto $L$. We compute this projection. First, $v_{1}$ and $v_{n+1}$ already lie in $L$. Second, for each $i\in[n]\setminus\{1\}$,
$P_{L}v_{i}=\left\langle v_{i},q_{1}\right\rangle q_{1}+\left\langle v_{i},a\right\rangle a=t\left\langle q_{1},q_{i}\right\rangle q_{1}+\sqrt{1-t^{2}}\cdot a=-\frac{t}{n-1}q_{1}+\sqrt{1-t^{2}}\cdot a,$
where $P_{L}$ denotes the orthogonal projection onto $L$. In particular, all $v_{2},\dots,v_{n}$ project onto the same point.
Thus, $P_{L}\Delta$ is the triangle with vertices $v_{1}$, $v_{n+1}$, and $P_{L}v_{2}$. Hence, ${\Delta}^{\circ}\cap L$ is a triangle, say $T$. Let us compute its vertices. Since $x_{1}$ and $x_{n+1}$ lie in $L$, they are two of the vertices of $T$. The remaining vertex $y$ can be found using duality in $L$:
$y=\frac{1}{n\gamma}\cdot\frac{q_{1}}{t}-\frac{1-n\gamma}{n\gamma}\cdot\frac{a}{\sqrt{1-t^{2}}}.$
Indeed,
$\left\langle y,v_{n+1}\right\rangle=\frac{1-n\gamma}{n\gamma}\cdot\frac{1}{\sqrt{1-t^{2}}}\left\langle a,v_{n+1}\right\rangle=1,$
and
$\left\langle y,v_{1}\right\rangle=\frac{1}{n\gamma}\cdot\left\langle\frac{q_{1}}{t},tq_{1}\right\rangle-\frac{1-n\gamma}{n\gamma}\cdot\left\langle\frac{a}{\sqrt{1-t^{2}}},\sqrt{1-t^{2}}a\right\rangle=\frac{1}{n\gamma}-\frac{1-n\gamma}{n\gamma}=1.$
The lemma follows. ∎
Lemma A.3.
Let $h$ be the foot of the altitude from the vertex $x_{n+1}$ of the triangle $T$, and let $z$ be the intersection point of the altitude $x_{n+1}h$ and the angle bisector at vertex $y$. Then the inradius $R(t)$ of ${\Delta}^{\circ}$ equals $\left|h-z\right|$. Moreover,
$R(t)=\frac{1}{n\gamma}\cdot\frac{1}{1+\sqrt{1-t^{2}}}.$
Proof.
Consider a regular simplex $K$ centered at the origin with a facet $G$ parallel to ${a}^{\bot}$. Let $g$ be the vertex opposite to $G$, and let $v$ be any vertex of $G$. By symmetry, the origin is the incenter of $K$. In particular, the altitude from $v$ passes through the origin and lies in the plane $L=\operatorname{span}\{g,v\}$.
After applying a directional dilation along the direction of $a$ that maps $K$ to a simplex $K^{\prime}$, the incenter of $K^{\prime}$ remains on the line with direction vector $a$. Consequently, the incenter and its orthogonal projection onto the facet opposite to $v$ remain in the plane $L$.
In our setting, this shows that the incenter of ${\Delta}^{\circ}$ lies in $T$ and is the intersection point $z$ of the altitude $x_{n+1}h$ and the angle bisector from vertex $y$.
Let us now compute $R(t)$. We use identities from Lemma A.1 and Lemma A.2 for $x_{1}$, $x_{n+1}$, and $y$.
By construction, the line passing through $y$ and $x_{1}$ is orthogonal to $a$, thus
$h=-\frac{1-n\gamma}{n\gamma}\cdot\frac{a}{\sqrt{1-t^{2}}}.$
Hence,
$\left|y-h\right|=\left|\frac{1}{n\gamma}\cdot\frac{q_{1}}{t}\right|=\frac{1}{n\gamma}\cdot\frac{1}{t}$
and
$\left|h-x_{n+1}\right|=\frac{1}{\sqrt{1-t^{2}}}\left(1+\frac{1-n\gamma}{n\gamma}\right)=\frac{1}{n\gamma}\cdot\frac{1}{\sqrt{1-t^{2}}}.$
From the right triangle $x_{n+1}hy$,
$\left|y-x_{n+1}\right|=\sqrt{\left|y-h\right|^{2}+\left|h-x_{n+1}\right|^{2}}=\frac{1}{n\gamma}\cdot\sqrt{\frac{1}{1-t^{2}}+\frac{1}{t^{2}}}.$
Since $yz$ is an angle bisector of the triangle $yhx_{n+1}$, we have:
$\frac{\left|h-z\right|}{\left|h-y\right|}=\frac{\left|z-x_{n+1}\right|}{\left|y-x_{n+1}\right|}.$
Therefore, the inradius of the simplex $x_{1},\dots,x_{n+1}$ is
$R(t)=\left|h-z\right|=\frac{\left|y-h\right|}{\left|y-h\right|+\left|y-x_{n+1}\right|}\cdot\left|h-x_{n+1}\right|.$
Substituting the computed values, we obtain:
$R(t)=\frac{1}{n\gamma}\cdot\frac{1}{\sqrt{1-t^{2}}}\cdot\frac{\frac{1}{t}}{\frac{1}{t}+\sqrt{\frac{1}{1-t^{2}}+\frac{1}{t^{2}}}}=\frac{1}{n\gamma}\cdot\frac{1}{\sqrt{1-t^{2}}+\sqrt{t^{2}+1-t^{2}}}=\frac{1}{n\gamma}\cdot\frac{1}{1+\sqrt{1-t^{2}}}.$
Denote $\sigma(t)=\sum\limits_{i\in[n]}\gamma\left|v_{i}\right|^{2}+(1-n\gamma)\left|v_{n+1}\right|^{2}.$ Using the obtained expressions:
$\sigma(t)=n\gamma+\frac{\left(n\gamma\right)^{2}}{1-n\gamma}\left(1-t^{2}\right)=n\gamma\left(1+\frac{n\gamma}{1-n\gamma}\left(1-t^{2}\right)\right).$
Theorem A.4.
Fix non-negative integers $d,s$ and $m$ with $m>d-s\geq 2$ and a positive $\varepsilon$. Then there is a configuration of $m$ vectors $b_{1},\dots,b_{m}$ in $\mathbb{R}^{d-s}$ and positive numbers $\beta_{1},\dots,\beta_{m}$ satisfying
$\sum\limits_{i\in[m]}\beta_{i}=1,\quad\sum\limits_{i\in[m]}\beta_{i}b_{i}=\mathbf{0},\quad\sum\limits_{i\in[m]}\beta_{i}\left|b_{i}\right|^{2}=\frac{d-s}{d},$
and such that the inradius of the polar of the convex hull of $\{b_{1}\dots,b_{m}\}$ is at least $(1-\varepsilon)\frac{d}{d-s}.$
Proof.
Set $\gamma=\frac{1}{d}$ and $n=d-s$. Then, with the notation introduced previously, inequality (A.1) holds and the following hold as $t$ tends to $1^{-}$:
  1. (1)
    the left-hand side of inequality (A.3) tends to zero;
  2. (2)
    $\sigma(t)$ is monotonically decreasing to $\frac{n}{d};$
  3. (3)
    $R(t)\to\frac{d}{n}.$
The computations of limits are straightforward.
For each $i\in[n],$ set $b_{i}=\sqrt{\frac{\theta}{\sigma(t)}}v_{i}$ with $\theta=\frac{d-s}{d}.$ Set $b_{n+1}=\ldots=b_{m}=\sqrt{\frac{\theta}{\sigma(t)}}v_{n+1}$. Also set $\beta_{i}=\gamma$ for $i\in[n]$ and $\beta_{i}=\frac{1-n\gamma}{m-n}$ for $i\in[m]\setminus[n]$. Then clearly, $\sum\limits_{i\in[m]}\beta_{i}=1$. Moreover,
[…]
and
$\sum\limits_{i\in[m]}\beta_{i}\left|b_{i}\right|^{2}=\gamma\sum_{i\in[n]}\left|b_{i}\right|^{2}+\frac{(1-n\gamma)}{m-n}\sum_{i\in[m]\setminus[n]}\left|b_{i}\right|^{2}=\theta.$
The inradius of the polar of $\operatorname{conv}\{b_{1},\dots,b_{m}\}$ is greater than that of the simplex $v_{1}\dots v_{d-s+1}$ for $t$ sufficiently close to 1.1.1 . On the other hand, there is a $t$ such that the latter is at least $(1-\varepsilon)\frac{d}{d-s}.$
Appendix B Majorization inequality for the volume bound
The following lemma yields (5.3) with the substitution $\left|P_{F}u_{i}\right|^{2}=x_{i}$ and $d=d^{\prime}.$ We leave it to the reader that the constraints on $x_{i}$ are satisfied.
Lemma B.1.
The functional $W(x_{1},\dots,x_{m},\delta_{1},\dots,\delta_{m})=\prod\limits_{i\in[m]}x_{i}^{\frac{\delta_{i}x_{i}}{2}}$ under the constraints
  • $\frac{1}{d+1}\leq x_{i}\leq 1\quad\forall\,i\in[m],$
  • $\sum\limits_{i\in[m]}\delta_{i}x_{i}=s+1,$
  • $\sum\limits_{i\in[m]}\delta_{i}=d+1,$
  • $0\leq\delta_{i}\leq 1$
on the variables satisfies the inequality
$W(x_{1},\dots,x_{m},\delta_{1},\dots,\delta_{m})\leq\frac{(s+1)^{\frac{s+1}{2(d+1)}}}{(d+1)^{\frac{d+1}{2(d+1)}}}.$
Proof.
Set $\Lambda=\frac{s+1}{d+1}.$ We will study the logarithm of our functional. We have to show that
$\ln W(x_{1},\dots,x_{m},\delta_{1},\dots,\delta_{m})\leq\frac{1}{2}\Lambda\ln\Lambda-\frac{d-s}{2(d+1)}\ln(d+1).$
Note that the function $\ln W$ is continuous on the compact set in $\mathbb{R}^{2m}$ defined by our constraints. Hence, it attains its maximum. For every $x=(x_{1},\ldots,x_{m})$ with $\frac{1}{d+1}\leq x_{i}\leq 1$ for all $i\in[m]$, let $\ln W_{x}(\delta_{1},\dots,\delta_{m})$ be the function
$\ln W_{x}(\delta_{1},\dots,\delta_{m})=\frac{1}{2}\sum\limits_{i\in[m]}\delta_{i}x_{i}\ln x_{i}.$
Notice that $\ln W_{x}$ is a convex function of $\delta=(\delta_{1},\dots,\delta_{m}).$ Since the set
[…]
is a compact convex set, $\ln W_{x}$ attains its maximum on some extreme point of $A_{x}$. These are the points of intersection of the 2-dimensional faces of the cube
[…]
with the $(m-2)$-dimensional affine subspace
$\left\{\delta\in\mathbb{R}^{m}:\sum\limits_{i\in[m]}\delta_{i}x_{i}=s+1,\quad\sum\limits_{i\in[m]}\delta_{i}=d+1\right\}.$
Therefore, a maximizer of the function $\ln W_{x}$ has to be a point of the form
[…]
for some $\frac{1}{2}\leq\lambda\leq 1$ (or a permutation of it), such that $\sum\limits_{i\in[m]}\delta_{i}x_{i}=s+1$ is satisfied. For every $\delta_{\lambda}$ with $\frac{1}{2}\leq\lambda\leq 1$, we will find the maximizer of the function
$\ln W_{\delta_{\lambda}}(x)=\frac{1}{2}\sum\limits_{i\in[m]}\delta_{\lambda,i}x_{i}\ln x_{i}$
on the compact convex set
$B_{\lambda}=\left\{x\in\mathbb{R}^{m}:\sum\limits_{i\in[m]}\delta_{\lambda,i}x_{i}=s+1,\quad\frac{1}{d+1}\leq x_{i}\leq 1\quad\forall\,i\in[m]\right\}.$
Set
$\tilde{x}=\left(\underbrace{1,1,\ldots,1}_{s},\ \Lambda,\ \underbrace{\frac{1}{d+1},\ldots,\frac{1}{d+1}}_{m-s-1}\right).$
We check that $\frac{1}{d+1}\leq\Lambda\leq 1$ and
$\sum\limits_{i\in[m]}\delta_{\lambda,i}\tilde{x}_{i}=s+\Lambda+\frac{d-s}{d+1}=s+1.$
It is easy to see that $\tilde{x}$ majorizes any $x=(x_{1},\dots,x_{m})\in B_{\lambda}.$ Thus, by the weighted Karamata inequality (see [Fuc47]), $\ln W_{\delta_{\lambda}}(x)\leq\ln W_{\delta_{\lambda}}\left(\tilde{x}\right)$ for all $x\in B_{\lambda}.$ By the definition of $\delta_{\lambda},$ we conclude that the maximum under our constraints is equal to
$\max\limits_{\lambda\in[1/2,1]}\ln W_{\delta_{\lambda}}(x)=\max\limits_{\lambda\in[1/2,1]}\ln W_{\delta_{\lambda}}=\frac{1}{2}\Lambda\ln\Lambda-\frac{d-s}{2(d+1)}\ln(d+1).$
The lemma follows. ∎
Acknowledgements
We thank Alexander Litvak and Roman Karasev for the fruitful conversations. We are also grateful to Susanna Dann, who organized the ‘Analysis and Convex Geometry Week at UniAndes’ in Bogotá in 2025, where the initial steps of this work were made.
G. I. is supported by Projeto Paz and by CAPES (Coordenação de Aperfeiçoamento de Pessoal de Nível Superior) - Brasil, grant number 23038.015548/2016-06. Z. L. and M. N. are partially supported by the NRDI research grant K147544. M. N. is also supported by the NRDI research grant K131529 as well as the ELTE TKP 2021-NKTA-62 funding scheme. Z. L. is supported also by the ERC Advanced Grant “ERMiD”, Á. S. is funded by the grant NKFIH 150151, a project that has been implemented with the support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the ADVANCED_24 funding scheme.